### v1

# On a pinning model in correlated Gaussian random environments

Abstract.
We consider a pinning model in correlated Gaussian random environments. For the model that is disorder relevant, we study its intermediate disorder regime and show that the rescaled partition functions converge to a non-trivial continuum limit in the Skorohod setting and in the Stratonovich setting, respectively. Our results partially confirm the prediction of Weinrib and Halperin [21] for disorder relevance/irrelevance.
1. Introduction
In this paper, we study a pinning model in Gaussian random environments (also called the disorder), of which the correlation has a power-law decay. We shall show that, under proper scalings, the rescaled partition functions converge weakly to a continuum limit which can be represented by a chaos expansion. Our results complement those of Berger [2] and are consistent with the prediction of Weinrib and Halperin [21] on the disorder relevance/irrelevance for physical systems with correlated disorder.
1.1. The disordered pinning model
The disordered pinning model is a type of random polymer models, which is defined via a Gibbs transform of the distribution of a renewal process. Pinning models are well suited models for studying phase transitions and critical phenomena and also have received widespread attention in physics, chemistry and biology. We refer to the monographs of Giacomin [9] and [10] for more details.
We now introduce the setting of our model. Let $\tau=\{\tau_{i}\text{,}~{}i\in\mathbb{N}_{0}\}$ be a persistent renewal process (i.e. ${\mathbb{P}}_{\tau}(\tau_{1}<\infty)=1$) on $\mathbb{N}$ with $\tau_{0}=0$ and the inter-arrival law
${\mathbb{P}}_{\tau}(\tau_{1}=n)=\frac{L(n)}{n^{1+{\alpha}}},\quad\text{for}~{}n\geq 1,$ (1.1)
where $\alpha>0$ and $L(\cdot):\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a slowly varying function, i.e., $\lim_{t\to\infty}L(at)/L(t)=1$ for any $a>0$ (see [5] for more details about slowly varying functions). The expectation with respect to ${\mathbb{P}}_{\tau}$ is denoted by $\mathbb{E}_{\tau}$. Note that $\mathbb{E}_{\tau}[\tau_{1}]<\infty$ for $\alpha>1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$ for $\alpha<1$, and by [5, Proposition 1.5.9a], for $\alpha=1$,
$\mathbb{E}_{\tau}[\tau_{1}\mathbf{1}_{\{\tau_{1}\leq N\}}]=l(N):=\sum_{n=1}^{N}\frac{L(n)}{n}$ (1.2)
is also slowly varying (whether it is finite or not depending on $L(\cdot)$). Hence, the renewal process $\tau$ can be classified by the following conditions:
(1) $10<{\alpha}<$ 1; (2) $\mathbb{E}_{\tau}[\tau_{1}]<\infty$; (3) $\alpha=1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$. (1.3)
Noting that our renewal process $\tau$ is aperiodic since $L(n)>0$ for all $n$, we have the renewal theorem (see [8] and [5, Theorem 8.7.5]):
${\mathbb{P}}_{\tau}(n\in\tau)\sim\begin{cases}\dfrac{C_{\alpha}}{L(n)n^{1-\alpha}},&\text{if }0<\alpha<1~{}(C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}),\\[10.0pt] \dfrac{1}{\mathbb{E}[\tau_{1}]},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \dfrac{1}{l(n)},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}\quad\qquad\text{as}~{}n\to\infty.$ (1.4)
Let $\omega=\{\omega_{n}\text{,}~{}n\in\mathbb{N}\}$ be a family of Gaussian random variables independent of $\tau$, representing the random environments. The probability and expectation for $\omega$ are denoted by ${\mathbb{P}}_{\omega}$ and $\mathbb{E}_{\omega}$, respectively. We assume that
$\mathbb{E}_{\omega}[\omega_{n}]=0\quad\text{and}\quad\mathbb{E}_{\omega}[\omega_{n}\omega_{n^{\prime}}]=\gamma(n-n^{\prime}),\quad\text{for }n,n^{\prime}\in\mathbb{N},$ (1.5)
where the covariance function $\gamma(\cdot)$ satisfies
[…]
Here, $H\in(1/2,1)$ is a fixed constant (called the Hurst parameter) throughout this paper. We further assume that for all $t\neq 0$,
$\lim\limits_{N\rightarrow\infty}N^{2-2H}\gamma(\left[Nt\right])=|t|^{2H-2}.$ (1.7)
We shall use ${\mathbb{P}}={\mathbb{P}}_{\tau}\otimes{\mathbb{P}}_{\omega}$ and $\mathbb{E}=\mathbb{E}_{\tau}\otimes\mathbb{E}_{\omega}$ on the product probability space of $(\tau,\omega)$. We may also use ${\mathbb{P}}$ and $\mathbb{E}$ to denote the probability and the expectation respectively of some random variables only involving $\tau$ or $\omega$ when there is no ambiguity.
Denote $\Omega_{N}:=\Big{\{}\frac{1}{N}\text{,}\frac{2}{N}\text{,}\cdots\text{,}\frac{N-1}{N}\text{,}1\Big{\}}$ and $\delta_{x}^{(N)}=\delta_{x}:=\mathbf{1}_{\{Nx\in\tau\}}$. The disordered pinning model with free boundary is defined by
$\frac{{\mathrm{d}}{\mathbb{P}}_{N,\beta}^{\omega,h}}{{\mathrm{d}}{\mathbb{P}}_{\tau}}(\tau):=\frac{1}{Z_{N,\beta}^{\omega,h}}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big{\}},$
where
$Z_{N,\beta}^{\omega,h}=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big{\}}\Big{]}$ (1.8)
is the partition function, $\beta>0$ is the inverse temperature and $h\in\mathbb{R}$ is an external field. One can also consider the model with conditioned boundary, of which the partition function is given by
$Z_{N,\beta}^{\omega,h,\text{c}}=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big{\}}\Big{|}N\in\tau\Big{]}.$ (1.9)
1.2. Disorder relevance/irrelevance and the Weinrib-Halperin prediction
When the disorder is i.i.d., the Harris criterion [11] indicates that disorder is irrelevant if ${\alpha}<\frac{1}{2}$ and is relevant if ${\alpha}>\frac{1}{2}$ (with $\alpha=\frac{1}{2}$ being the marginal case). When the disorder is correlated and the correlation is of the form $n^{-p}$ ($p>0$ and $n$ is the distance between two points), Weinrib and Halperin [21] predicted that the disorder should be relevant if ${\alpha}>\min(p,1)/2$ and irrelevant if ${\alpha}<\min(p,1)/2$. Berger [2] confirmed the Weinrib-Halperin prediction for the case $p>1$ (note that in this case the Harris criterion still applies), and showed that the case $p<1$ is different in the sense that classical methods (see, e.g., [10]) in the study of disorder relevance/irrelevance by comparing critical points/exponents no longer apply: when $p<1$, the disordered system does not possess a localization/delocalization phase transition, since the annealed free energy is always infinite and the quenched free energy is always strictly positive for all $h$ as long as $\beta>0$, and thus the system is always localized (see Theorem 2.5 and Section 2.3 in [2]). This phenomenon is due to the fact, as pointed out by Berger [3], that when $p<1$, there are large and frequent regions in the environments with arbitrarily high attractions (called the “infinite disorder” in [3]). These results hint at scalings of $\beta$ and $h$ when studying disorder relevance/irrelevance for the case $p<1$.
On the other hand, Caravenna, Sun and Zygouras [6] proposed a new perspective on disorder relevance/irrelevance: noting that disorder relevance means that an arbitrarily small amount of disorder (i.e., for any $\beta>0$) changes the features of the underlying homogeneous model, it is possible that if the strength of disorder is tuned down to zero at a proper rate (i.e., $\beta_{N}\to 0$) as the system size tends to infinity, then the disorder will persist in the resulting continuum model. This phenomenon has been shown for the i.i.d. disorder by studying weak limits of rescaled partition functions for the directed polymer model in [1] and for pinning models, directed polymer models and the Ising model in a unified way in [6]. For related results on directed polymers, we refer to [18, 7] for the case when the disorder is correlated in space but independent in time, to [19] for the case when the disorder is correlated in time but independent in space, and to [20] for the case when the disorder is space-time correlated. In those models, the Weinrib-Halperin prediction is not clear (the decay rates in space and in time of the correlation of the disorder in [20] do not agree).
Inspired by the above-mentioned works, we study the weak convergence of the rescaled partition functions. In our setting, noting that $p=2-2H\in(0,1)$ with $H\in(\frac{1}{2},1)$, the Weinrib-Halperin prediction suggests that the disorder relevent if $\alpha>1-H$, which coincides with the condition for the existence of the weak limit for each chaos (see Proposition 3.1).
Finally, we make a few more remarks on disorder relevance/irrelevance for pinning models in correlated random environment. Poisat [17] showed that the Harris criterion still applies to the pinning model in random environments with finite range correlation. When disorders have long range correlation, Berger and Poisat [4] conformed the Weinrib-Halperin prediction for the case when the correlations are summable, which extended the result in [2].
1.3. Main results
Noting that
$\mathbb{E}_{\omega}\Big{[}\exp\Big{\{}\sum\limits_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big{\}}\Big{]}=\exp\Big{\{}\frac{1}{2}\mathrm{Var}_{\omega}\Big{(}\sum\limits_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big{)}\Big{\}},$
we first consider the following Wick ordered partition function
$\tilde{Z}_{N,\beta}^{\omega,h}:=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}_{\omega}\Big{(}\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big{)}\Big{\}}\Big{]}.$ (1.10)
Note that the term $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{1}{2}\beta^{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n,m\in\tau\}}$ is a direct normalization of the partition function, instead of the Gibbs measure itself, since it depends on the realization of $\tau$. We adopt this normalization because the disorders are correlated, and thus a polynomial chaos expansion as has been done in [1, 6] cannot be employed for our pinning model. The normalization in the Wick ordered partition function (1.10) results in a Wick exponential, which allows us to perform a Wick expansion (see Section 3 and also the discussion in [13]). We point out that the Wick ordered partition function (1.10) reduces to [6, equation (3.3)] when the disorders are i.i.d., noting that $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{\beta^{2}}{2}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}$.
Let $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$ be fixed constants throughout this paper. Consider the following scalings:
$\beta_{N}=\begin{cases}\hat{\beta}\dfrac{L(N)}{N^{\alpha+H-1}},&\text{if }0<\alpha<1,\\[10.0pt] \hat{\beta}\dfrac{1}{N^{H}},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \hat{\beta}\dfrac{l(N)}{N^{H}},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}h_{N}=\begin{cases}\hat{h}\dfrac{L(N)}{N^{\alpha}},&\text{if }0<\alpha<1,\\[10.0pt] \hat{h}\dfrac{1}{N},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \hat{h}\dfrac{l(N)}{N},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}$ (1.11)
where we recall that $l(n)$ given in (1.2) is a slowly varying function.
Let $\phi_{k}(t_{1},\dots,t_{k})$ be a symmetric function, defined for $10<t_{1}<\cdots<t_{k}<$ 1 by
$\phi_{k}(t_{1},\dots,t_{k}):=\begin{cases}\displaystyle\frac{C_{\alpha}^{k}}{t_{1}^{1-\alpha}(t_{2}-t_{1})^{1-\alpha}\cdots(t_{k}-t_{k-1})^{1-\alpha}},&\text{ if }0<\alpha<1,\\[10.0pt] \displaystyle\Big{(}\frac{1}{\mathbb{E}[\tau_{1}]}\Big{)}^{k},&\text{ if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \displaystyle 1,&\text{ if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}$ (1.12)
with $C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}$ as in (1.4).
Now we present our main results for the Wick ordered partition function $\tilde{Z}_{N,\beta}^{\omega,h}$ given in (1.10) and for the original partition function $Z_{N,\beta}^{\omega,h}$ given in (1.8), respectively.
Theorem 1.1.
Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H and $\alpha$ be parameters satisfying
$H\in\Big{(}\frac{1}{2},1\Big{)}\text{ and }~{}\alpha>\frac{1}{2}.$
Recalling the Wick ordered partition function $\tilde{Z}^{\omega,h}_{N,\beta}$ given in (1.10) and the scalings (1.11), we have that
$\tilde{Z}^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}~{},\text{ as }N\to\infty,$
where $\tilde{Z}_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
$\tilde{Z}_{\hat{\beta},\hat{h}}:=1+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod\limits_{i=1}^{k}\diamond\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right),$ (1.13)
where $\phi_{k}$ is given by (1.12), $\dot{W}$ is Gaussian noise with $\mathbb{E}[\dot{W}(t)\dot{W}(s)]=|t-s|^{2H-2}$ and the stochastic integral is of Skorohod type.
Theorem 1.2.
Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H and $\alpha$ be parameters satisfying
$H\in\Big{(}\frac{1}{2},1\Big{)}\text{ and }~{}\alpha+H>\frac{3}{2}.$
Recalling the partition function $Z^{\omega,h}_{N,\beta}$ given in (1.8) and the scalings (1.11), we have that
$Z^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}Z_{\hat{\beta},\hat{h}}~{},~{}as~{}N\to\infty,$
where $Z_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
$Z_{\hat{\beta},\hat{h}}:=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod_{i=1}^{k}\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right),$ (1.14)
where $\phi_{k}$ is given by (1.12), $\dot{W}$ is Gaussian noise with $\mathbb{E}[\dot{W}(t)\dot{W}(s)]=|t-s|^{2H-2}$, and the stochastic integral is of Stratonovich type.
Remark 1.3.
The results in Theorems 1.1 and 1.2 also hold for the conditioned partition functions
$\tilde{Z}_{N}^{\text{c}}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N},\text{c}}:=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}\Big{[}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx})\delta_{x}\Big{]}\Big{\}}\Big{|}N\in\tau\Big{]}$
and
${Z}_{N}^{\text{c}}={Z}_{N,\beta_{N}}^{\omega,h_{N},\text{c}}:=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big{\}}\Big{|}N\in\tau\Big{]},$
with the integrand $\phi_{k}$ in (1.13) and (1.14) being replaced by a symmetric function $\phi_{k}^{c}$, defined for $10<t_{1}<\cdots<t_{k}<$ 1 by
$\phi_{k}^{c}(t_{1},\dots,t_{k}):=\begin{cases}\displaystyle\frac{C_{\alpha}^{k}}{t_{1}^{1-\alpha}(t_{2}-t_{1})^{1-\alpha}\cdots(t_{k}-t_{k-1})^{1-\alpha}(1-t_{k})^{1-\alpha}},&\text{ if }1-H<\alpha<1,\\[10.0pt] \displaystyle\Big{(}\frac{1}{\mathbb{E}[\tau_{1}]}\Big{)}^{k},&\text{ if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \displaystyle 1,&\text{ if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty.\end{cases}$
Remark 1.4.
A natural condition for the rescaled Wick ordered partition function to converge weakly is $\alpha+H-1>0$, where $\alpha+H-1$ is the scaling exponent for $\beta$ (see (1.11)). Note that for directed polymers, under the condition that the corresponding scaling exponent is positive, the continuum partition converges in $L^{2}$ as a Wiener chaos (see [6, 7, 19]). However, for our pinning model, assuming $\alpha+H-1>0$ we can only prove that each chaos is a well-defined Skorohod integral which is $L^{2}$-integrable (see Lemma 3.2), and we need a stronger condition $H\in(\frac{1}{2},1),\alpha>\frac{1}{2}$ to guarantee the $L^{2}$-convergence as stated in Theorem 1.1 (also see Remark 3.3 for the criticality of $\alpha=\frac{1}{2}$). This is a new phenomenon that does not appear in the models studied in [1, 6, 18, 19, 7]: in those models, the rescaled partition functions are always uniformly bounded in $L^{2}$ if the models are in the disorder relevant regime, no matter the random environments are independent or correlated. This hints that the pinning model in correlated environment indeed carries some new features.
Motivated by the Weinrib-Halperin prediction, we conjecture that under the condition $\alpha+H-1>0$ with $\alpha\leq\frac{1}{2}$, the rescaled Wick ordered partition function $\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}$ converges weakly to some $L^{1}$-integrable random variable. We will discuss further on the uniform integrability and the weak limit of $\big{\{}\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big{\}}_{N}$ in Section 3.2.
To close this section, let us highlight our contributions in this paper. Firstly, we obtain some finer property of a disordered pinning model with environment correlation of the form $n^{-p}$, where $p<1$, which complements the result of Berger [2]. To be specific, we find scaling limits of rescaled (Wick ordered) partition functions in the disorder relevant regime. In particular, the Weinrib-Halperin prediction is partially conformed when it differs from the Harris criterion. We also show that the rescaled (Wick ordered) partition functions are not $L^{2}$-bounded in the entire disorder relevance regime (see Remark 3.3), which is a new feature compared to other disordered models. Secondly, we develop a decoupling approach to prove uniform integrability (see Section 3.2), and provide some sharp upper bounds for the norm of traces (see Lemma 4.1) and for the second moments of multiple Stratonovich integrals (see Lemma 4.4). These auxiliary results may be of independent interest.
Notations 1.5.
[…]
The rest of the paper is organized as follows. In Section 2, we introduce some preliminaries. The proofs of Theorem 1.1 and Theorem 1.2 will be presented in Section 3 and Section 4, respectively. We will only deal with the case $10<\alpha<$ 1 in the proofs, noting that the other case $\alpha\geq 1$ can be treated in the same way. Some technical estimates are put in Appendix A.
2. Preliminaries on Gaussian spaces and U-statistics
2.1. Gaussian spaces
In this subsection, we collect some preliminaries on the analysis of Gaussian space. We refer to [15, 12] for more details.
On a probability space $(\Omega,\mathcal{F},{\mathbb{P}})$ satisfying usual conditions, let $\dot{W}=\{\dot{W}(t):t\in[0,1]\}$ be a real-valued centered Gaussian noise with covariance
$\mathbb{E}[\dot{W}(t)\dot{W}(s)]=K(t-s)=|t-s|^{2H-2},$ (2.1)
where $H\in(\frac{1}{2},1)$ is the Hurst parameter appearing in (1.6) and (1.7). Thus, $\dot{W}$ is the continuum counterpart of the disorder $\omega$ introduced in Section 1.1.
The Hilbert space $\mathcal{H}$ associated with $\dot{W}$ is the completion of smooth functions with compact support under the inner product:
$f,g\right>_{\mathcal{H}}$ $=\int_{0}^{1}\int_{0}^{1}f(s)K(s-t)g(t){\mathrm{d}}s{\mathrm{d}}t=\int_{0}^{1}\int_{0}^{1}f(s)g(t)|s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t.$ (2.2)
Noting that $\mathcal{H}$ contains distributions (see [16]), it is convenient to consider $|\mathcal{H}|$ which is a linear subspace of measurable functions $f$ on $[0,1]$ satisfying
$\|f\|_{|\mathcal{H}|}^{2}:=\int_{0}^{1}\int_{0}^{1}|f(s)g(t)||s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t<\infty.$ (2.3)
Clearly, $|\mathcal{H}|$ is a Banach space and is dense in $\mathcal{H}$.
Let $\{W(f),f\in\mathcal{H}\}$ be an isonormal Gaussian process with covariance $\mathbb{E}[W(f)W(g)]=\left\langle f,g\right\rangle_{\mathcal{H}}.$ The Wiener integral $W(f)$ is often denoted in the integral form:
$\int_{0}^{1}f(s)W({\mathrm{d}}s):=W(f).$
For $n\in\mathbb{N}_{0}$, the $n$-th Hermite polynomial is given by
$H_{n}(x):=(-1)^{n}\mathrm{e}^{x^{2}/2}\frac{{\mathrm{d}}^{n}}{{\mathrm{d}}x^{n}}\mathrm{e}^{-x^{2}/2},\,x\in\mathbb{R}.$
For $f\in\mathcal{H}$ with $\|f\|_{\mathcal{H}}=1$, the multiple Wiener integral of $f^{\otimes n}$ is defined by
$\mathbf{I}_{n}(f^{\otimes n}):=H_{n}(W(f)).$ (2.4)
Let ^${\mathcal{H}}^{\hat{\otimes}k}$ be the subspace of ${\mathcal{H}}^{\otimes k}$ containing only symmetric functions. For $f\in{\mathcal{H}}^{\hat{\otimes}k}$, the $k$-th multiple Wiener integral $\mathbf{I}_{k}(f)$ can be defined by (2.4) via a limiting argument, and moreover,
${\mathbb{E}}[\mathbf{I}_{m}(f)\mathbf{I}_{n}(g)]=n!\langle f,g\rangle_{\mathcal{H}^{\otimes n}}\mathbf{1}_{\{m=n\}},~{}{\rm for}~{}f\in{\mathcal{H}}^{\hat{\otimes}m},g\in{\mathcal{H}}^{\hat{\otimes}n}.$
For general $f\in{\mathcal{H}}^{\otimes k}$, we simply set $\mathbf{I}_{k}(f):=\mathbf{I}_{k}(\hat{f}),$ where $\hat{f}$ denotes the symmetrization of $f$. We also use the following integral form for multiple Wiener integrals:
$\int_{[0,1]^{n}}f({\mathbf{t}})W({\mathrm{d}}{t_{1}})\diamond\cdots\diamond W({\mathrm{d}}{t_{n}})=\int_{[0,1]^{n}}f({\mathbf{t}})\prod_{i=1}^{n}\diamond W({\mathrm{d}}{t_{i}}):=\mathbf{I}_{n}(f).$
For $f\in\mathcal{H}^{\otimes n}$ and $g\in\mathcal{H}^{\otimes m}$, the Wick product of $\mathbf{I}_{n}(f)$ and $\mathbf{I}_{m}(g)$ is defined by
$\mathbf{I}_{n}(f)\diamond\mathbf{I}_{m}(g):=\mathbf{I}_{n+m}(f\otimes g).$ (2.5)
For square integrable random variables $F$ and $G$ with Wiener chaos expansions $F=\sum_{k=0}^{\infty}\mathbf{I}_{k}(f_{k})$ and $G=\sum_{k=0}^{\infty}\mathbf{I}_{k}(g_{k})$ respectively, we have $F\diamond G=\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\mathbf{I}_{n+m}(f_{n}\otimes g_{m})$, as long as the right-hand side term exists in some proper sense. For a centered Gaussian random variable $F$, its Wick exponential is defined by
$\exp^{\diamond}(F):=\exp\left(F-\frac{\mathbb{E}[F^{2}]}{2}\right)=\sum_{m=0}^{\infty}\frac{1}{m!}F^{\diamond m}.$ (2.6)
The multiple Stratonovich integral of $g^{\otimes n}$ for $g\in\mathcal{H}$ is defined by $\mathbb{I}_{m}(g^{\otimes n}):=W(g)^{n},$ and $\mathbb{I}_{n}(g)$ for $g\in\mathcal{H}^{\hat{\otimes}n}$ by a limiting argument. We also take the following notation for $\mathbb{I}_{n}(f)$
$\int_{[0,1]^{n}}f({\mathbf{t}})W({\mathrm{d}}{t_{1}})\cdots W({\mathrm{d}}{t_{n}})=\int_{[0,1]^{n}}f({\mathbf{t}})\prod_{i=1}^{n}W({\mathrm{d}}{t_{i}}):=\mathbb{I}_{n}(f).$
For $f\in\mathcal{H}^{\otimes n}$ and $k=0,1,\cdots,\lfloor\frac{n}{2}\rfloor$, define the $k$-th order trace $\mathrm{Tr}^{k}f$ of $f$ by
[…]
where $K(x-y)$ is given in (2.1). Multiple Stratonovich integrals $\mathbb{I}_{m}(f)$ and multiple Wiener integrals $\mathbf{I}_{m}(f)$ are related via the following celebrated Hu-Meyer’s formula:
$\mathbb{I}_{n}(f)=\sum_{k=0}^{[\frac{n}{2}]}\frac{n!}{k!(n-2k)!2^{k}}\mathbf{I}_{n-2k}(\mathrm{Tr}^{k}\hat{f}),\quad\text{for}~{}f\in\mathcal{H}^{\otimes n},$ (2.8)
as long as $\vvvert f\vvvert_{n}<\infty$, where
$\vvvert f\vvvert^{2}_{n}:=\mathbb{E}[|\mathbb{I}_{n}(f)|^{2}]=\sum_{k=0}^{[\frac{n}{2}]}\frac{1}{(n-2k)!}\left(\frac{n!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\hat{f}\|^{2}_{\mathcal{H}^{\otimes(n-2k)}}.$ (2.9)
2.2. U-statistics
Definition 2.1.
We denote
[…]
Given $f\in|\mathcal{H}|^{\otimes m}$, we denote by $\mathcal{A}_{N}(f)$ its conditional expectation with respect to the $\sigma$-field generated by $\mathcal{R}_{N}^{m}$, i.e.,
[…]
where $|B|$ is the Lebesgue measure of B.
We introduce the U-statistics for our model, which appears in the chaos expansion for the Wick ordered partition function (1.10).
Definition 2.2.
For $f\in|\mathcal{H}|^{\otimes m}$, let
$\mathbf{I}_{m}^{(N)}(f)$ $:=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right)\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}$ (2.11)
$:=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)(\boldsymbol{t})\,\omega_{\boldsymbol{n}}^{\diamond},$
where we denote
$\omega_{\boldsymbol{n}}^{\diamond}:=\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}~{}\text{ and }~{}\boldsymbol{t}=(t_{1}\cdots t_{m}):=\frac{\boldsymbol{n}}{N}=\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right).$ (2.12)
Lemma 2.3.
For $m,N\in\mathbb{N}$ and $f\in|\mathcal{H}|^{\otimes m}$, let $\mathcal{A}_{N}(f)$ be defined as above. Then we have,
$\|\mathcal{A}_{N}(f)\|_{|\mathcal{H}|^{\otimes m}}\leq C^{m}\|f\|_{|\mathcal{H}|^{\otimes m}}$
where C is a constant depending only on H.
Proof.
The desired result follows directly from [19, Lemma C.1]. ∎
Lemma 2.4.
There exists a constant $C$ depending only on $H$ such that for all $f\in|\mathcal{H}|^{\otimes m}$ and $N\in\mathbb{N}$,
$\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{m}^{(N)}(f)\Big{)}^{2}\Big{]}\leq C^{m}m!\|\hat{f}\|_{|\mathcal{H}|^{\otimes m}}^{2},$
where $\hat{f}$ is the symmetrization of f.
Proof.
For $\boldsymbol{n},\boldsymbol{n}^{\prime}\in\llbracket N\rrbracket^{m}$, by Lemma A.2, we have:
$\mathbb{E}\left[\omega_{\boldsymbol{n}}^{\diamond}~{}\omega_{\boldsymbol{n^{\prime}}}^{\diamond}\right]=\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big{(}n_{i}-n^{\prime}_{\sigma(i)}\big{)},$
where $\mathcal{P}_{m}$ denotes the set of all permutations on $\llbracket m\rrbracket$. Thus, we have
$\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{m}^{(N)}(f)\Big{)}^{2}\Big{]}$ $=\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{m}^{(N)}(\hat{f})\Big{)}^{2}\Big{]}$
$=N^{-2mH}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big{(}n_{i}-n^{\prime}_{\sigma(i)}\big{)}$
$=m!N^{-2mH}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\prod_{i=1}^{m}\gamma(n_{i}-n^{\prime}_{i}),$
where the last equality follows from the symmetry of $\mathcal{A}_{N}(\hat{f})$.
For $t\in\mathbb{R}$, let
$\gamma_{N}(t):=N^{2-2H}\gamma([|t|N]).$ (2.13)
By (1.6) and (1.7), we have
$\gamma_{N}(t)\lesssim|t|^{2H-2}\text{ and }\lim\limits_{N\rightarrow\infty}\gamma_{N}(t)=|t|^{2H-2}.$
Therefore, we get by a Riemann sum approximation,
$\mathbb{E}\Big{[}\left(\mathbf{I}_{m}^{(N)}(f)\right)^{2}\Big{]}$ $=m!N^{-2m}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\prod_{i=1}^{m}N^{2-2H}\gamma(n_{i}-n^{\prime}_{i})$
$=m!N^{-2m}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\prod_{i=1}^{m}\gamma_{N}\Big{(}\frac{n_{i}-n^{\prime}_{i}}{N}\Big{)}$
$\lesssim m!\int_{[0,1]^{2m}}|\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})|\prod\limits_{i=1}^{m}|t_{i}-t_{i}^{\prime}|^{2H-2}{\mathrm{d}}t_{i}{\mathrm{d}}t_{i}^{\prime}$
$=m!\|\mathcal{A}_{N}(\hat{f})\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq m!C^{m}\|\hat{f}\|_{|\mathcal{H}|^{\otimes m}}^{2},$
where the last inequality follows from Lemma 2.3. ∎
Proposition 2.5.
For $f\in|\mathcal{H}|$, $\mathbf{I}_{1}^{(N)}(f)$ converges weakly to the Wiener integral $\mathbf{I}_{1}(f)$ as $N\rightarrow\infty$. Moreover, for any $k\in\mathbb{N}$ and $f_{1},\cdots,f_{k}\in|\mathcal{H}|$, we have the joint convergence in distribution:
$\Big{(}\mathbf{I}_{1}^{(N)}(f_{1}),\cdots,\mathbf{I}_{1}^{(N)}(f_{k})\Big{)}\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{1}(f_{1}),\cdots,\mathbf{I}_{1}(f_{k})\right),\quad\text{as}~{}N\rightarrow\infty.$
Proof.
First we consider the case where $f$ is an indicator function: $f(x)=\mathrm{I}_{(s,t]}(x)$ with $s,t\in\Omega_{N}$. Then
$\mathcal{A}_{N}(f)(r)=\mathrm{I}_{(s,t]}(r)\quad\text{and}\quad\mathbf{I}_{1}^{(N)}(f)=N^{-H}\sum_{n=1}^{N}\omega_{n}\mathrm{I}_{(Ns,Nt]}(n),$
and we have
$\mathbb{E}_{\omega}\Big{[}\Big{(}\mathbf{I}_{1}^{(N)}(f)\Big{)}^{2}\Big{]}$ $=\mathbb{E}_{\omega}\Big{[}N^{-2H}\Big{(}\sum_{n=1}^{N}\omega_{n}\mathbf{I}_{(Ns,Nt]}(n)\Big{)}^{2}\Big{]}$
$=N^{-2H}\mathbb{E}_{\omega}\Big{[}\sum_{n,n^{\prime}\in(Ns,Nt]}\omega_{n}\omega_{n^{\prime}}\Big{]}$
$=N^{-2H}\sum_{n,n^{\prime}\in(Ns,Nt]}\gamma(n-n^{\prime}).$
Denoting $r=\frac{n}{N},r^{\prime}=\frac{n^{\prime}}{N}$, we get
$\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{1}^{(N)}(f)\Big{)}^{2}\Big{]}=N^{-2}\sum_{r,r^{\prime}\in(s,t]\cap\Omega_{N}}\gamma_{N}(r-r^{\prime})\overset{N\rightarrow\infty}{\longrightarrow}\int_{s}^{t}\int_{s}^{t}|r-r^{\prime}|^{2H-2}{\mathrm{d}}r{\mathrm{d}}r^{\prime}=\|f\|_{\mathcal{H}}^{2}.$
Noting that $\mathbf{I}_{1}^{(N)}(f)$ and $\mathbf{I}_{1}(f)$ are Gaussian random variables and $\mathbf{I}_{1}(f)$ has zero mean and variance $\|f\|_{\mathcal{H}}^{2}$, we have $\mathbf{I}_{1}^{(N)}(f)\overset{(d)}{\longrightarrow}\mathbf{I}_{1}(f)$ as $N\rightarrow\infty.$
Similarly, one can show that $\mathbf{I}_{1}^{(N)}(f)\overset{(d)}{\longrightarrow}\mathbf{I}_{1}(f)$ as $N\rightarrow\infty$ also holds for a simple function $f$. For a general function $f\in|\mathcal{H}|$, since simple functions are dense in $|\mathcal{H}|$, there exists a sequence of simple functions $\{f_{n}\}_{n\in\mathbb{N}}$ converging to $f$ in $|\mathcal{H}|$. On the one hand, for each fixed $n$, we have shown that $\mathbf{I}_{1}^{(N)}(f_{n})\overset{(d)}{\longrightarrow}\mathbf{I}_{1}(f_{n})$ as $N\rightarrow\infty.$ On the other hand, for each fixed $N$, by Lemma 2.4, we have $\mathbf{I}_{1}^{(N)}(f_{n})\rightarrow\mathbf{I}_{1}^{(N)}(f)$ as $n\rightarrow\infty$ in $L^{2}$ uniformly in $N$. Meanwhile, we also have $\mathbf{I}_{1}(f_{n})\rightarrow\mathbf{I}_{1}(f)$ as $n\rightarrow\infty$ in $L^{2}$ by the property of the Wiener integral. Combining the above convergences and by Lemma A.3, we have $\mathbf{I}_{1}^{(N)}(f)\overset{(d)}{\longrightarrow}\mathbf{I}_{1}(f)$ as $N\rightarrow\infty$ for $f\in|\mathcal{H}|$. Finally, the joint convergence follows from the linearity of $\mathbf{I}_{1}^{(N)}$ and the Cramér-Wold Theorem. ∎
Proposition 2.6.
For any $m\in\mathbb{N}$ and $f\in|\mathcal{H}|^{\otimes m}$, we have
$\mathbf{I}_{m}^{(N)}(f)\overset{(d)}{\longrightarrow}\mathbf{I}_{m}(f),\quad\text{as}~{}N\rightarrow\infty,$ (2.14)
where $\mathbf{I}_{m}(f)$ is the $m$-th multiple Wiener integral of f. Moreover, for any $k\in\mathbb{N}$ and $f_{1},\cdots,f_{k}$ with $f_{i}\in|\mathcal{H}|^{\otimes l_{i}}$ for $l_{i}\in\mathbb{N}$, we have the joint convergence in distribution
$\left(\mathbf{I}_{l_{1}}^{(N)}(f_{1}),\cdots,\mathbf{I}_{l_{k}}^{(N)}(f_{k})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{l_{1}}(f_{1}),\cdots,\mathbf{I}_{l_{k}}(f_{k})\right),\quad\text{as}~{}N\rightarrow\infty.$
Proof.
We only prove (2.14). The joint convergence follows from the linearity of $\mathbf{I}_{l}^{(N)}$ and the Cramér-Wold Theorem.
First, we consider $f=h_{1}^{\otimes p_{1}}\otimes\cdots\otimes h_{l}^{\otimes p_{l}}$ with $h_{i}\in|\mathcal{H}|$ and $p_{1}+\cdots+p_{l}=m$. For such $f$, we have
$\mathbf{I}_{m}^{(N)}(f)$ $=N^{-2mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(h_{1}^{\otimes p_{1}}\otimes\cdots\otimes h_{l}^{\otimes p_{l}})\omega_{\boldsymbol{n}}^{\diamond}$
$=N^{-2mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(h_{1})^{\otimes p_{1}}\otimes\cdots\otimes\mathcal{A}_{N}(h_{l})^{\otimes p_{l}}\omega_{\boldsymbol{n}}^{\diamond}$
$=\left(\mathbf{I}_{1}^{(N)}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}^{(N)}(h_{l})\right)^{\diamond p_{l}}$
By Lemma 2.4, we have $\mathbb{E}[(\mathbf{I}_{1}^{(N)}(h_{i}))^{2}]\leq C\|h_{i}\|_{|\mathcal{H}|}^{2}$ uniformly in $N$. Note that $\mathbf{I}_{1}^{(N)}(h_{i})$ is a centered Gaussian random variable, and thus, we have
$\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{1}^{(N)}(h_{i})\Big{)}^{2q}\Big{]}\leq C^{q}(2q-1)!!\|h_{i}\|_{|\mathcal{H}|}^{2q}.$
Combining the joint convergence in Proposition 2.5, we can apply [19, Lemma B.3] to get
$\mathbf{I}_{m}^{(N)}(f)$ $=\left(\mathbf{I}_{1}^{(N)}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}^{(N)}(h_{l})\right)^{\diamond p_{l}}\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{1}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}(h_{l})\right)^{\diamond p_{l}},\quad\text{as}~{}N\rightarrow\infty,$
where the weak limit is equal to, noting (2.5),
$\left(\mathbf{I}_{p_{1}}(h_{1}^{\otimes p_{1}})\right)\diamond\cdots\diamond\left(\mathbf{I}_{p_{l}}(h_{l}^{\otimes p_{l}})\right)=\mathbf{I}_{m}\left(h_{1}^{\otimes p_{1}}\otimes\cdots\otimes h_{l}^{\otimes p_{l}}\right)=\mathbf{I}_{m}(f).$
For a general function $f\in|\mathcal{H}|^{\otimes m}$, we can find a sequence of functions $\{f_{n}\}_{n\in\mathbb{N}}$ which are the linear combinations of the functions of the form $h_{1}^{\otimes p_{1}}\otimes\cdots\otimes h_{l}^{\otimes p_{l}}$ satisfying $f_{n}\rightarrow f$ in $|\mathcal{H}|^{\otimes m}$. Similarly to the proof of Proposition 2.5, we have $\mathbf{I}_{m}^{(N)}(f)\overset{(d)}{\longrightarrow}\mathbf{I}_{m}(f)\text{, as}~{}N\rightarrow\infty$. ∎
3. The Skorohod case
In Section 3.1, we study the weak convergence for the Wick ordered partition function defined in (1.10) and prove Theorem 1.1 which guarantees the weak convergence of the rescaled Wick ordered partition functions in $L^{2}$-regime when $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$. In Section 3.2, we discuss the uniform integrability and characterize the weak limit of the Wick ordered partition functions for the case $\alpha\leq\frac{1}{2}$.
By Taylor expansion and Wick exponential (2.6), we get
$\tilde{Z}_{N,\beta}^{\omega,h}$ $=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}h\sum_{x\in\Omega_{N}}\delta_{x}\Big{\}}\exp\Big{\{}\diamond\Big{(}\beta\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big{)}\Big{\}}\Big{]}$ (3.1)
$=\mathbb{E}_{\tau}\Big{[}\Big{(}\sum_{k=0}^{\infty}\frac{1}{k!}\Big{(}h\sum_{x\in\Omega_{N}}\delta_{x}\Big{)}^{k}\Big{)}\Big{(}\sum_{m=0}^{\infty}\frac{1}{m!}\Big{(}\beta\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big{)}^{\diamond m}\Big{)}\Big{]}$
$=\sum_{k=0}^{\infty}\sum_{m=0}^{\infty}\frac{h^{k}\beta^{m}}{k!m!}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{k+m}}{\mathbb{P}}(n_{1}\in\tau,n_{2}\in\tau,\cdots,n_{k+m}\in\tau)\omega_{n_{1}}\diamond\omega_{n_{2}}\diamond\cdots\diamond\omega_{n_{m}}.$
We shall write $\tilde{Z}_{N}:=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}$ throughout the rest of the paper for simplicity. By (3.1), we have
$\tilde{Z}_{N}=\sum_{k=0}^{\infty}\sum_{m=0}^{\infty}\frac{1}{k!m!}S_{m,k}^{(N)},$ (3.2)
where
$S_{m,k}^{(N)}:=\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{k+m}}{\mathbb{P}}\big{(}n_{1}\in\tau,n_{2}\in\tau,\cdots,n_{k+m}\in\tau\big{)}\,\omega_{n_{1}}\diamond\omega_{n_{2}}\diamond\cdots\diamond\omega_{n_{m}}.$ (3.3)
By recalling the definition of $\beta_{N}$ and $h_{N}$ from (1.11), Denote
$\begin{split}&f_{m,k}^{(N)}(t_{1},\cdots,t_{m})\\ &:=\begin{cases}\displaystyle\sum_{\boldsymbol{q}\in\llbracket N\rrbracket^{m}}\sum_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}P_{\tau}(\boldsymbol{q},\boldsymbol{s})\prod_{i=1}^{m}\mathrm{I}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),\quad&\text{for}~{}1-H<\alpha<1,\\[10.0pt] \displaystyle\sum\limits_{\boldsymbol{q}\in\llbracket N\rrbracket^{m}}\sum\limits_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}N^{-k}P_{\tau}(\boldsymbol{q},\boldsymbol{s})\prod\limits_{i=1}^{m}\mathrm{I}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),\quad&\text{for}~{}\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \displaystyle\sum\limits_{\boldsymbol{q}\in\llbracket N\rrbracket^{m}}\sum\limits_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}N^{-k}l(N)^{m+k}P_{\tau}(\boldsymbol{q},\boldsymbol{s})\prod\limits_{i=1}^{m}\mathrm{I}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),\quad&\text{if}~{}\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}\end{split}$ (3.4)
where
$P_{\tau}(\boldsymbol{q},\boldsymbol{s}):={\mathbb{P}}(\{q_{1},\cdots,q_{m},s_{1},\cdots,s_{k}\}\subset\tau).$ (3.5)
Then, recalling the definition (2.11) for $\mathbf{I}_{m}^{(N)}(f)$, we have
$S_{m,k}^{(N)}=\hat{\beta}^{m}\hat{h}^{k}\mathbf{I}^{(N)}_{m}(f_{m,k}^{(N)}).$ (3.6)
3.1. Weak convergence
First, we prove the weak convergence of each chaos, i.e., the weak convergence of $\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})$, where $\mathbf{I}_{m}^{(N)}$ is given by (2.11) and $f_{m,k}^{(N)}$ by (3.4).
Assumption (A1).
Assume $H\in(\frac{1}{2},1)$ and $\alpha+H>1$.
Proposition 3.1.
Assume (A1). Then, we have
$\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\overset{(d)}{\longrightarrow}\mathbf{I}_{m}(\psi_{m,k}),\quad\text{as}~{}N\to\infty.$ (3.7)
Here
$\psi_{m,k}(t_{1},\cdots,t_{m}):=\begin{cases}\displaystyle\int_{[0,1]^{k}}\phi_{m+k}(t_{1},\cdots,t_{m+k}){\mathrm{d}}t_{m+1}\cdots{\mathrm{d}}t_{m+k},\quad&\text{if}~{}1-H<\alpha<1,\\ \displaystyle\Big{(}\frac{1}{\mathbb{E}[\tau_{1}]}\Big{)}^{m+k},&\text{if }~{}\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] 1\,,&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}$ (3.8)
where the integrand $\phi_{m+k}$ is the symmetric function given by (1.12) with $\alpha<1$.
Proof.
We shall focus on the case $11-H<{\alpha}<$ 1 and the proofs for the other two cases are similar.
Firstly, observe that $\psi_{m,k}(t_{1},\cdots,t_{m})$ belongs to the space $|\mathcal{H}|^{\otimes m}$ due to the following lemma.
Lemma 3.2.
There exists a constant $C=C_{\alpha,H}$ such that for all $m,k\in\mathbb{N}$,
$\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}^{2}\leq\frac{C^{m+k}[(m+k)!]^{2H}}{\Gamma\left((m+k)\frac{\alpha+H-1}{H}+1\right)^{2H}}.$
Proof of Lemma 3.2.
Recall (2.3) and apply Lemma A.4 and Jensen’s inequality,
$\begin{split}&\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|\psi_{m,k}\|_{L^{1/H}(\mathbb{R}^{m})}^{2}\leq C_{H}^{m}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}^{2}.\end{split}$
Then by permuting $t_{1},\cdots,t_{m+k}$ and by Lemma A.5, we obtain the desired bound. ∎
By Proposition 2.6, we have $\mathbf{I}_{m}^{(N)}(\psi_{m,k})\overset{(d)}{\longrightarrow}\mathbf{I}_{m}(\psi_{m,k})\text{, as}~{}N\rightarrow\infty.$ To prove (3.7), we denote
$Y_{m,k}^{(N)}=\mathbf{I}_{m}^{(N)}(\psi_{m,k})-\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})=\mathbf{I}_{m}^{(N)}(\psi_{m,k}-f_{m,k}^{(N)}),$
and then it suffices to show $Y_{m,k}^{(N)}\rightarrow 0$ in $L^{2}$ as $N\rightarrow\infty$. By Lemma 2.4 and noting the symmetry of $f_{m,k}^{(N)}$ and $\psi_{m,k}$, we get
$\mathbb{E}\Big{[}\Big{(}Y_{m,k}^{(N)}\Big{)}^{2}\Big{]}\leq m!C^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}.$
Thus, it remains to show
$\lim_{N\to\infty}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}=0.$
Define
$g_{m,k}^{(N)}(t_{1},\cdots,t_{m+k}):=\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}{\mathbb{P}}\left(\{n_{1},\cdots,n_{m+k}\}\subset\tau\right)\prod_{i=1}^{m+k}\mathbf{I}_{(\frac{n_{i}-1}{N},\frac{n_{i}}{N}]}(t_{i}).$ (3.9)
Recall (3.4) and we can get
$f_{m,k}^{(N)}(t_{1},\cdots,t_{m})=\int_{[0,1]^{k}}g_{m,k}^{(N)}(t_{1},\cdots,t_{m+k})\prod\limits_{i=1}^{k}{\mathrm{d}}t_{m+i}.$ (3.10)
By (1.4) and the renewal property, we have $\lim_{N\to\infty}g_{m,k}^{(N)}(t_{1},\cdots,t_{m+k})=\phi_{m+k}(t_{1},\cdots,t_{m+k})$. By Karamata’s representation theorem for slowly varying function, we have for any $\epsilon>0$, there exists a constant $A=A_{\epsilon}\in(0,\infty)$ such that
$\frac{1}{A}\left(\frac{n}{m}\right)^{-\epsilon}\leq\frac{L(n)}{L(m)}\leq A\left(\frac{n}{m}\right)^{\epsilon},~{}\text{ for all }n,m\in\mathbb{N}~{}\text{with}~{}m\leq n.$ (3.11)
Also by (1.4), for a larger constant $A^{\prime}\in(0,\infty)$, we have
$\frac{C_{\alpha}}{A^{\prime}(t-s)^{1-{\alpha}-\epsilon}}\leq\frac{L(N)}{N^{{\alpha}-1}}{\mathbb{P}}_{\tau}\left(N(t-s)\in\tau\right)\leq\frac{A^{\prime}C_{\alpha}}{(t-s)^{1-{\alpha}+\epsilon}},~{}\text{ for all}~{}t,s\in\Omega_{N}.$
Together with (1.6) and (1.7), we have that ($t_{i}$ is unnecessarily belonging to $\Omega_{N}$).
$g_{m,k}^{(N)}(t_{1},\cdots,t_{m+k})\leq C^{m+k}\left(\phi_{m+k}(t_{1},\cdots,t_{m+k})\right)^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}.$ (3.12)
By choosing $\epsilon$ sufficiently small such that $(\alpha-\epsilon)+H>1$, the term on the right-hand side above belongs to $L^{\frac{1}{H}}$, as has been shown in Lemma  3.2. Then, by the dominated convergence theorem, we have $\lim_{N\to\infty}\|g_{m,k}^{(N)}-\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}=0$.
Finally, Jensen’s inequality yields
$\int_{[0,1]^{m}}\left|f_{m,k}^{(N)}-\psi_{m,k}\right|^{\frac{1}{H}}{\mathrm{d}}t_{1}\cdots{\mathrm{d}}t_{m}\leq\int_{[0,1]^{m+k}}\left|g_{m,k}^{(N)}-\phi_{m+k}\right|^{\frac{1}{H}}{\mathrm{d}}t_{1}\cdots{\mathrm{d}}t_{m+k},$
and thus by Lemma A.4, $\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{L^{1/H}(\mathbb{R}^{m})}^{2}\to 0$, as $N\to\infty$. ∎
Now we are ready to prove our first main result.
Proof of Theorem 1.1.
We assume (A1) and $\alpha>\frac{1}{2}$, i.e., $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$. Note that
$\tilde{Z}_{\hat{\beta},\hat{h}}$ $=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod\limits_{i=1}^{k}\diamond\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right)$
$=\sum_{k=0}^{\infty}\sum_{p=0}^{k}\frac{\hat{\beta}^{p}\hat{h}^{k-p}}{k!}{\binom{k}{p}}\mathbf{I}_{p}\bigg{(}\int_{[0,1]^{k-p}}\phi_{k}(t_{1},\cdots,t_{k}){\mathrm{d}}t_{p+1}\cdots{\mathrm{d}}t_{k}\bigg{)}$
$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbf{I}_{p}\bigg{(}\int_{[0,1]^{q}}\phi_{k}(t_{1},\cdots,t_{k}){\mathrm{d}}t_{p+1}\cdots{\mathrm{d}}t_{k}\bigg{)}$
$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbf{I}_{p}(\psi_{p,q})=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}(\psi_{m,k}),$
where $\psi_{m,k}$ is given in (3.8).
Recall that $\tilde{Z}_{N}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}$ is given in (3.2) (see also (3.6)):
$\tilde{Z}_{N}:=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{m!k!}S_{m,k}^{(N)}=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)}).$
We define, recalling (3.4) and (3.8) for the definitions of $f_{m,k}^{(N)}$ and $\psi_{m,k}$ respectively,
$\tilde{Z}_{N}^{(R)}:=\sum_{k+m\leq R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\quad\text{and}\quad\tilde{Z}^{(R)}:=\sum_{k+m\leq R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}(\psi_{m,k}).$
By Proposition 3.1, we have that $\tilde{Z}_{N}^{(R)}\overset{(d)}{\longrightarrow}\tilde{Z}^{(R)}$ as $N\to\infty$. To complete the proof, we will apply Lemma A.3. Hence, it suffices to prove, as $R\rightarrow\infty$, $\tilde{Z}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}$ and $\tilde{Z}_{N}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{N}$ uniformly in $N$, which is done in the following two steps respectively.
Step (1): We have
$\Big{\|}\tilde{Z}^{(R)}-\tilde{Z}_{\hat{\beta},\hat{h}}\Big{\|}_{L^{2}}$ $=\Big{\|}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\Big{\|}_{L^{2}}\leq\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\Big{\|}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\Big{\|}_{L^{2}}$
$\leq C^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}\leq C_{H}^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}$
$\leq\sum_{k+m>R}(C_{\hat{\beta},\hat{h}}^{\alpha,H})^{m+k}\frac{[(m+k)!]^{H}}{\sqrt{m!}k!\Gamma\left((m+k)\frac{\alpha+H-1}{H}+1\right)^{H}},$
where the second inequality is due to the Riemann sum approximation and the third inequality follows from Lemma 3.2. By Stirling’s formula, each term in the sum on the right-hand side is bounded above by
$C^{m+k}\frac{(m+k)^{H/2}}{\sqrt{m^{1/2}k}\left((m+k)\frac{\alpha+H-1}{H}\right)^{H/2}}\frac{(m+k)^{(m+k)H}}{m^{m/2}k^{k}(m+k)^{(m+k)(\alpha+H-1)}}\leq C^{m+k}\frac{(m+k)^{(m+k)(1-\alpha)}}{m^{m/2}k^{k}}.$ (3.13)
Note that for $a,b,c>0,(a+b)^{c}\leq(2\max\{a,b\})^{c}\leq 2^{c}(a^{c}+b^{c})$. Hence, the right-hand side of (3.13) is bounded above by:
$C^{m+k}\Big{(}\frac{m^{(1-\alpha)m}m^{(1-\alpha)k}}{m^{m/2}k^{k}}+\frac{k^{(1-\alpha)k}k^{(1-\alpha)m}}{m^{m/2}k^{k}}\Big{)},$
which is summable in $m,k$ assuming $\alpha>\frac{1}{2}$ (this is where the additional condition $\alpha>\frac{1}{2}$ is used) by Lemma A.6. Hence, we have proved $\tilde{Z}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}$ as $R\rightarrow\infty$.
Step (2): To prove the uniform convergence $\tilde{Z}_{N}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{N}$ as $R\to\infty$, recalling (3.4) and (3.6),
$\begin{split}&\Big{\|}\tilde{Z}^{(R)}_{N}-\tilde{Z}_{N}\Big{\|}_{L^{2}}^{2}=\mathbb{E}\Big{[}\Big{(}\sum_{k+m>R}\frac{1}{m!k!}S_{m,k}^{(N)}\Big{)}^{2}\Big{]}\\ =&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}\mathbb{E}\left[\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\mathbf{I}_{m}^{(N)}(f_{m,k^{\prime}}^{(N)})\right]\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}C^{m}m!\|f_{m,k}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}\|f_{m,k^{\prime}}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}.\end{split}$ (3.14)
where the inequality follows from Cauchy-Schwartz inequality and Lemma 2.4. Recalling (3.9) and (3.10), by Jensen’s inequality and Lemma A.4, then (3.14) is bounded above by
$\begin{split}&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\|g_{m,k}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\|g_{m,k^{\prime}}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\Big{\|}\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\Big{\|}\phi_{m+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum_{k+m>R}C^{m+k}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\Big{\|}\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\sum_{k^{\prime}+m^{\prime}>R}C^{m^{\prime}+k^{\prime}}\frac{\hat{\beta}^{m^{\prime}}\hat{h}^{k^{\prime}}}{\sqrt{m^{\prime}!}k^{\prime}!}\Big{\|}\phi_{m^{\prime}+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m^{\prime}+k^{\prime}})},\end{split}$
where the second inequality follows from (3.12) and the term on the right-hand side goes to 0 as $R\to\infty$ by choosing $\epsilon$ sufficiently small such that $(\alpha-\epsilon)+H>1$ and ${\alpha}-\epsilon>\frac{1}{2}$ and the result in Step (1). ∎
Remark 3.3.
Note that $\alpha=\frac{1}{2}$ is indeed a criticality. For simplicity, assume $h_{N}\equiv 0$ and $L(\cdot)\equiv 1$ in (1.11). Denote by $\tau^{\prime}$ an independent copy of $\tau$. Then
$\begin{split}\mathbb{E}\Big{[}\big{(}\tilde{Z}_{N}\big{)}^{2}\Big{]}&=\mathbb{E}_{\tau,\tau^{\prime}}\Big{[}\exp\Big{\{}\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{\}}\Big{]}\\ &\geq\exp\Big{\{}\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\Big{\}}{\mathbb{P}}(\{1,\cdots,N\}\subset\tau\cap\tau^{\prime})\\ &\geq\exp\Big{(}C\hat{\beta}^{2}\frac{1}{N^{2(\alpha+H-1)}}\sum\limits_{n,m=1}^{N}|n-m|^{2H-2}\Big{)}({\mathbb{P}}(\tau_{1}=1))^{2N}\\ &\geq\exp(CN^{2(1-\alpha)})({\mathbb{P}}(\tau_{1}=1))^{2N},\end{split}$
which tends to infinity as $N\to\infty$ if $\alpha<\frac{1}{2}$. In this case, $\{\tilde{Z}_{N}\}$ is not $L^{2}$ uniformly bounded.
3.2. Some discussion on uniform integrability
The Weinrib-Halperin prediction suggests that our pinning model should be disorder relevant under (A1), i.e., $H\in(\frac{1}{2},1)$ and $\alpha+H>1$. This is partially confirmed by our results in the following sense: Proposition 3.1 yields the weak convergence of each chaos under (A1); while for the weak convergence of the partition functions, an extra condition $\alpha>\frac{1}{2}$ is imposed in order to guarantee the $L^{2}$-convergence of the Wiener chaos expansion (see Step (1) in the proof of Theorem 1.1 and Remark 3.3). This implies that when $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$, the family $\{\tilde{Z}_{N}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\}_{N\in\mathbb{N}}$ of the rescaled partition functions is bounded in $L^{2}$-norm and in particular, for the limiting distribution we have $\mathbb{E}[\tilde{Z}_{\hat{\beta},\hat{h}}]=1$. Inspired by the Weinrib-Halperin prediction, we conjecture that the family $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is uniformly integrable and converges weakly to an $L^{1}$-integrable limit under (A1).
In the rest of this section, we shall prove the uniform integrability of $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ and characterize its weak limit, under a stronger condition:
Assumption (A2).
Assume $H\in(\frac{1}{2},1)$ and $\alpha+2H>2$.
Condition (A2) implies condition (A1) since $H<1$. Note that (A2) is also the optimal condition that ensures the convergence of each chaos in the Stratonovich case (see Section 4).
Lemma 3.4.
Assuming (A2), we have
$\sup_{N}\beta_{N}^{2}\mathbb{E}\Big{[}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{]}<\infty,$ (3.15)
and
$\sup_{N}\mathbb{E}\Big{[}\exp\Big{\{}h_{N}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}\Big{\}}\Big{]}<\infty.$ (3.16)
Proof.
We prove the result for the case $\alpha<1$ and the other cases can be proven in a similarly way. For (3.15), we have
$\quad\beta_{N}^{2}\mathbb{E}\Big{[}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{]}$
$=\hat{\beta}^{2}\sum_{n,m=1}^{N}N^{2-2H}\gamma(n-m)\frac{L(N)^{2}}{N^{2{\alpha}-2}}{\mathbb{P}}_{\tau}(n\in\tau,m\in\tau)\frac{1}{N^{2}}$
$\leq C\int_{0}^{1}\int_{0}^{1}|s-t|^{2H-2}\phi_{2}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}(s,t){\mathrm{d}}s{\mathrm{d}}t<\infty$
for $\alpha+2H>2$, where $\phi_{2}$ is defined in (1.12) and the inequality follows from (1.6), (1.7) and (3.12). For (3.16), we have that for some $\alpha^{\prime}\in(0,\alpha)$
$\begin{split}\mathbb{E}\Big{[}\exp\Big{\{}\sum_{n=1}^{N}h_{N}\sigma_{n}\Big{\}}\Big{]}&=\sum\limits_{k=0}^{\infty}\frac{1}{k!}(h_{N})^{k}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{k}}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{k}\in\tau)\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{\Gamma(k\alpha^{\prime})}<\infty,\end{split}$ (3.17)
where we have used (3.12) and Lemma A.5. ∎
Proposition 3.5.
Assuming (A2), the family $\big{\{}\tilde{Z}_{N}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big{\}}_{N\in\mathbb{N}}$ of the rescaled partition functions is uniformly integrable.
Proof.
Recall that
$\tilde{Z}_{N}=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{\}}\exp\Big{\{}\diamond\Big{(}\beta_{N}\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big{)}\Big{\}}\Big{]},$
and we have
$\mathbb{E}_{\omega}\Big{[}\Big{(}\tilde{Z}_{N}\Big{)}^{2}\Big{]}=\mathbb{E}_{\tau,\tau^{\prime}}\Big{[}\exp\Big{\{}h_{N}\sum_{n=1}^{N}(1_{\{n\in\tau\}}+\mathbf{1}_{\{n\in\tau^{\prime}\}})\Big{\}}\exp\Big{\{}\beta_{N}^{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{\}}\Big{]}.$
Note that
$\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau,m\in\tau^{\prime}\}}=\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\mathbb{E}[\omega_{n}\omega_{m}]\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}$
$\leq$ $\beta_{N}^{2}\bigg{(}\mathbb{E}_{\omega}\Big{[}\Big{(}\sum_{n=1}^{N}\omega_{n}\mathbf{1}_{\{n\in\tau\}}\Big{)}^{2}\Big{]}\mathbb{E}_{\omega}\Big{[}\Big{(}\sum_{n=1}^{N}\omega_{n}\mathbf{1}_{\{n\in\tau^{\prime}\}}\Big{)}^{2}\Big{]}\bigg{)}^{\frac{1}{2}}$
$=$ $\beta_{N}^{2}\Big{(}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau^{\prime}\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{)}^{\frac{1}{2}}.$
We have that by $ab\leq\frac{1}{2}(a^{2}+b^{2})$,
$\mathbb{E}_{\omega}\Big{[}\Big{(}\tilde{Z}_{N}\Big{)}^{2}\Big{]}$ $\leq\mathbb{E}_{\tau,\tau^{\prime}}\bigg{[}\exp\Big{\{}h_{N}\sum_{n=1}^{N}1_{\{n\in\tau\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{\}}$
×$\quad\times\exp\Big{\{}h_{N}\sum_{n=1}^{N}1_{\{n\in\tau^{\prime}\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau^{\prime}\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{\}}\bigg{]}.$
Denote
$A_{l}^{N}:=\Big{\{}h_{N}\sum_{n=1}^{N}1_{\{n\in\tau\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\leq l\Big{\}}.$
Lemma 3.4 and Markov’s inequality together yields:
$\lim_{l\to\infty}\sup_{N}{\mathbb{P}}\left((A_{l}^{N})^{c}\right)=0.$ (3.18)
Define
$\tilde{Z}_{N}(A_{l}^{N}):=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\mathbf{1}_{A_{l}^{N}}\Big{\}}\exp\Big{\{}\diamond\Big{(}\beta_{N}\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\mathbf{1}_{A_{l}^{N}}\Big{)}\Big{\}}\Big{]}.$
Then, by Lemma A.7, we get
$\left|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})\right|\leq\mathbb{E}_{\tau}\left[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\right]+{\mathbb{P}}((A_{l}^{N})^{c}),$
where we denote
$\mathscr{E}_{N}:=\exp\Big{\{}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{\}}\exp\Big{\{}\diamond\Big{(}\beta_{N}\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big{)}\Big{\}}.$
Thus we have
$\mathbb{E}_{\omega}\left|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})\right|$ $\leq\mathbb{E}_{\omega}\mathbb{E}_{\tau}\left[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\right]+{\mathbb{P}}((A_{l}^{N})^{c})$ (3.19)
$=\mathbb{E}_{\tau}\Big{[}\mathbf{1}_{(A_{l}^{N})^{c}}\exp\Big{(}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{)}\Big{]}+{\mathbb{P}}((A_{l}^{N})^{c})$
$\leq{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}\mathbb{E}_{\tau}\Big{(}\Big{[}\exp\Big{(}2h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{)}\Big{]}\Big{)}^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c})$
$\leq C{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c}),$
where the constant $C$ in the last inequality is independent of $N$ by (3.16).
Finally, noting (3.18), (3.19), and the fact that $\mathbb{E}[\tilde{Z}_{N}(A_{l}^{N})^{2}]\leq e^{2l}$ for all $N$, we can apply Lemma A.8 to obtain the uniform integrability of $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$. ∎
A direct consequence of the uniform integrability is the following.
Proposition 3.6.
The sequence $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is tight. Assuming (A2), we have $\mathbb{E}[\mathcal{Z}]=1$ for any weak limit $\mathcal{Z}$.
Proof.
The tightness follows from $\tilde{Z}_{N}>0$ a.s. and $\mathbb{E}\big{[}\tilde{Z}_{N}\big{]}=1$, for all $N\in\mathbb{N}$. Assuming (A2), the family $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is uniformly integrable due to Proposition 3.5 and thus $\mathbb{E}[\mathcal{Z}]=1$. ∎
In the proof of Proposition 3.5, we apply Cauchy-Schwartz inequality to decouple $\tau$ and $\tau^{\prime}$, which requires the stronger Assumption (A2). We believe this is a technical issue and (A1) should be sufficient for the uniform integrability.
When $\alpha<\frac{1}{2}$, though $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is tight, we are not able to prove the weak convergence as in the proof of Theorem 1.1. This is because, when $\alpha<\frac{1}{2}$, $\tilde{Z}_{N}$ is not uniformly bounded in $L^{2}$ (see Remark 3.3) and thus $\mathbb{E}[\mathcal{Z}^{2}]=\infty$ for any weak limit $\mathcal{Z}$. Therefore, we cannot employ the theory of Wiener chaos, which applies only to $L^{2}$-integrable random variables, to obtain a weak limit.
4. The Stratonovich case
In this section, we shall study the weak convergence for the original partition function defined in (1.8) and prove Theorem 1.2. Our approach is similar to that in Section 3.1 but with more involved details. First, we show that the limit (1.14) in Theorem 1.2 is well-defined in $L^{2}$. To do that, we need good enough upper bounds for the norms of each chaos and the trace defined in (2.7). Then we show that each term in the Taylor expansion of the partition function converges weakly to a corresponding chaos. Theorem 1.2 then follows from a classical truncation procedure.
4.1. $L^{2}$-convergence of the continuum partition function.
Define the following continuum partition function
$Z_{\hat{\beta},\hat{h}}$ $=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod_{i=1}^{k}\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right)$ (4.1)
[…]
$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbb{I}_{p}(\psi_{p,q})=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbb{I}_{m}(\psi_{m,k}).$
Recall the functions $\psi$ in (3.8), $\phi$ in (1.12), and the trace defined in (2.7).
Lemma 4.1.
Assume (A2). Then we have $\mathrm{Tr}^{j}{\psi}_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $j=0,1,...,\lfloor\frac{m}{2}\rfloor$. Moreover, we have
$\|\mathrm{Tr}^{j}{\psi}_{m,k}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}\leq\|\mathrm{Tr}^{j}{\phi}_{m+k}\|^{2}_{|\mathcal{H}|^{\otimes(m+k-2j)}}.$ (4.2)
Proof.
We only prove the case $\alpha\in(0,1)$. When $k=0$, we have $\psi_{m,0}=\phi_{m}$ and thus
$\|\mathrm{Tr}^{j}{\phi}_{m}$ ||($\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg{(}\int_{[0,1]^{2j}}\phi_{m}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
×$\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\phi_{m}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
First, we split $[0,1]^{2m}$ into $(2m)!$ regions such that in each region, the numbers $\{t_{i},t_{i}^{\prime},s_{j},s_{j}^{\prime}:i=1,\dots,m-2j,j=1,\dots,2j\}$ are listed in an increasing order. To get an upper bound, we replace, for instance, each $|t_{i}-t_{i}^{\prime}|^{2H-2}$ by $|t_{i}^{*}-t_{i}^{\star}|^{2H-2}$ where $t^{*}_{i}:=\max\{t_{i},t_{i}^{\prime}\}$ and $t_{i}^{\star}$ is the largest number that is smaller than $t^{*}_{i}$ (it could be some $t_{j}$, $t^{\prime}_{j}$, $s_{j}$ or $s^{\prime}_{j}$), and we perform the same operation for all the other $|s_{2i}-s_{2i-1}|^{2H-2}$, $|s^{\prime}_{2i}-s^{\prime}_{2i-1}|^{2H-2}$ and the factors in $\phi_{m}$ (see Figures 1 and 2 for illustration). Then on each region, the above integral is smaller than an integral with $2m$ variables in Lemma A.5 with $m$ exponents equal to ${\alpha}-1$ and $m$ exponents equal to $32H+{\alpha}-$ 3. Noting that $12H+\alpha-3>-$ 1 under (A2), we get by Lemma A.5
$\|\mathrm{Tr}^{j}{\phi}_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}\leq(2m)!\frac{\Gamma({\alpha})^{m}\Gamma({\alpha}+2H-2)^{m}}{\Gamma(m(2H+2\alpha-2)+1)}<\infty.$ (4.3)
[…]
Figure 1. An example for $m=6$ and $j=2$ before modifications, the red segments represent the renewal kernels (with exponent $\alpha-1$) and the blue segments represent the kernels of the fractional Brownian motion (with exponent $22H-$ 2).
[…]
Figure 2. After modifications, the red segments represent the renewal kernels and the blue segments represent the kernels of the fractional Brownian motion.
When $k>0$, we have
$\quad\|\mathrm{Tr}^{j}{\psi}_{m,k}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}$
$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg{(}\int_{[0,1]^{2j}}\psi_{m,k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
×$\quad\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\psi_{m,k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}$
×$\qquad\times\bigg{(}\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j},x_{1},\cdots,x_{k})\mathrm{d}\boldsymbol{x}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
×$\qquad\times\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime},x_{1}^{\prime},\cdots,x_{k}^{\prime})\mathrm{d}\boldsymbol{x^{\prime}}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
$\leq\int_{[0,1]^{2(m+k-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\prod_{p=1}^{k}|x_{p}-x_{p}^{\prime}|^{2H-2}\mathrm{d}x_{p}\mathrm{d}x_{p}^{\prime}$
×$\quad\times\bigg{(}\int_{[0,1]^{2j}}\phi_{m+k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j},x_{1},\cdots,x_{k})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
×$\quad\times\int_{[0,1]^{2j}}\phi_{m+k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime},x_{1}^{\prime},\cdots,x_{k}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
$=\|\mathrm{Tr}^{j}{\phi}_{m+k}\|^{2}_{|\mathcal{H}|^{\otimes(m+k-2j)}}.$
This proves (4.2) and hence $\mathrm{Tr}^{j}{\psi}_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ from the first step. ∎
Remark 4.2.
Assuming (A2), Lemma 4.1 yields ${\vvvert\psi_{m,k}\vvvert_{m}<\infty}$ (see (2.9)) and hence $\mathbb{I}_{m}(\psi_{m,k})$ is well defined due to Hu-Meyer’s formula (2.8). On the other hand, $\alpha+2H>2$ is the optimal condition for ${\vvvert\psi_{m,k}\vvvert_{m}<\infty}$. Indeed, for $k=0,m=2$ and $j=1$, we have
[…]
which is finite if and only if ${\alpha}+2H>2$.
We have shown that $\mathbb{I}_{m}(\psi_{m,k})$ is well defined under (A2). In order to make sense of $Z_{\hat{\beta},\hat{h}}$, we impose the following condition to obtain the $L^{2}$-convergence for the series on the right-hand side of (4.1).
Assumption (A3).
Assume $H\in(\frac{1}{2},1)$ and $\alpha+H>\frac{3}{2}$.
Note that the Assumption (A3) implies the Assumption (A2) due to $H>\frac{1}{2}$.
Proposition 4.3.
[…]
Proof.
Recalling
$Z_{\hat{\beta},\hat{h}}=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbb{I}_{m}(\psi_{m,k}),$
we have
$\|Z_{\hat{\beta},\hat{h}}\|_{2}$ $\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\|\mathbb{I}_{m}(\psi_{m,k})\|_{2}$
$\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\sqrt{2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}}$
$\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{C^{m+k}}{\sqrt{m!}k!}\sqrt{\frac{[2(m+k)]!}{\Gamma((m+k)(2H+2\alpha-2)+1)}}\,,$
where the second inequality follows from Lemma 4.4 below and the last inequality is due to (4.2) and (4.3) below.
Note that $\alpha+H>\frac{3}{2}$ implies $\zeta=4-2H-2\alpha<1$. By Stirling’s formula, we get
$\frac{C^{m+k}}{\sqrt{m!}k!}\sqrt{\frac{[2(m+k)]!}{\Gamma((m+k)(2H+2\alpha-2))}}\leq\sqrt{\frac{C^{m+k}(m+k)^{\zeta(m+k)}}{m^{m}k^{2k}}}\leq\sqrt{C^{m+k}\left(\frac{m^{\zeta(m+k)}}{m^{m}k^{2k}}+\frac{k^{\zeta(m+k)}}{m^{m}k^{2k}}\right)}.$
This together with Lemma A.6 yields $\|Z_{\hat{\beta},\hat{h}}\|_{2}<\infty$. ∎
Lemma 4.4.
Recall $\phi$ in (1.12) and $\psi$ in (3.8). We have
$\mathbb{E}[\mathbb{I}_{m}(\phi_{m})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\phi}_{2m},$
and
$\mathbb{E}[\mathbb{I}_{m}(\psi_{m,k})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}.$
Proof.
For the first inequality, Hu-Meyer’s formula (2.8) implies
$\mathbb{E}[|\mathbb{I}_{m}(\phi_{m})|^{2}]$ $=\sum_{k=0}^{[\frac{m}{2}]}\frac{1}{(m-2k)!}\left(\frac{m!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.4)
$=\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}\binom{m}{2k}\binom{2k}{k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}\leq\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}2^{m+2k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$
$=2^{m}m!\sum_{k=0}^{[\frac{m}{2}]}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}},$
where the inequality follows from the fact $\binom{m}{n}\leq 2^{m}$. For each term on the right-hand side, we have
$\|\mathrm{Tr}^{k}{\phi}_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.5)
$=$ $\int_{[0,1]^{2m}}\prod_{i=1}^{m-2k}|t_{i}-t_{i}^{\prime}|^{2H-2}\prod_{j=1}^{k}|s_{2j}-s_{2j-1}|^{2H-2}|s_{2j}^{\prime}-s_{2j-1}^{\prime}|^{2H-2}$
×$\times\quad\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}){\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{s^{\prime}}{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{t^{\prime}}$
$=$ $\mathrm{Tr}^{m}{\Phi}_{m}\leq\mathrm{Tr}^{m}{\phi}_{2m},$
where
$\Phi_{m}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime})$
::=:= $\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime})$
$\leq$ ${\phi}_{2m}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime}).$
Substituting (4.5) into (4.4) yields the desired inequality.
Similar to the above, noting that
$\quad\psi_{2m,2k}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime})$
$\geq\psi_{m,k}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\psi_{m,k}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}),$
we can prove the second inequality in the same way. ∎
4.2. Weak convergence of the rescaled partition functions.
First, note that Taylor expansion of the partition function $Z_{N}:=Z_{N,\beta_{N}}^{\omega,h_{N}}$ given in (1.8) yields
$Z_{N}$ $=\mathbb{E}_{\tau}\bigg{[}\Big{(}\sum_{k=0}^{\infty}\frac{1}{k!}\Big{(}\sum_{n=1}^{N}h_{N}\sigma_{n}\Big{)}^{k}\Big{)}\Big{(}\sum_{m=0}^{\infty}\frac{1}{m!}\Big{(}\sum_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{m}\Big{)}\bigg{]}$ (4.6)
$=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m},\cdots,n_{m+k}\in\tau)\prod_{i=1}^{m}\omega_{n_{i}},$
$=:\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\mathbb{S}_{m,k}^{(N)}.$
We then prove the weak convergence of each chaos.
Proposition 4.5.
Assume (A2). Then, for each $m,k\in\mathbb{N}_{0}$,
$\mathbb{S}_{m,k}^{(N)}\overset{(d)}{\longrightarrow}\hat{\beta}^{m}\hat{h}^{k}\mathbb{I}_{m}(\psi_{m,k}),\quad\text{as}~{}N\rightarrow\infty,$ (4.7)
where $\psi_{m,k}$ is defined in (3.8) and $\mathbb{I}_{m}(\psi_{m,k})$ is an $m$-th multiple Stratonovich integral. Moreover, for any $p\in\mathbb{N}$ and $l_{1},\cdots,l_{p},k_{1},\cdots,k_{p}\in\mathbb{N}_{0}$, we have the joint convergence in distribution
$\left(\mathbb{S}_{l_{1},k_{1}}^{(N)},\cdots,\mathbb{S}_{l_{p},k_{p}}^{(N)})\right)\overset{(d)}{\longrightarrow}\left(\hat{\beta}^{l_{1}}\hat{h}^{k_{1}}\mathbb{I}_{l_{1}}(\psi_{l_{1},k_{1}}),\cdots,\hat{\beta}^{l_{p}}\hat{h}^{k_{p}}\mathbb{I}_{l_{p}}(\psi_{l_{p},k_{p}})\right),\quad\text{as}~{}N\rightarrow\infty.$ (4.8)
Proof.
[…]
$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}]$
$=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}].$
For each $j=0,1,\cdots,\lfloor\frac{m}{2}\rfloor$, all the terms in the summation $\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket,|B|=m-2j\end{subarray}}$ are equal by symmetry. Fix $j\in\{0,1,...,\lfloor\frac{m}{2}\rfloor\}$ and note that there are in total $\binom{m}{m-2j}$ subsets of $\llbracket m\rrbracket$ with cardinality $m-2j$. Without loss of generality, we may assume $B=\{2j+1,...,m\}$. Then, we have
$\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}]=\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket 2j\rrbracket}}]=\sum_{V}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{n_{l_{1}}}\omega_{n_{l_{2}}}],$
where the sum $\sum_{V}$ is taken over all pair partitions $V$ of $\llbracket 2j\rrbracket$. By symmetry, the summations
$\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{n_{l_{1}}}\omega_{n_{l_{2}}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)$
equal to each other for all pair partitions $V$. Note that there are in total $(2j-1)!!$! pair partitions of $\llbracket 2j\rrbracket$, and thus
$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\binom{m}{m-2j}(2j-1)!!\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)$ (4.9)
$=\sum_{j=0}^{[\frac{m}{2}]}\frac{m!}{j!(m-2j)!2^{j}}\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau).$
Compare (4.9) with Hu-Meyer’s formula (2.8) and note that $\mathrm{Tr}^{j}\psi_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $k=0,1,...,\lfloor\frac{m}{2}\rfloor$ by Lemma 4.1. From now on, without loss of generality, we assume that $m$ is odd and the case of even $m$ can be treated in the same manner. By Proposition 2.6, we have
$\left(\mathbf{I}_{1}^{(N)}(\mathrm{Tr}^{\frac{m-1}{2}}\psi_{m,k}),\cdots,\mathbf{I}_{m}^{(N)}(\psi_{m,k})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{1}(\mathrm{Tr}^{\frac{m-1}{2}}\psi_{m,k}),\cdots,\mathbf{I}_{m}(\psi_{m,k})\right),\quad\text{as}~{}N\rightarrow\infty.$ (4.10)
Thus, in order to prove (4.7), it suffices to show that for all $j=0,1,\cdots,\frac{m-1}{2}$,
$Y_{N}:$ $=\frac{L(N)^{m+k}}{N^{m(\alpha-1+H)+\alpha k}}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)-\mathbf{I}_{m-2j}^{(N)}(\mathrm{Tr}^{j}\psi_{m,k})$ (4.11)
$=\mathbf{I}_{m-2j}^{(N)}(f_{\mathrm{Tr}^{j},m,k}^{(N)})-\mathbf{I}_{m-2j}^{(N)}(\mathrm{Tr}^{j}\psi_{m,k})\overset{L^{2}}{\longrightarrow}0\quad\text{as}~{}N\to\infty,$
with
$f_{\mathrm{Tr}^{j},m,k}^{(N)}=f_{\mathrm{Tr}^{j},m,k}^{(N)}(t_{1},\cdots,t_{m-2j})$ (4.12)
$:=N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}\sum_{\boldsymbol{u}\in\llbracket N\rrbracket^{2j}}\sum_{\boldsymbol{q}\in\llbracket N\rrbracket^{m-2j}}\sum_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})\prod_{i=1}^{m-2j}\mathbf{1}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),$
where we recall that $\mathbf{I}_{m-2j}^{(N)}$ is defined in (2.11), $P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})$ is defined in (3.5), and note that
$\frac{L(N)^{m+k}}{N^{m(\alpha-1+H)+\alpha k}}=N^{-(m-2j)H}N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}.$
For $j=0$, we have $Y_{N}=\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)}-\psi_{m,k})$, which converges to 00 in $L^{2}$ as $N\to\infty$ as has been proved in the proof of Proposition 3.1. For $j=1,...,\frac{m-1}{2}$, by Lemma 2.4 and Lemma A.4, we get
$\mathbb{E}[Y_{N}^{2}]$ $\lesssim(m-2j)!\|f_{\mathrm{Tr}^{j},m,k}^{(N)}-\mathrm{Tr}^{j}\psi_{m,k}\|_{|\mathcal{H}|^{\otimes(m-2j)}}^{2}$ (4.13)
$\lesssim(m-2j)!\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t^{\prime}_{i}|^{2H-2}|\boldsymbol{D}_{N}\boldsymbol{D}^{\prime}_{N}|{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{t}^{\prime}\lesssim(m-2j)!\|\boldsymbol{D}_{N}\|_{\frac{1}{H}}^{2}$
where
$\boldsymbol{D}_{N}=f_{\mathrm{Tr}^{j},m,k}^{(N)}-\mathrm{Tr}^{j}\psi_{m,k}\quad\text{and}\quad{\boldsymbol{D}}^{\prime}_{N}=\boldsymbol{D}_{N}(\boldsymbol{t}^{\prime}).$ (4.14)
Then, in order to prove (4.11), it suffices to show
$f_{\mathrm{Tr}^{j},m,k}^{(N)}\xrightarrow{L^{1/H}}\mathrm{Tr}^{j}\psi_{m,k}\quad\text{ as }N\to\infty.$ (4.15)
We denote
$g_{\mathrm{Tr}^{j},m,k}^{(N)}(t_{1},\cdots,t_{m+k})$
$:=\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\left(\prod_{l=1}^{j}\gamma_{N}\left(\frac{n_{2l-1}}{N}-\frac{n_{2l}}{N}\right){\mathbb{P}}_{\tau}\left(n_{1},\cdots,n_{m+k}\in\tau\right)\prod_{i=1}^{m+k}\mathbf{I}_{(\frac{n_{i}-1}{N},\frac{n_{i}}{N}]}(t_{i})\right),$
where $\gamma_{N}(t)$ is defined in (2.13), and denote
$\quad\quad\phi_{\mathrm{Tr}^{j},m+k}(t_{1},\cdots,t_{m+k}):=\prod_{l=1}^{j}|t_{2j-1}-t_{2j}|^{2H-2}\phi_{m+k}(t_{1},\cdots,t_{m+k}).$
Then we have
$f_{\mathrm{Tr}^{j},m,k}^{(N)}(t_{1},\cdots,t_{m-2j})$ $=\int_{[0,1]^{2j+k}}g_{\mathrm{Tr}^{j},m,k}^{(N)}(t_{1},\cdots,t_{m+k}){\mathrm{d}}t_{1}\cdots{\mathrm{d}}t_{2j}{\mathrm{d}}t_{m+1}\cdots{\mathrm{d}}t_{m+k},$
$\mathrm{Tr}^{j}\psi_{m,k}(t_{1},\cdots,t_{m-2j})$ $=\int_{[0,1]^{2j+k}}\phi_{\mathrm{Tr}^{j},m+k}(t_{1},\cdots,t_{m+k}){\mathrm{d}}t_{1}\cdots{\mathrm{d}}t_{2j}{\mathrm{d}}t_{m+1}\cdots{\mathrm{d}}t_{m+k}.$
To prove (4.15), by Jensen’s inequality, it suffices to show $g_{\mathrm{Tr}^{j},m,k}^{(N)}\xrightarrow{L^{1/H}}\phi_{\mathrm{Tr}^{j},m+k}$, as $N\to\infty$. First, by (1.4) and (1.7), we have $g_{\mathrm{Tr}^{j},m,k}^{(N)}\to\phi_{\mathrm{Tr}^{j},m+k}$ pointwisely as $N\to\infty$. By Lemma  4.1, $\phi_{\mathrm{Tr}^{j},m+k}\in L^{1/H}$. Hence, it suffices to exhibit an $L^{1/H}$ domination of $\big{\{}g_{\mathrm{Tr}^{j},m,k}^{(N)}\big{\}}_{N\geq 1}$. By (1.4) and (3.11), and together with (1.6) and (1.7), we have that for some $\alpha^{\prime}\in(0,\alpha)$,
$g_{\mathrm{Tr}^{j},m,k}^{(N)}(t_{1},\cdots,t_{m+k})\leq C^{m+k}\prod_{l=1}^{j}|t_{2l-1}-t_{2l}|^{2H-2}\big{(}\phi_{m+k}(t_{1},\cdots,t_{m+k})\big{)}^{\frac{1-{\alpha}^{\prime}}{1-{\alpha}}}.$ (4.16)
By a similar argument used in the proof of Lemma 4.1, we can show that the right-hand side of (4.16) belongs to $L^{1/H}$. This is provides the desired $L^{1/H}$ domination.
Finally, the proof is completed by the continuous mapping theorem and (4.10). ∎
4.3. $L^{1}$-bound of the rescaled partition functions
In this subsection, we aim to prove the uniform $L^{1}$-boundedness, under Assumption (A3), for the partition function $Z_{N}$. Recall its Taylor expansion (4.6) and by Cauchy-Schwartz inequality, we have that
$\begin{split}\mathbb{E}[Z_{N}]\leq\sum\limits_{k=0}^{\infty}\sum\limits_{m=0}^{\infty}\frac{1}{k!}\frac{1}{m!}\mathbb{E}_{\tau}\Big{[}\Big{(}\sum\limits_{n=1}^{N}h_{N}\sigma_{n}\Big{)}^{2k}\Big{]}^{\frac{1}{2}}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}.\end{split}$ (4.17)
Using the same argument as in (3.17), we get that for some $\alpha^{\prime}\in(0,\alpha)$,
$\sum\limits_{k=0}^{\infty}\frac{1}{k!}\mathbb{E}_{\tau}\Big{[}\Big{(}\sum\limits_{k=0}^{\infty}h_{N}\sigma_{N}\Big{)}^{2k}\Big{]}^{\frac{1}{2}}\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k!}\sqrt{\frac{(2k)!}{\Gamma(2k\alpha^{\prime})}}\lesssim\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k^{k\alpha^{\prime}}}<\infty.$ (4.18)
For each term in the summation over $m$ in (4.17), we have that
$\begin{split}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}=\Big{(}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{2m}}\beta_{N}^{2m}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{2m}\in\tau)\mathbb{E}_{\omega}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]\Big{)}^{\frac{1}{2}}.\end{split}$ (4.19)
Similar to the proof of Proposition 4.5, $\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]=\sum_{V}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{l_{1}}\omega_{l_{2}}]$, where $V$ is all pair partitions of $\llbracket 2m\rrbracket$. Note that $|V|=(2m-1)!!$!. By (1.4) and (3.11), a Riemann sum approximation and the kernel modification procedure in the proof of Lemma 4.1, (4.19) is bounded above by
$\sqrt{\frac{C^{2m}(2m)!(2m-1)!!}{\Gamma(m(2\alpha^{\prime}+2H-2))}}\lesssim C^{m}m^{m}\sqrt{\frac{m}{m^{m(2\alpha^{\prime}-2H-3)}}}$ (4.20)
for some $\alpha^{\prime}<\alpha$ and $\alpha^{\prime}+H>\frac{3}{2}$. Hence, we have that
$\sum\limits_{m=0}^{\infty}\frac{1}{m!}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}\leq\sum\limits_{m=0}^{\infty}\frac{C^{m}}{m^{m(\alpha^{\prime}+H-\frac{3}{2})}}<\infty.$ (4.21)
Combining (4.17), (4.18) and (4.21) we conclude the $L^{1}$-boundedness of $Z_{N}$ and we can also deduce from the proof that
$\lim_{M\to\infty}\sup_{N\in\mathbb{N}}\sum_{l=M}^{\infty}\sum_{m+k=l}\frac{1}{k!m!}\mathbb{E}\Big{[}\Big{|}h_{N}\sum_{n=1}^{N}\sigma_{n}\Big{|}^{k}\Big{|}\beta_{N}\sum_{n=1}^{N}\omega_{n}\sigma_{n}\Big{|}^{m}\Big{]}=0.$ (4.22)
Now we are ready to prove our second main result.
Proof of Theorem 1.2.
By Proposition 4.5, we have that for all $R\in\mathbb{N}$,
$Z^{(R)}_{N}=\sum_{l=0}^{R}\sum_{m+k=l}\frac{1}{k!m!}\mathbb{S}_{m,k}^{(N)}\overset{(d)}{\longrightarrow}Z^{(R)}:=\sum_{l=0}^{R}\sum_{m+k=l}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbb{I}_{m}(\psi_{m,k}),\quad\text{as}~{}N\to\infty.$
By Proposition 4.3, $Z^{(R)}\overset{L^{2}}{\longrightarrow}Z$ as $N\to\infty$. By Lemma A.3, we only need to show that $Z^{(R)}_{N}$ converges to $Z_{N}$ (see (4.6)) in probability uniformly in $N$ as $R\to\infty$. This follows from (4.22). ∎
Appendix A Some miscellaneous results
Lemma A.1.
Let $(\omega_{n})_{n\in\mathbb{N}}$ be Gaussian. Then
$\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}=\sum_{j=0}^{[\frac{m}{2}]}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}],$
where $\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}:=\prod_{i=1}^{m}\omega_{n_{i}}$ is an ordinary product and $\omega_{\boldsymbol{n}_{B}}^{\diamond}:=\prod_{n_{i}\in B}\diamond\omega_{n_{i}}$ is a Wick product.
Lemma A.2.
Let $(\omega_{n})_{n\in\mathbb{N}}$ be Gaussian. Then, for $\boldsymbol{n}=(n_{1},\dots,n_{p})$ and $\boldsymbol{m}=(m_{1},\dots,m_{q})$,
$\mathbb{E}[\omega_{\boldsymbol{n}}^{\diamond}\omega_{\boldsymbol{m}}^{\diamond}]=\begin{cases}\displaystyle 0&\text{ if }p\neq q,\\[10.0pt] \displaystyle\sum_{\pi}\prod_{\{i,j\}\in\pi}\mathbb{E}[\omega_{i}\omega_{j}],&\text{ if }p=q,\end{cases}$
where $\pi$ is a pairwise partition of the set $\{n_{1},\dots,n_{p},m_{1},\dots,m_{p}\}$ such that for each $\{i,j\}\in\pi$, $i\in\{n_{1},\dots,n_{p}\}$ and $j\in\{m_{1},\dots,m_{p}\}$.
Lemma A.3.
Consider random variables $Y_{n}^{(N)},Y^{(N)},Y_{n}$ and $Y$ such that $Y^{(N)}_{n}$ converges weakly to $Y_{n}$ as $N\to\infty$, $Y_{n}$ converges weakly to $Y$ as $n\to\infty$, and $Y_{n}^{(N)}$ converges in probability to $Y^{(N)}$ uniformly in $N$ as $n\to\infty$, then we have $Y^{(N)}$ converges weakly to $Y$ as $N\to\infty$.
The following Hardy-Littlewood-Sobolev inequality was proved in [14].
Lemma A.4.
For $H\in(\frac{1}{2},1]$, the following inequality holds:
$\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(\boldsymbol{t})f(\boldsymbol{s})\prod_{i=1}^{m}|t_{i}-s_{i}|^{2H-2}{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{s}\leq C_{H}^{m}\|f\|_{L^{1/H}(\mathbb{R}^{m})}^{2},$
where $C_{H}$ is a constant depending only on $H$.
The following result can be obtained by direct computation.
Lemma A.5.
Suppose ${\alpha}_{i}>-1$ for $i=1,\cdots,m$. Denoting ${\alpha}:=\sum_{i=1}^{m}{\alpha}_{i}$, then
$\int_{0=r_{0}<r_{1}<r_{2}<\cdots<r_{m}<t}\prod_{i=1}^{m}(r_{i}-r_{i-1})^{{\alpha}_{i}}{\mathrm{d}}\boldsymbol{r}=\frac{\prod_{i=1}^{m}\Gamma(1+{\alpha}_{i})}{\Gamma(m+{\alpha}+1)}t^{m+{\alpha}},$
where $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}{\mathrm{d}}t$ is the Gamma function.
Lemma A.6.
Let $a,b$ and $c$ be positive constants. Assuming $c>a$, then we have for all $A>0$,
$\sum_{k+m=1}^{\infty}A^{m+k}\frac{m^{ak}}{m^{bm}k^{ck}}<\infty.$
Proof.
For some large $k_{0}$ and $m_{0}$ which will be determined later, we write
$\sum_{k+m=1}^{\infty}A^{m+k}\frac{m^{ak}}{m^{bm}k^{ck}}\leq\Big{(}\sum_{k<k_{0}}\sum_{m\geq 1}+\sum_{m<m_{0}}\sum_{k\geq 1}+\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}\Big{)}A^{m+k}\frac{m^{ak}}{m^{bm}k^{ck}}.$
It is not hard to see that the first two sums are finite. For the third term, note that $(bm)^{ck}(ck)^{bm}\leq(bm)^{bm}(ck)^{ck}$. Let $d=c-a$ and we have that
$\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}A^{m+k}\frac{m^{ak}}{m^{bm}k^{ck}}\leq\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}A^{m+k}\frac{m^{ak}}{m^{ck}k^{bm}}\Big{(}\frac{b}{c}\Big{)}^{bm}\Big{(}\frac{c}{b}\Big{)}^{ck}$
$=$ $\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}\frac{1}{\left(C_{A}m\right)^{dk}\left(C^{\prime}_{A}k\right)^{bm}}\leq\sum_{m\geq m_{0}}\frac{1}{\left(C^{\prime}_{A}k_{0}\right)^{bm}}\sum_{k\geq k_{0}}\frac{1}{\left(C_{A}m_{0}\right)^{dk}}<\infty,$
by choosing $C_{A}^{\prime}k_{0}>1$ and $C_{A}m_{0}>1$. The proof is now completed. ∎
Lemma A.7.
Let $X$ be an exponentially integrable random variable on $(\Omega,\mathcal{F},{\mathbb{P}})$ and $A\in\mathcal{F}$. Then we have
$\mathbb{E}\left[e^{X}-e^{X\mathbf{1}_{A}}\right]={\mathbb{P}}(A^{c})\left(\mathbb{E}\left[e^{X}\Big{|}A^{c}\right]-1\right)=\mathbb{E}\left[e^{X}\mathbf{1}_{A^{c}}\right]-{\mathbb{P}}(A^{c}).$
Proof.
In general, we have
$\mathbb{E}\left[f(X)-f(X\mathbf{1}_{A})\right]$ $={\mathbb{P}}(A)\mathbb{E}[f(X)-f(X\mathbf{1}_{A})|A]+{\mathbb{P}}(A^{c})\mathbb{E}[f(X)-f(X\mathbf{1}_{A})|A^{c}]$
$={\mathbb{P}}(A)\cdot 0+{\mathbb{P}}(A^{c})\mathbb{E}[f(X)-f(0)|A^{c}]$
$=\mathbb{E}[f(X)\mathbf{1}_{A^{c}}]-f(0){\mathbb{P}}(A^{c}),$
which yields the desired result by choosing $f(x)=e^{x}$. ∎
Lemma A.8.
Consider a family $\{Y_{n},n\in\mathbb{N}\}$ of non-negative and integrable random variables such that $\sup_{n}\mathbb{E}|Y_{n}-Y_{n}^{\varepsilon}|<\varepsilon$, where $\{Y_{n}^{\varepsilon},n\in\mathbb{N},\varepsilon\in(0,1)\}$ satisfies $\sup_{n}\mathbb{E}[|Y_{n}^{\varepsilon}|^{2}]=M_{\varepsilon}<\infty$. Then, $\{Y_{n},n\in\mathbb{N}\}$ is uniformly integrable.
Proof.
By triangular inequality and Cauchy-Schwartz inequality, we get
$\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}|$ $\leq\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}-Y^{\varepsilon}_{n}\mathbf{1}_{[Y_{n}>k]}|+\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y_{n}>k]}-Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|+\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|$
$\leq\mathbb{E}|Y_{n}-Y_{n}^{\varepsilon}|+\big{(}\mathbb{E}|Y_{n}^{\varepsilon}|^{2}\big{)}^{\frac{1}{2}}\big{[}{\mathbb{P}}(Y_{n}>k)^{\frac{1}{2}}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{\frac{1}{2}}\big{]}+\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|$
$\leq\varepsilon+\sqrt{M_{\varepsilon}}\big{[}{\mathbb{P}}(Y_{n}>k)^{\frac{1}{2}}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{\frac{1}{2}}\big{]}+\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|.$
For any fixed $\varepsilon\in(0,1)$, we may find $K=K(\varepsilon)$ such that for all $k>K$ and $n\in\mathbb{N}$,
${\mathbb{P}}(Y_{n}>k)^{\frac{1}{2}}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{\frac{1}{2}}<\frac{\varepsilon}{\sqrt{M_{\varepsilon}}}\quad\text{and}\quad\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|<\varepsilon$
by Markov’s inequality and Cauchy-Schwartz inequality. Therefore, for all $k>K$ and $n\in\mathbb{N}$, $\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}|<3\varepsilon$. This proves the uniform integrability of $\{Y_{n},n\in\mathbb{N}\}$. ∎
Acknowledgements. J. Song is partially supported by NSFC Tianyuan Key Program Project (No. 12226001), National Natural Science Foundation of China (No. 12471142), and the Fundamental Research Funds for the Central Universities. R. Wei is supported by Xi’an Jiaotong-Liverpool University Research Development Fund (No. RDF-23-01-024) and NSFC (No. 12401170). The authors thank Quentin Berger, Rongfeng Sun and Jinjiong Yu for reading the first version of this paper and providing invaluable suggestions.

### v2

@@
 For the model that is disorder relevant, we study its intermediate disorder regime and show that the rescaled partition functions converge to a non-trivial continuum limit in the Skorohod setting and in the Stratonovich setting, respectively.
-Our results partially confirm the prediction of Weinrib and Halperin [21] for disorder relevance/irrelevance.
+Our results partially confirm the Weinrib–Halperin prediction for disorder relevance/irrelevance.
 1. Introduction
 In this paper, we study a pinning model in Gaussian random environments (also called the disorder), of which the correlation has a power-law decay.
-We shall show that, under proper scalings, the rescaled partition functions converge weakly to a continuum limit which can be represented by a chaos expansion.
-Our results complement those of Berger [2] and are consistent with the prediction of Weinrib and Halperin [21] on the disorder relevance/irrelevance for physical systems with correlated disorder.
+We shall show that, under proper scalings, the rescaled partition functions converge weakly to continuum limits which can be represented by chaos expansions.
+Our results complement those of [4] and are consistent with the prediction of [33] on the disorder relevance/irrelevance for physical systems with correlated disorders.
 1.1. The disordered pinning model
@@
 Pinning models are well suited models for studying phase transitions and critical phenomena and also have received widespread attention in physics, chemistry and biology.
-We refer to the monographs of Giacomin [9] and [10] for more details.
+We refer to the monographs of [15] and [16] for more details.
 We now introduce the setting of our model.
@@
 The expectation with respect to ${\mathbb{P}}_{\tau}$ is denoted by $\mathbb{E}_{\tau}$.
-Note that $\mathbb{E}_{\tau}[\tau_{1}]<\infty$ for $\alpha>1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$ for $\alpha<1$, and by [5, Proposition 1.5.9a], for $\alpha=1$,
+Note that $\mathbb{E}_{\tau}[\tau_{1}]<\infty$ for $\alpha>1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$ for $\alpha<1$, and by [7, Proposition 1.5.9a], for $\alpha=1$,
 $\mathbb{E}_{\tau}[\tau_{1}\mathbf{1}_{\{\tau_{1}\leq N\}}]=l(N):=\sum_{n=1}^{N}\frac{L(n)}{n}$ (1.2)
-is also slowly varying (whether it is finite or not depending on $L(\cdot)$). Hence, the renewal process $\tau$ can be classified by the following conditions:
-(1) $10<{\alpha}<$ 1; (2) $\mathbb{E}_{\tau}[\tau_{1}]<\infty$; (3) $\alpha=1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$. (1.3)
-Noting that our renewal process $\tau$ is aperiodic since $L(n)>0$ for all $n$, we have the renewal theorem (see [8] and [5, Theorem 8.7.5]):
-${\mathbb{P}}_{\tau}(n\in\tau)\sim\begin{cases}\dfrac{C_{\alpha}}{L(n)n^{1-\alpha}},&\text{if }0<\alpha<1~{}(C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}),\\[10.0pt] \dfrac{1}{\mathbb{E}[\tau_{1}]},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \dfrac{1}{l(n)},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}\quad\qquad\text{as}~{}n\to\infty.$ (1.4)
-Let $\omega=\{\omega_{n}\text{,}~{}n\in\mathbb{N}\}$ be a family of Gaussian random variables independent of $\tau$, representing the random environments.
+is also slowly varying (whether it is finite or not depends on $L(\cdot)$). Hence, the renewal process $\tau$ can be classified by the following conditions:
+(1) $0<{\alpha}<1$; (2) $\mathbb{E}_{\tau}[\tau_{1}]<\infty$; (3) $\alpha=1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$. (1.3)
+Noting that our renewal process $\tau$ is aperiodic since $L(n)>0$ for all $n$, we have the renewal theorem (see [14] and [7, Theorem 8.7.5]):
+${\mathbb{P}}_{\tau}(n\in\tau)\sim\begin{cases}\dfrac{C_{\alpha}}{L(n)n^{1-\alpha}},&\text{if }0<\alpha<1~\Big(C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}\Big),\\[10.0pt] \dfrac{1}{\mathbb{E}[\tau_{1}]},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \dfrac{1}{l(n)},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}\quad\qquad\text{as}~n\to\infty.$ (1.4)
+Let $\omega=\{\omega_{n}\text{,}~n\in\mathbb{N}\}$ be a family of Gaussian random variables independent of $\tau$, representing the random environment.
 The probability and expectation for $\omega$ are denoted by ${\mathbb{P}}_{\omega}$ and $\mathbb{E}_{\omega}$, respectively.
 We assume that
-$\mathbb{E}_{\omega}[\omega_{n}]=0\quad\text{and}\quad\mathbb{E}_{\omega}[\omega_{n}\omega_{n^{\prime}}]=\gamma(n-n^{\prime}),\quad\text{for }n,n^{\prime}\in\mathbb{N},$ (1.5)
+[…]
 where the covariance function $\gamma(\cdot)$ satisfies
@@
 We may also use $\mathbb{P}$} and $\mathbb{E}$ to denote the probability and the expectation respectively of some random variables only involving $\tau$ or $\omega$ when there is no ambiguity.
-Denote $\Omega_{N}:=\Big{\{}\frac{1}{N}\text{,}\frac{2}{N}\text{,}\cdots\text{,}\frac{N-1}{N}\text{,}1\Big{\}}$ and $\delta_{x}^{(N)}=\delta_{x}:=\mathbf{1}_{\{Nx\in\tau\}}$.
-The disordered pinning model with free boundary is defined by
-$\frac{{\mathrm{d}}{\mathbb{P}}_{N,\beta}^{\omega,h}}{{\mathrm{d}}{\mathbb{P}}_{\tau}}(\tau):=\frac{1}{Z_{N,\beta}^{\omega,h}}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big{\}},$
+Denote $\Omega_{N}:=\{\frac{1}{N}\text{,}\frac{2}{N}\text{,}\cdots\text{,}\frac{N-1}{N}\text{,}1\}$ and $\delta_{x}^{(N)}=\delta_{x}:=\mathbf{1}_{\{Nx\in\tau\}}$.
+The disordered pinning model with free boundary is defined via Gibbs transform by
+$\frac{{\mathrm{d}}{\mathbb{P}}_{N,\beta}^{\omega,h}}{{\mathrm{d}}{\mathbb{P}}_{\tau}}(\tau):=\frac{1}{Z_{N,\beta}^{\omega,h}}\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big\},$ (1.8)
 where
@@
 $Z_{N,\beta}^{\omega,h,\text{c}}=\mathbb{E}_{\tau}\Big[\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big\}\Big|N\in\tau\Big].$ (1.10)
-1.2. Disorder relevance/irrelevance and the Weinrib-Halperin prediction
+1.2. Disorder relevance/irrelevance and the Weinrib–Halperin prediction
 When the disorder is i.i.d., the Harris criterion [17] indicates that disorder is irrelevant if ${\alpha}<\frac{1}{2}$ and is relevant if ${\alpha}>\frac{1}{2}$ (with $\alpha=\frac{1}{2}$ being the marginal case).
-When the disorder is correlated and the correlation is of the form $n^{-p}$ ($p>0$ and $n$ is the distance between two points), Weinrib and Halperin [21] predicted that the disorder should be relevant if ${\alpha}>\min(p,1)/2$ and irrelevant if ${\alpha}<\min(p,1)/2$.
-Berger [2] confirmed the Weinrib-Halperin prediction for the case $p>1$ (note that in this case the Harris criterion still applies), and showed that the case $p<1$ is different in the sense that classical methods (see, e.g., [10]) in the study of disorder relevance/irrelevance by comparing critical points/exponents no longer apply: when $p<1$, the disordered system does not possess a localization/delocalization phase transition, since the annealed free energy is always infinite and the quenched free energy is always strictly positive for all $h$ as long as $\beta>0$, and thus the system is always localized (see Theorem 2.5 and Section 2.3 in [2]).
-This phenomenon is due to the fact, as pointed out by Berger [3], that when $p<1$, there are large and frequent regions in the environments with arbitrarily high attractions (called the “infinite disorder” in [3]).
-These results hint at scalings of $\beta$ and $h$ when studying disorder relevance/irrelevance for the case $p<1$.
-On the other hand, Caravenna, Sun and Zygouras [6] proposed a new perspective on disorder relevance/irrelevance: noting that disorder relevance means that an arbitrarily small amount of disorder (i.e., for any $\beta>0$) changes the features of the underlying homogeneous model, it is possible that if the strength of disorder is tuned down to zero at a proper rate (i.e., $\beta_{N}\to 0$) as the system size tends to infinity, then the disorder will persist in the resulting continuum model.
+Here, “disorder is relevant” means that for any $\beta>0$, the behavior of $\tau$ under ${\mathbb{P}}_{N,\beta}^{\omega,h}$ differs drastically from that under ${\mathbb{P}}_{\tau}$.
+To be specific, some crucial quantities, such as the correlation length and the critical point shift will be changed by introducing disorder in the Gibbs transform (1.8).
+When disorder is irrelevant, the behavior of $\tau$ under ${\mathbb{P}}_{N,\beta}^{\omega,h}$ is comparable to that under ${\mathbb{P}}_{\tau}$, at least for small enough $\beta>0$.
+Readers may refer to [9, 8, 2, 11] for more detailed discussion about disorder relevance and irrelevance.
+When the disorder is correlated and the correlation is of the form $n^{-p}$ ($p>0$ and $n$ is the distance between two points), [33] predicted that the disorder should be relevant if ${\alpha}>\min(p,1)/2$ and irrelevant if ${\alpha}<\min(p,1)/2$.
+[4] confirmed the Weinrib–Halperin prediction for the case $p>1$ (note that in this case the Harris criterion still applies), and showed that the case $p<1$ is different in the sense that classical methods (see, e.g., [16]) in the study of disorder relevance/irrelevance by comparing critical points/exponents no longer apply: when $p<1$, the disordered system does not possess a localization/delocalization phase transition, since the annealed free energy is always infinite and the quenched free energy is always strictly positive for all $h$ as long as $\beta>0$, and thus the system is always localized (see Theorem 2.5 and Section 2.3 in [4]).
+This phenomenon is due to the fact, as pointed out by [5], that when $p<1$, there are large and frequent regions in the environments with arbitrarily high attractions (called the “infinite disorder” in [5]).
+These results indicate that different methods are needed to study disorder relevance/irrelevance for the case $p<1$.
+On the other hand, [9] proposed a new perspective on disorder relevance/irrelevance: it is possible that if the strength of disorder is tuned down to zero at a proper rate (i.e., $\beta_{N}\to 0$) as the system size tends to infinity, then the disorder will persist in the resulting continuum model.
 This phenomenon has been shown for the i.i.d.
@@
 For related results on directed polymers, we refer to [30, 13] for the case when the disorder is correlated in space but independent in time, to [29] for the case when the disorder is correlated in time but independent in space, and to [31] for the case when the disorder is space-time correlated.
-In those models, the Weinrib-Halperin prediction is not clear (the decay rates in space and in time of the correlation of the disorder in [20] do not agree).
+In those models, the Weinrib–Halperin prediction is not clear (the decay rates in space and in time of the correlation of the disorder in [31] do not agree).
 Inspired by the above-mentioned works, we study the weak convergence of the rescaled partition functions.
-In our setting, noting that $p=2-2H\in(0,1)$ with $H\in(\frac{1}{2},1)$, the Weinrib-Halperin prediction suggests that the disorder relevent if $\alpha>1-H$, which coincides with the condition for the existence of the weak limit for each chaos (see Proposition 3.1).
+In our setting, noting that $p=2-2H\in(0,1)$ with $H\in(\frac{1}{2},1)$, the Weinrib–Halperin prediction suggests that disorder is relevant if $\alpha>1-H$, which coincides with the condition for the existence of the weak limit for each chaos (see Proposition 3.1).
+When $H\in(\frac{1}{2},1)$, under the stronger condition $\alpha\geq\frac{1}{2}$, we establish in Theorem 1.1 that the rescaled Wick-ordered partition functions converge weakly to an $L^{2}$ random variable.
+We further conjecture that for $H\in(\frac{1}{2},1)$, these rescaled Wick-ordered partition functions converge weakly to an $L^{1}$ random variable under the Weinrib–Halperin condition $\alpha>1-H$ (see Remark 1.6 and Section 3.2 for further details).
 Finally, we make a few more remarks on disorder relevance/irrelevance for pinning models in correlated random environment.
-Poisat [17] showed that the Harris criterion still applies to the pinning model in random environments with finite range correlation.
-When disorders have long range correlation, Berger and Poisat [4] conformed the Weinrib-Halperin prediction for the case when the correlations are summable, which extended the result in [2].
+[26] showed that the Harris criterion still applies to the pinning model in random environments with finite range correlation.
+When disorders have long range correlation, [3] confirmed the Weinrib–Halperin prediction for the case when the correlations are summable, which extended the result in [4], and [27] obtained sharp asymptotics of the annealed free energy for doubly summable or exponentially decaying correlations.
+We also mention that [12] studied a disordered pinning model, where the random environment is given by another independent renewal process.
+In that case, the random environment is correlated and it is different from the more typical setting where the random environment is a family of random variables.
+They found that in their model, the chaos expansion heuristics and the Weinrib–Halperin prediction provide different phase diagram.
 1.3. Main results
-Noting that
+Rather than studying the Gibbs measure directly, which is highly involved, we focus on its partition function, which already encapsulates sufficient information about the model. Noting that
 $\mathbb{E}_{\omega}\Big[\exp\Big\{\sum\limits_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big\}\Big]=\exp\Big\{\frac{1}{2}\mathrm{Var}_{\omega}\Big(\sum\limits_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big)\Big\},$
-we first consider the following Wick ordered partition function
-$\tilde{Z}_{N,\beta}^{\omega,h}:=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}_{\omega}\Big{(}\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big{)}\Big{\}}\Big{]}.$ (1.10)
-Note that the term $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{1}{2}\beta^{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n,m\in\tau\}}$ is a direct normalization of the partition function, instead of the Gibbs measure itself, since it depends on the realization of $\tau$.
-We adopt this normalization because the disorders are correlated, and thus a polynomial chaos expansion as has been done in [1, 6] cannot be employed for our pinning model.
-The normalization in the Wick ordered partition function (1.10) results in a Wick exponential, which allows us to perform a Wick expansion (see Section 3 and also the discussion in [13]).
-We point out that the Wick ordered partition function (1.10) reduces to [6, equation (3.3)] when the disorders are i.i.d., noting that $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{\beta^{2}}{2}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}$.
+we first consider the following Wick-ordered partition function
+$\tilde{Z}_{N,\beta}^{\omega,h}:=\mathbb{E}_{\tau}\Big[\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}_{\omega}\Big(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big)\Big\}\Big],$ (1.11)
+where the term $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{1}{2}\beta^{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n,m\in\tau\}}$ is a direct normalization of the partition function (we call it Wick normalization), which depends on the realization of $\tau$.
+We adopt this normalization because disorder is correlated, and thus a polynomial chaos expansion as has been done in [1, 9] cannot be employed for our pinning model.
+The normalization in the Wick-ordered partition function (1.11) results in a Wick exponential, which allows us to perform a Wick expansion (see Section 3 and also the discussion in [21]).
+We point out that the Wick-ordered partition function (1.11) reduces to [9, equation (3.3)] when the disorders are i.i.d., noting that $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{1}{2}\beta^{2}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}$.
+After addressing the Wick-ordered partition function, we turn to the original partition function $Z_{N,\beta}^{\omega,h}$ given in (1.9), for which we carry out a Taylor expansion.
+The finite-order chaos in the Wick and Taylor expansions is connected by the Hu–Meyer formula (see Eq.
+(2.10) in Section 2).
 Let $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$ be fixed constants throughout this paper. Consider the following scalings:
-$\beta_{N}=\begin{cases}\hat{\beta}\dfrac{L(N)}{N^{\alpha+H-1}},&\text{if }0<\alpha<1,\\[10.0pt] \hat{\beta}\dfrac{1}{N^{H}},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \hat{\beta}\dfrac{l(N)}{N^{H}},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}h_{N}=\begin{cases}\hat{h}\dfrac{L(N)}{N^{\alpha}},&\text{if }0<\alpha<1,\\[10.0pt] \hat{h}\dfrac{1}{N},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \hat{h}\dfrac{l(N)}{N},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}$ (1.11)
+[…]
 where we recall that $l(n)$ given in (1.2) is a slowly varying function.
-Let $\phi_{k}(t_{1},\dots,t_{k})$ be a symmetric function, defined for $10<t_{1}<\cdots<t_{k}<$ 1 by
-$\phi_{k}(t_{1},\dots,t_{k}):=\begin{cases}\displaystyle\frac{C_{\alpha}^{k}}{t_{1}^{1-\alpha}(t_{2}-t_{1})^{1-\alpha}\cdots(t_{k}-t_{k-1})^{1-\alpha}},&\text{ if }0<\alpha<1,\\[10.0pt] \displaystyle\Big{(}\frac{1}{\mathbb{E}[\tau_{1}]}\Big{)}^{k},&\text{ if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \displaystyle 1,&\text{ if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}$ (1.12)
-with $C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}$ as in (1.4).
-Now we present our main results for the Wick ordered partition function $\tilde{Z}_{N,\beta}^{\omega,h}$ given in (1.10) and for the original partition function $Z_{N,\beta}^{\omega,h}$ given in (1.8), respectively.
+Let $\phi_{k}(t_{1},\dots,t_{k})$ be a symmetric function.
+For any given $(t_{1},\cdots,t_{k})\in(0,1)^{k}$ with $t_{i}\neq t_{j}$ for any $i\neq j$ (if there exist $i\neq j$, s.t.
+$t_{i}=t_{j}$, we set $\phi_{k}=0$), there exists a unique permutation $\sigma$ on $\llbracket k\rrbracket$ such that $0<t_{\sigma(1)}<t_{\sigma(2)}<\cdots<t_{\sigma(k)}<1$.
+We define $\phi_{k}$ as follows:
+[…]
+where $C_{\alpha}$ is the same as that in (1.4).
+Now we present our main results for the Wick-ordered partition function $\tilde{Z}_{N,\beta}^{\omega,h}$ given in (1.11) and for the original partition function $Z_{N,\beta}^{\omega,h}$ given in (1.9), respectively.
+The weak limits $\tilde{Z}_{\hat{\beta},\hat{h}}$ in (1.14) and $Z_{\hat{\beta},\hat{h}}$ in (1.15) involve multiple stochastic integrals in the sense of Skorohod and Stratonovich (see Section 2 for their definitions), respectively.
+Roughly speaking, the multiple Skorohod integrals in (1.14) (resp.
+the multiple Stratonovich integrals in (1.15)) can be defined as the limit of Riemann sums, where the products are taken as Wick products (resp.
+ordinary products).
+Although multiple Skorohod and Stratonovich integrals are related through the Hu–Meyer formula (2.10), it is not straightforward to establish a direct connection between $\tilde{Z}_{\hat{\beta},\hat{h}}$ and $Z_{\hat{\beta},\hat{h}}$.
 Theorem 1.1.
-Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H and $\alpha$ be parameters satisfying
-$H\in\Big{(}\frac{1}{2},1\Big{)}\text{ and }~{}\alpha>\frac{1}{2}.$
-Recalling the Wick ordered partition function $\tilde{Z}^{\omega,h}_{N,\beta}$ given in (1.10) and the scalings (1.11), we have that
+Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1.
+Let H be the parameter in (1.6)-(1.7) with $H\in(\frac{1}{2},1)$.
+Recall the Wick-ordered partition function $\tilde{Z}^{\omega,h}_{N,\beta}$ given in (1.11) and the scalings (1.12).
+We have the following three cases for different $\alpha>0$ in (1.1).
+Case (1): If $\alpha>\frac{1}{2}$, then for any $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$, we have that
+$\tilde{Z}^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}~,\text{ as }N\to\infty.$
+Here $\tilde{Z}_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
+$\tilde{Z}_{\hat{\beta},\hat{h}}:=1+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod\limits_{i=1}^{k}\diamond\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right),$ (1.14)
+where $\phi_{k}$ is given by (1.13), $\prod\diamond$ is the Wick product, $\dot{W}$ is Gaussian noise with $\mathbb{E}[\dot{W}(t)\dot{W}(s)]=|t-s|^{2H-2}$ and the stochastic integral is of Skorohod type.
+Case (2): If $\alpha=\frac{1}{2}$ and we further assume that the slowly varying function $L(\cdot)$ in (1.1) satisfies $0<\inf_{n\in\mathbb{N}}L(n)\leq\sup_{n\in\mathbb{N}}L(n)<+\infty$, then there exist some $\hat{\beta}_{c},\hat{h}_{c}\in(0,+\infty)$, such that for $\hat{\beta}\in(0,\hat{\beta}_{c})$ and $|\hat{h}|<\hat{h}_{c}$, we have the same weak convergence as in Case (1), that is,
 $\tilde{Z}^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}~,\text{ as }N\to\infty,$
-where $\tilde{Z}_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
-$\tilde{Z}_{\hat{\beta},\hat{h}}:=1+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod\limits_{i=1}^{k}\diamond\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right),$ (1.13)
-where $\phi_{k}$ is given by (1.12), $\dot{W}$ is Gaussian noise with $\mathbb{E}[\dot{W}(t)\dot{W}(s)]=|t-s|^{2H-2}$ and the stochastic integral is of Skorohod type.
+where $\tilde{Z}_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series (1.14).
+Case (3): If $\alpha<\frac{1}{2}$, then for any $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$, we have
+$\lim\limits_{N\to\infty}\mathbb{E}\Big[\Big(\tilde{Z}^{\omega,h_{N}}_{N,\beta_{N}}\Big)^{2}\Big]=+\infty.$
+Remark 1.2.
+Note that $\alpha=\tfrac{1}{2}$ is the threshold between the $L^{2}$ regime and the non-$L^{2}$ regime, which will be further discussed in Remark 3.3.
+The case $\alpha<\tfrac{1}{2}$ will be addressed in Remark 1.6 below.
 Theorem 1.3.
-Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H and $\alpha$ be parameters satisfying
-$H\in\Big{(}\frac{1}{2},1\Big{)}\text{ and }~{}\alpha+H>\frac{3}{2}.$
-Recalling the partition function $Z^{\omega,h}_{N,\beta}$ given in (1.8) and the scalings (1.11), we have that
-$Z^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}Z_{\hat{\beta},\hat{h}}~{},~{}as~{}N\to\infty,$
-where $Z_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
+Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1.
+Let H and $\alpha$ be parameters satisfying $H\in(\frac{1}{2},1)$ and $\alpha+H>\frac{3}{2}$.
+Recalling the partition function $Z^{\omega,h}_{N,\beta}$ given in (1.9) and the scalings (1.12), we have that for any $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$,
+$Z^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}Z_{\hat{\beta},\hat{h}}~,~as~N\to\infty.$
+Here $Z_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
 $Z_{\hat{\beta},\hat{h}}:=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod_{i=1}^{k}\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right),$ (1.15)
@@
 Remark 1.4.
-The results in Theorems 1.1 and 1.2 also hold for the conditioned partition functions
-$\tilde{Z}_{N}^{\text{c}}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N},\text{c}}:=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}\Big{[}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx})\delta_{x}\Big{]}\Big{\}}\Big{|}N\in\tau\Big{]}$
-and
+Assuming the conditions $H\in(\frac{1}{2},1)$ and $\alpha+H>\frac{3}{2}$ as in Theorem 1.3, the rescaled original partition function $Z^{\omega,h_{N}}_{N,\beta_{N}}$ is actually uniformly bounded in $L^{1}$ (see Section 4.3), and thus no normalization is needed in this situation.
+In contrast, under the weaker conditions $H\in\bigl(\tfrac{1}{2},1\bigr)$ and $\alpha\geq\frac{1}{2}$ assumed in Cases (1) and (2) in Theorem 1.1, Wick normalization is employed to obtain weak convergence to a nontrivial random variable.
+In our setting with colored Gaussian noise, Wick normalization provides more convenience for computations than the “standard” normalization.
+Furthermore, since the polymer measure acts as a random measure of the underlying renewal process, it is natural to adopt Wick normalization so that the normalized partition function has a mean equal to the homogeneous partition function (i.e., $Z_{N,\beta=0}^{\omega,h}$), which is $1$ when $h=0$.
+This viewpoint has also been adopted very recently in [28] for the parabolic Anderson model on $\mathbb{R}^{2}$, where the classical Skorohod integral was extended to the $L^{1}$ setting and the global existence of an $L^{1}$ Skorohod solution of the stochastic heat equation with multiplicative planar white noise was proved.
+For the directed polymer model in time-correlated random environment, the continuum Skorohod (resp.
+Stratonovich) partition function corresponds to the Skorohod (resp.
+Stratonovich) solution of the stochastic heat equation driven by time-colored Gaussian noise (see [29]).
+However, the literature on SPDEs related to pinning models in random environments is very limited.
+A stochastic heat equation with multiplicative white noise, corresponding to the marginally relevant pinning model with $\alpha=\tfrac{1}{2}$ in an independent random environment, was recently studied in [32].
+In our forthcoming work [22], we investigate a class of fractional stochastic heat equations in the sense of Skorohod (resp.
+Stratonovich) associated with the Skorohod (resp.
+Stratonovich) continuum partition functions of random pinning models.
+Remark 1.5.
+The results in Theorems 1.1 and 1.3 also hold for the conditioned partition functions
+$\tilde{Z}_{N}^{\text{c}}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N},\text{c}}:=\mathbb{E}_{\tau}\Big[\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}\Big[\sum_{x\in\Omega_{N}}(\beta\omega_{Nx})\delta_{x}\Big]\Big\}\Big|N\in\tau\Big],~\text{and}$
 ${Z}_{N}^{\text{c}}={Z}_{N,\beta_{N}}^{\omega,h_{N},\text{c}}:=\mathbb{E}_{\tau}\Big[\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big\}\Big|N\in\tau\Big],$
-with the integrand $\phi_{k}$ in (1.13) and (1.14) being replaced by a symmetric function $\phi_{k}^{c}$, defined for $10<t_{1}<\cdots<t_{k}<$ 1 by
-$\phi_{k}^{c}(t_{1},\dots,t_{k}):=\begin{cases}\displaystyle\frac{C_{\alpha}^{k}}{t_{1}^{1-\alpha}(t_{2}-t_{1})^{1-\alpha}\cdots(t_{k}-t_{k-1})^{1-\alpha}(1-t_{k})^{1-\alpha}},&\text{ if }1-H<\alpha<1,\\[10.0pt] \displaystyle\Big{(}\frac{1}{\mathbb{E}[\tau_{1}]}\Big{)}^{k},&\text{ if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \displaystyle 1,&\text{ if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty.\end{cases}$
+with the integrand $\phi_{k}$ in (1.14) and (1.15) being replaced by a symmetric function $\phi_{k}^{c}$ with the same symmetrization as for $\phi_{k}$, which for $0<t_{1}<\cdots<t_{k}<1$ is defined by
+[…]
 Remark 1.6.
-A natural condition for the rescaled Wick ordered partition function to converge weakly is $\alpha+H-1>0$, where $\alpha+H-1$ is the scaling exponent for $\beta$ (see (1.11)).
+A natural condition for the rescaled Wick-ordered partition function to converge weakly is $\alpha+H-1>0$, where $\alpha+H-1$ is the scaling exponent for $\beta$ (see (1.12)).
 Note that for directed polymers, under the condition that the corresponding scaling exponent is positive, the continuum partition converges in $L^{2}$ as a Wiener chaos (see [9, 13, 29]).
-However, for our pinning model, assuming $\alpha+H-1>0$ we can only prove that each chaos is a well-defined Skorohod integral which is $L^{2}$-integrable (see Lemma 3.2), and we need a stronger condition $H\in(\frac{1}{2},1),\alpha>\frac{1}{2}$ to guarantee the $L^{2}$-convergence as stated in Theorem 1.1 (also see Remark 3.3 for the criticality of $\alpha=\frac{1}{2}$).
-This is a new phenomenon that does not appear in the models studied in [1, 6, 18, 19, 7]: in those models, the rescaled partition functions are always uniformly bounded in $L^{2}$ if the models are in the disorder relevant regime, no matter the random environments are independent or correlated.
+However, for our pinning model, assuming $\alpha+H-1>0$ we can only prove that each chaos is a well-defined Skorohod integral which is $L^{2}$-integrable (see Lemma 3.2), and we need a stronger condition $H\in(\frac{1}{2},1),\alpha\geq\frac{1}{2}$ to guarantee the $L^{2}$-convergence as stated in Theorem 1.1.
+On the other hand, when $H\in(\frac{1}{2},1)$, we show in Proposition 3.5 that, assuming $\alpha>2-2H$, $\big\{\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big\}_{N}$ is uniformly integrable.
+Note that in this situation, when $H\in(\frac{3}{4},1)$, $\alpha$ may be strictly less than $\frac{1}{2}$, and thus $\big\{\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big\}_{N}$ is not $L^{2}$-bounded.
+This is a new phenomenon that does not appear in the models studied in [1, 9, 30, 29, 13]: in those models, the rescaled partition functions are always uniformly bounded in $L^{2}$ if the models are in the disorder relevant regime, no matter whether the random environments are independent or correlated.
 This hints that the pinning model in correlated environment indeed carries some new features.
-Motivated by the Weinrib-Halperin prediction, we conjecture that under the condition $\alpha+H-1>0$ with $\alpha\leq\frac{1}{2}$, the rescaled Wick ordered partition function $\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}$ converges weakly to some $L^{1}$-integrable random variable.
-We will discuss further on the uniform integrability and the weak limit of $\big{\{}\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\big{\}}_{N}$ in Section 3.2.
+We note that a phase transition from the $L^{2}$ regime to an $L^{p}$ regime with $1<p<2$ has been studied in the directed polymer model with fixed inverse temperature $\beta$.
+In particular, [20] showed that, while the polymer measure exhibits the same macroscopic behavior in both the $L^{2}$ and $L^{p}$ regimes, its microscopic behavior differs, as evidenced by distinct exponents in the corresponding local limit theorems.
+By analogy, we expect that this new phenomenon of the gap between the $L^{2}$ regime and the $L^{1}$ regime in our model should relate to some phase transition.
+Motivated by the Weinrib–Halperin prediction and the above discussion, we conjecture that under the condition $\alpha+H-1>0$ with $\alpha\leq\frac{1}{2}$, the rescaled Wick-ordered partition function $\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}$ converges weakly to some $L^{1}$-integrable random variable for any $\hat{\beta}>0$.
+We will discuss the conjecture in more detail in Section 3.2.
 To close this section, let us highlight our contributions in this paper.
-Firstly, we obtain some finer property of a disordered pinning model with environment correlation of the form $n^{-p}$, where $p<1$, which complements the result of Berger [2].
-To be specific, we find scaling limits of rescaled (Wick ordered) partition functions in the disorder relevant regime.
-In particular, the Weinrib-Halperin prediction is partially conformed when it differs from the Harris criterion.
-We also show that the rescaled (Wick ordered) partition functions are not $L^{2}$-bounded in the entire disorder relevance regime (see Remark 3.3), which is a new feature compared to other disordered models.
+Firstly, we obtain some finer property of a disordered pinning model with environment correlation of the form $n^{-p}$, where $p<1$, which complements the result of [4].
+To be specific, we find scaling limits of rescaled (Wick-ordered) partition functions in the disorder relevant regime.
+In particular, the Weinrib–Halperin prediction is partially confirmed when it differs from the Harris criterion.
+We also show that the rescaled (Wick-ordered) partition functions are not $L^{2}$-bounded in the entire disorder relevance regime (see Remark 3.3), which is a new feature compared to other disordered models.
 Secondly, we develop a decoupling approach to prove uniform integrability (see Section 3.2), and provide some sharp upper bounds for the norm of traces (see Lemma 4.1) and for the second moments of multiple Stratonovich integrals (see Lemma 4.4).
 These auxiliary results may be of independent interest.
-Notations 1.5.
-[…]
+Notations 1.7.
+Let $\mathbb{N}$ denote the set of natural numbers without 0, i.e., $\mathbb{N}:=\{1,2,\cdots\}$ and $\mathbb{N}_{0}:=\{0,1,2,\cdots\}$; for $N\in\mathbb{N}$, $\llbracket N\rrbracket:=\{1,2,\cdots,N\}$; For $\in\mathbb{R}, [a]\left[a\right$]means the greatest integer that is not greater than $a$; $\|\cdot\|$ is used for the Euclidean norm; let $\boldsymbol{k}:=(k_{1},k_{2},\cdots,k_{d}),\boldsymbol{x}:=(x_{1},x_{2},\cdots,x_{d})$ etc.
+stand for vectors in ${\mathbb{Z}}^{d}$ or $\mathbb{R}^{d}$ depending on the context; we use C to denote a generic positive constant that may vary in different lines; we say $f(x)\lesssim g(x)$ if $f(x)\leq Cg(x)$ for all $x$; we write $f(x)\sim g(x)(as\ x\rightarrow\infty)$ if $\lim_{x\rightarrow\infty}f(x)/g(x)=1$; we use $\overset{(d)}{\longrightarrow}$ to denote the convergence in distribution (also called the weak convergence) for random variables or random vectors.
+For a random variable X, $\|X\|_{L^{p}}:=(\mathrm{E}\left[|X|^{p}\right])^{1/p}$ for $p\geq 1$.
+We use $\mathrm{I}_{A}$ to denote the indicator function, i.e., $\mathrm{I}_{A}(\boldsymbol{t})=1$ if $\boldsymbol{t}\in A$ and $\mathrm{I}_{A}(\boldsymbol{t})=0$ if $\boldsymbol{t}\notin A$.
 The rest of the paper is organized as follows.
@@
 The proofs of Theorem 1.1 and Theorem 1.3 will be presented in Section 3 and Section 4, respectively.
-We will only deal with the case $10<\alpha<$ 1 in the proofs, noting that the other case $\alpha\geq 1$ can be treated in the same way.
+We will only deal with the case $0<\alpha<1$ in the proofs, noting that the other case $\alpha\geq 1$ can be treated in the same way.
 Some technical estimates are put in Appendix A.
@@
 In this subsection, we collect some preliminaries on the analysis of Gaussian space. We refer to [24, 18] for more details.
-On a probability space $(\Omega,\mathcal{F},{\mathbb{P}})$ satisfying usual conditions, let $\dot{W}=\{\dot{W}(t):t\in[0,1]\}$ be a real-valued centered Gaussian noise with covariance
+On a probability space $(\Omega,\mathcal{F},{\mathbb{P}})$ satisfying the usual conditions, let $\dot{W}=\{\dot{W}(t):t\in[0,1]\}$ be a real-valued centered Gaussian noise with covariance
 $\mathbb{E}[\dot{W}(t)\dot{W}(s)]=K(t-s)=|t-s|^{2H-2},$ (2.1)
@@
 The Hilbert space $\mathcal{H}$ associated with $\dot{W}$ is the completion of smooth functions with compact support under the inner product:
-$f,g\right>_{\mathcal{H}}$ $=\int_{0}^{1}\int_{0}^{1}f(s)K(s-t)g(t){\mathrm{d}}s{\mathrm{d}}t=\int_{0}^{1}\int_{0}^{1}f(s)g(t)|s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t.$ (2.2)
+$\left<f,g\right$$_{\mathcal{H}}$ $=\int_{0}^{1}\int_{0}^{1}f(s)K(s-t)g(t){\mathrm{d}}s{\mathrm{d}}t=\int_{0}^{1}\int_{0}^{1}f(s)g(t)|s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t.$ (2.2)
 Noting that $\mathcal{H}$ contains distributions (see [25]), it is convenient to consider $|\mathcal{H}|$ which is a linear subspace of measurable functions $f$ on $[0,1]$ satisfying
-$\|f\|_{|\mathcal{H}|}^{2}:=\int_{0}^{1}\int_{0}^{1}|f(s)g(t)||s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t<\infty.$ (2.3)
+$\|f\|_{|\mathcal{H}|}^{2}:=\int_{0}^{1}\int_{0}^{1}|f(s)f(t)||s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t<\infty.$ (2.3)
 Clearly, $|\mathcal{H}|$ is a Banach space and is dense in $\mathcal{H}$.
+Remark 2.1.
+The inner product defined in (2.2) is indeed positive definite since for $H\in(1/2,1)$,
+$\left<f,f\right$$_{\mathcal{H}}$ $=\int_{0}^{1}\int_{0}^{1}f(s)f(t)|s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t$
+$=\int_{0}^{1}\int_{0}^{1}f(s)f(t)\left(C_{H}\int_{\mathbb{R}}|s-r|^{H-\frac{3}{2}}|t-r|^{H-\frac{3}{2}}{\mathrm{d}}r\right){\mathrm{d}}s{\mathrm{d}}t$
+$=C_{H}\int_{\mathbb{R}}\left(\int_{0}^{1}f(s)|s-r|^{H-\frac{3}{2}}{\mathrm{d}}s\right)^{2}{\mathrm{d}}r\geq 0.$
+Also note that
+$\|f\|_{L^{1}(\mathbb{R})}^{2}\leq\int_{0}^{1}\int_{0}^{1}|f(s)f(t)||s-t|^{2H-2}{\mathrm{d}}s{\mathrm{d}}t\leq C_{H}\|f\|_{L^{1/H}(\mathbb{R})}^{2},$
+where the second inequality is due to Lemma A.4. Hence, we have $L^{1/H}\subset|\mathcal{H}|\subset L^{1}$. For more details, we refer to Section 5.1.3 in [24].
 Let $\{W(f),f\in\mathcal{H}\}$ be an isonormal Gaussian process with covariance $\mathbb{E}[W(f)W(g)]=\left\langle f,g\right\rangle_{\mathcal{H}}.$ The Wiener integral $W(f)$ is often denoted in the integral form:
@@
 For $f\in\mathcal{H}$ with $\|f\|_{\mathcal{H}}=1$, the multiple Wiener integral of $f^{\otimes n}$ is defined by
-$\mathbf{I}_{n}(f^{\otimes n}):=H_{n}(W(f)).$ (2.4)
-Let ^${\mathcal{H}}^{\hat{\otimes}k}$ be the subspace of ${\mathcal{H}}^{\otimes k}$ containing only symmetric functions.
+$\mathbf{I}_{n}(f^{\otimes n}):=H_{n}(W(f)),$ (2.4)
+where $f^{\otimes n}$ is a function on $0,1]^{n}$ defined by $f^{\otimes n}(t_{1},\cdots,t_{n}):=\prod_{k=1}^{n}f(t_{k})$.
+As in (2.2) and (2.3), let the Hilbert space ${\mathcal{H}}^{\otimes k}$ be the completion of smooth functions with compact support under the inner product:
+$\left<f,g\right>_{\mathcal{H}^{\otimes k}}$ $=\int_{[0,1]^{2k}}f(s_{1},\cdots,s_{k})g(t_{1},\cdots,t_{k})\prod_{i=1}^{k}|s_{i}-t_{i}|^{2H-2}{\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{t},$ (2.5)
+and we also denote by $|\mathcal{H}|^{\otimes k}$ the linear subspace of measurable functions $f$ on $0,1]^{k}$ satisfying
+$\|f\|_{|\mathcal{H}|^{\otimes k}}^{2}:=\int_{[0,1]^{2k}}|f(s_{1},\cdots,s_{k})g(t_{1},\cdots,t_{k})|\prod_{i=1}^{k}|s_{i}-t_{i}|^{2H-2}{\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{t}<\infty.$ (2.6)
+Let ${\mathcal{H}}^{\hat{\otimes}k}$ be the subspace of ${\mathcal{H}}^{\otimes k}$ containing only symmetric functions.
 For $f\in{\mathcal{H}}^{\hat{\otimes}k}$, the $k$-th multiple Wiener integral $\mathbf{I}_{k}(f)$ can be defined by (2.4) via a limiting argument, and moreover,
 ${\mathbb{E}}[\mathbf{I}_{m}(f)\mathbf{I}_{n}(g)]=n!\langle f,g\rangle_{\mathcal{H}^{\otimes n}}\mathbf{1}_{\{m=n\}},~{\rm for}~f\in{\mathcal{H}}^{\hat{\otimes}m},g\in{\mathcal{H}}^{\hat{\otimes}n}.$
-For general $f\in{\mathcal{H}}^{\otimes k}$, we simply set $\mathbf{I}_{k}(f):=\mathbf{I}_{k}(\hat{f}),$ where $\hat{f}$ denotes the symmetrization of $f$.
-We also use the following integral form for multiple Wiener integrals:
+[…]
 $\int_{[0,1]^{n}}f({\mathbf{t}})W({\mathrm{d}}{t_{1}})\diamond\cdots\diamond W({\mathrm{d}}{t_{n}})=\int_{[0,1]^{n}}f({\mathbf{t}})\prod_{i=1}^{n}\diamond W({\mathrm{d}}{t_{i}}):=\mathbf{I}_{n}(f).$
@@
 $\mathbf{I}_{n}(f)\diamond\mathbf{I}_{m}(g):=\mathbf{I}_{n+m}(f\otimes g).$ (2.7)
-For square integrable random variables $F$ and $G$ with Wiener chaos expansions $F=\sum_{k=0}^{\infty}\mathbf{I}_{k}(f_{k})$ and $G=\sum_{k=0}^{\infty}\mathbf{I}_{k}(g_{k})$ respectively, we have $F\diamond G=\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\mathbf{I}_{n+m}(f_{n}\otimes g_{m})$, as long as the right-hand side term exists in some proper sense.
+For square integrable random variables $F$ and $G$ with Wiener chaos expansions $F=\sum_{k=0}^{\infty}\mathbf{I}_{k}(f_{k})$ and $G=\sum_{k=0}^{\infty}\mathbf{I}_{k}(g_{k})$ respectively, we have $F\diamond G=\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\mathbf{I}_{n+m}(f_{n}\otimes g_{m})$, as long as the right-hand side term is well defined (i.e., the series converges in $L^{2}$).
 For a centered Gaussian random variable $F$, its Wick exponential is defined by
 $\exp^{\diamond}(F):=\exp\left(F-\frac{\mathbb{E}[F^{2}]}{2}\right)=\sum_{m=0}^{\infty}\frac{1}{m!}F^{\diamond m}.$ (2.8)
-The multiple Stratonovich integral of $g^{\otimes n}$ for $g\in\mathcal{H}$ is defined by $\mathbb{I}_{m}(g^{\otimes n}):=W(g)^{n},$ and $\mathbb{I}_{n}(g)$ for $g\in\mathcal{H}^{\hat{\otimes}n}$ by a limiting argument.
+The multiple Stratonovich integral of $g^{\otimes n}$ for $g\in\mathcal{H}$ is defined by $\mathbb{I}_{n}(g^{\otimes n}):=W(g)^{n},$ and $\mathbb{I}_{n}(g)$ for $g\in\mathcal{H}^{\hat{\otimes}n}$ by taking limits for linear combinations (see [18] for more details).
 We also take the following notation for $\mathbb{I}_{n}(f)$
@@
 For $f\in\mathcal{H}^{\otimes n}$ and $k=0,1,\cdots,\lfloor\frac{n}{2}\rfloor$, define the $k$-th order trace $\mathrm{Tr}^{k}f$ of $f$ by
-[…]
-where $K(x-y)$ is given in (2.1). Multiple Stratonovich integrals $\mathbb{I}_{m}(f)$ and multiple Wiener integrals $\mathbf{I}_{m}(f)$ are related via the following celebrated Hu-Meyer’s formula:
+$\mathrm{Tr}^{k}f(t_{1},.,t_{n-2k})$ $\overset{\text{def}}{=}\int_{[0,1]^{2k}}f(s_{1},.,s_{2k},t_{1},.,t_{n-2k})\prod\limits_{i=1}^{k}K(s_{2i}-s_{2i-1}){\mathrm{d}}s_{2i-1}{\mathrm{d}}s_{2i},$ (2.9)
+where K(xy)K(x-y) is given in (2.1). Multiple Stratonovich integrals $\mathbb{I}_{m}(f)$ and multiple Wiener integrals $\mathbf{I}_{m}(f)$ are related via the following celebrated Hu–Meyer formula:
 $\mathbb{I}_{n}(f)=\sum_{k=0}^{[\frac{n}{2}]}\frac{n!}{k!(n-2k)!2^{k}}\mathbf{I}_{n-2k}(\mathrm{Tr}^{k}\hat{f}),\quad\text{for}~f\in\mathcal{H}^{\otimes n},$ (2.10)
-as long as $\vvvert f\vvvert_{n}<\infty$, where
-$\vvvert f\vvvert^{2}_{n}:=\mathbb{E}[|\mathbb{I}_{n}(f)|^{2}]=\sum_{k=0}^{[\frac{n}{2}]}\frac{1}{(n-2k)!}\left(\frac{n!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\hat{f}\|^{2}_{\mathcal{H}^{\otimes(n-2k)}}.$ (2.9)
+as long as $\interleave f\interleave_{n}<\infty$, where
+$\interleave f\interleave^{2}_{n}=\mathbb{E}[|\mathbb{I}_{n}(f)|^{2}]=\sum_{k=0}^{[\frac{n}{2}]}\frac{1}{(n-2k)!}\left(\frac{n!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\hat{f}\|^{2}_{\mathcal{H}^{\otimes(n-2k)}}.$ (2.11)
 2.2. U-statistics
@@
 We denote
-[…]
+$\mathcal{R}_{N}^{m}:=\Big\{\Big(\frac{\boldsymbol{i}-\vec{1}_{m}}{N},\frac{\boldsymbol{i}}{N}\Big]:\boldsymbol{i}\in\llbracket N\rrbracket^{m},\vec{1}_{m}\text{ is the $m$-dimensional vector of all ones}\Big\}.$
 Given $f\in|\mathcal{H}|^{\otimes m}$, we denote by $\mathcal{A}_{N}(f)$ its conditional expectation with respect to the $\sigma$-field generated by $\mathcal{R}_{N}^{m}$, i.e.,
-[…]
+$\mathcal{A}_{N}(f)(\boldsymbol{t})=\sum_{B\in\mathcal{R}_{N}^{m}}\frac{1}{|B|}\int_{B}f(\boldsymbol{s})d\boldsymbol{s}\cdot\mathrm{I}_{B}(\boldsymbol{t}),$ (2.12)
 where $|B|$ is the Lebesgue measure of B.
-We introduce the U-statistics for our model, which appears in the chaos expansion for the Wick ordered partition function (1.10).
+We introduce the U-statistics for our model, which appears in the chaos expansion for the Wick-ordered partition function (1.11).
 Definition 2.3.
 For $f\in|\mathcal{H}|^{\otimes m}$, let
-$\mathbf{I}_{m}^{(N)}(f)$ $:=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right)\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}$ (2.11)
-$:=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)(\boldsymbol{t})\,\omega_{\boldsymbol{n}}^{\diamond},$
+$\mathbf{I}_{m}^{(N)}(f)$ $=N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right)\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}=:N^{-mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(f)(\boldsymbol{t})\,\omega_{\boldsymbol{n}}^{\diamond},$ (2.13)
 where we denote
-$\omega_{\boldsymbol{n}}^{\diamond}:=\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}~{}\text{ and }~{}\boldsymbol{t}=(t_{1}\cdots t_{m}):=\frac{\boldsymbol{n}}{N}=\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right).$ (2.12)
+$\omega_{\boldsymbol{n}}^{\diamond}:=\omega_{n_{1}}\diamond\cdots\diamond\omega_{n_{m}}\quad\text{ and }\quad\boldsymbol{t}=(t_{1}\cdots t_{m}):=\frac{\boldsymbol{n}}{N}=\left(\frac{n_{1}}{N},\cdots,\frac{n_{m}}{N}\right).$ (2.14)
 Lemma 2.4.
 For $m,N\in\mathbb{N}$ and $f\in|\mathcal{H}|^{\otimes m}$, let $\mathcal{A}_{N}(f)$ be defined as above. Then we have,
-$\|\mathcal{A}_{N}(f)\|_{|\mathcal{H}|^{\otimes m}}\leq C^{m}\|f\|_{|\mathcal{H}|^{\otimes m}}$
+$\|\mathcal{A}_{N}(f)\|_{|\mathcal{H}|^{\otimes m}}\leq C^{m}\|f\|_{|\mathcal{H}|^{\otimes m}},$
 where C is a constant depending only on H.
 Proof.
-The desired result follows directly from [19, Lemma C.1]. ∎
+The desired result follows directly from [29, Lemma C.1]. ∎
 Lemma 2.5.
@@
 Proof.
-For $\boldsymbol{n},\boldsymbol{n}^{\prime}\in\llbracket N\rrbracket^{m}$, by Lemma A.2, we have:
-$\mathbb{E}\left[\omega_{\boldsymbol{n}}^{\diamond}~{}\omega_{\boldsymbol{n^{\prime}}}^{\diamond}\right]=\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big{(}n_{i}-n^{\prime}_{\sigma(i)}\big{)},$
-where $\mathcal{P}_{m}$ denotes the set of all permutations on $\llbracket m\rrbracket$. Thus, we have
-$\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{m}^{(N)}(f)\Big{)}^{2}\Big{]}$ $=\mathbb{E}\Big{[}\Big{(}\mathbf{I}_{m}^{(N)}(\hat{f})\Big{)}^{2}\Big{]}$
-$=N^{-2mH}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big{(}n_{i}-n^{\prime}_{\sigma(i)}\big{)}$
+For $\boldsymbol{n},\boldsymbol{n}^{\prime}\in\llbracket N\rrbracket^{m}$, by Lemma A.2, we have $\mathbb{E}\left[\omega_{\boldsymbol{n}}^{\diamond}~\omega_{\boldsymbol{n^{\prime}}}^{\diamond}\right]=\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big(n_{i}-n^{\prime}_{\sigma(i)}\big)$, where $\mathcal{P}_{m}$ denotes the set of all permutations on $\llbracket m\rrbracket$.
+Thus, we have
+$\mathbb{E}\Big[\Big(\mathbf{I}_{m}^{(N)}(f)\Big)^{2}\Big]$ $=\mathbb{E}\Big[\Big(\mathbf{I}_{m}^{(N)}(\hat{f})\Big)^{2}\Big]=N^{-2mH}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\sum_{\sigma\in\mathcal{P}_{m}}\prod_{i=1}^{m}\gamma\big(n_{i}-n^{\prime}_{\sigma(i)}\big)$
 $=m!N^{-2mH}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\prod_{i=1}^{m}\gamma(n_{i}-n^{\prime}_{i}),$
-where the last equality follows from the symmetry of $\mathcal{A}_{N}(\hat{f})$.
+[…]
 For $t\in\mathbb{R}$, let
 $\gamma_{N}(t):=N^{2-2H}\gamma([|t|N]).$ (2.15)
-By (1.6) and (1.7), we have
-$\gamma_{N}(t)\lesssim|t|^{2H-2}\text{ and }\lim\limits_{N\rightarrow\infty}\gamma_{N}(t)=|t|^{2H-2}.$
-Therefore, we get by a Riemann sum approximation,
+By (1.6) and (1.7), we have $\gamma_{N}\lesssim|t|^{2H-2}$ and $\lim_{N\rightarrow\infty}\gamma_{N}=|t|^{2H-2}$. Therefore, we get by a Riemann sum approximation,
 $\mathbb{E}\Big[\left(\mathbf{I}_{m}^{(N)}(f)\right)^{2}\Big]$ $=m!N^{-2m}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\prod_{i=1}^{m}N^{2-2H}\gamma(n_{i}-n^{\prime}_{i})$
-$=m!N^{-2m}\sum_{\boldsymbol{n},\boldsymbol{n^{\prime}}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})\prod_{i=1}^{m}\gamma_{N}\Big{(}\frac{n_{i}-n^{\prime}_{i}}{N}\Big{)}$
 $\lesssim m!\int_{[0,1]^{2m}}|\mathcal{A}_{N}(\hat{f})(\boldsymbol{t})\mathcal{A}_{N}(\hat{f})(\boldsymbol{t^{\prime}})|\prod\limits_{i=1}^{m}|t_{i}-t_{i}^{\prime}|^{2H-2}{\mathrm{d}}t_{i}{\mathrm{d}}t_{i}^{\prime}$
@@
 Proof.
-First we consider the case where $f$ is an indicator function: $f(x)=\mathrm{I}_{(s,t]}(x)$ with $s,t\in\Omega_{N}$. Then
+First we consider indicator functions of the form $f(x)=\mathrm{I}_{(s,t]}(x)$ with $s,t\in\Omega_{N}$. Then
 $\mathcal{A}_{N}(f)(r)=\mathrm{I}_{(s,t]}(r)\quad\text{and}\quad\mathbf{I}_{1}^{(N)}(f)=N^{-H}\sum_{n=1}^{N}\omega_{n}\mathrm{I}_{(Ns,Nt]}(n),$
 and we have
-$\mathbb{E}_{\omega}\Big{[}\Big{(}\mathbf{I}_{1}^{(N)}(f)\Big{)}^{2}\Big{]}$ $=\mathbb{E}_{\omega}\Big{[}N^{-2H}\Big{(}\sum_{n=1}^{N}\omega_{n}\mathbf{I}_{(Ns,Nt]}(n)\Big{)}^{2}\Big{]}$
-$=N^{-2H}\mathbb{E}_{\omega}\Big{[}\sum_{n,n^{\prime}\in(Ns,Nt]}\omega_{n}\omega_{n^{\prime}}\Big{]}$
-$=N^{-2H}\sum_{n,n^{\prime}\in(Ns,Nt]}\gamma(n-n^{\prime}).$
+$\mathbb{E}_{\omega}\Big[\Big(\mathbf{I}_{1}^{(N)}(f)\Big)^{2}\Big]$ $=\mathbb{E}_{\omega}\Big[N^{-2H}\Big(\sum_{n=1}^{N}\omega_{n}\mathbf{I}_{(Ns,Nt]}(n)\Big)^{2}\Big]=N^{-2H}\sum_{n,n^{\prime}\in(Ns,Nt]}\gamma(n-n^{\prime}).$
 Denoting $r=\frac{n}{N},r^{\prime}=\frac{n^{\prime}}{N}$, we get
@@
 Moreover, for any $k\in\mathbb{N}$ and $f_{1},\cdots,f_{k}$ with $f_{i}\in|\mathcal{H}|^{\otimes l_{i}}$ for $l_{i}\in\mathbb{N}$, we have the joint convergence in distribution
-$\left(\mathbf{I}_{l_{1}}^{(N)}(f_{1}),\cdots,\mathbf{I}_{l_{k}}^{(N)}(f_{k})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{l_{1}}(f_{1}),\cdots,\mathbf{I}_{l_{k}}(f_{k})\right),\quad\text{as}~{}N\rightarrow\infty.$
+$\left(\mathbf{I}_{l_{1}}^{(N)}(f_{1}),\cdots,\mathbf{I}_{l_{k}}^{(N)}(f_{k})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{l_{1}}(f_{1}),\cdots,\mathbf{I}_{l_{k}}(f_{k})\right),\quad\text{as}~N\rightarrow\infty.$ (2.17)
 Proof.
@@
 $\mathbf{I}_{m}^{(N)}(f)$ $=N^{-2mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(h_{1}^{\otimes p_{1}}\otimes\cdots\otimes h_{l}^{\otimes p_{l}})\omega_{\boldsymbol{n}}^{\diamond}$
-$=N^{-2mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(h_{1})^{\otimes p_{1}}\otimes\cdots\otimes\mathcal{A}_{N}(h_{l})^{\otimes p_{l}}\omega_{\boldsymbol{n}}^{\diamond}$
-$=\left(\mathbf{I}_{1}^{(N)}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}^{(N)}(h_{l})\right)^{\diamond p_{l}}$
+$=N^{-2mH}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m}}\mathcal{A}_{N}(h_{1})^{\otimes p_{1}}\otimes\cdots\otimes\mathcal{A}_{N}(h_{l})^{\otimes p_{l}}\omega_{\boldsymbol{n}}^{\diamond}=\left(\mathbf{I}_{1}^{(N)}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}^{(N)}(h_{l})\right)^{\diamond p_{l}}$
 By Lemma 2.5, we have $\mathbb{E}[(\mathbf{I}_{1}^{(N)}(h_{i}))^{2}]\leq C\|h_{i}\|_{|\mathcal{H}|}^{2}$ uniformly in $N$.
-Note that $\mathbf{I}_{1}^{(N)}(h_{i})$ is a centered Gaussian random variable, and thus, we have
+Note that $\mathbf{I}_{1}^{(N)}(h_{i})$ is a centered Gaussian random variable, and thus its higher moments can be expressed via its variance.
+Hence, we have for any $q\in\mathbb{N}$,
 $\mathbb{E}\Big[\Big(\mathbf{I}_{1}^{(N)}(h_{i})\Big)^{2q}\Big]\leq C^{q}(2q-1)!!\|h_{i}\|_{|\mathcal{H}|}^{2q}.$
-Combining the joint convergence in Proposition 2.5, we can apply [19, Lemma B.3] to get
+Combining the joint convergence in Proposition 2.6, we can apply [29, Lemma B.3] to get
 $\mathbf{I}_{m}^{(N)}(f)$ $=\left(\mathbf{I}_{1}^{(N)}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}^{(N)}(h_{l})\right)^{\diamond p_{l}}\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{1}(h_{1})\right)^{\diamond p_{1}}\diamond\cdots\diamond\left(\mathbf{I}_{1}(h_{l})\right)^{\diamond p_{l}},\quad\text{as}~N\rightarrow\infty,$
@@
 3. The Skorohod case
-In Section 3.1, we study the weak convergence for the Wick ordered partition function defined in (1.10) and prove Theorem 1.1 which guarantees the weak convergence of the rescaled Wick ordered partition functions in $L^{2}$-regime when $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$.
-In Section 3.2, we discuss the uniform integrability and characterize the weak limit of the Wick ordered partition functions for the case $\alpha\leq\frac{1}{2}$.
-By Taylor expansion and Wick exponential (2.6), we get
+In Section 3.1, we study the weak convergence for the Wick-ordered partition function defined in (1.11) and prove Theorem 1.1 which guarantees the weak convergence of the rescaled Wick-ordered partition functions in the $L^{2}$-regime when $H\in(\frac{1}{2},1)$ and $\alpha\geq\frac{1}{2}$.
+In Section 3.2, we discuss the uniform integrability and characterize the weak limit of the Wick-ordered partition functions for the case $\alpha\leq\frac{1}{2}$.
+By a Taylor expansion and Wick exponential (2.8), we get
 $\tilde{Z}_{N,\beta}^{\omega,h}$ $=\mathbb{E}_{\tau}\Big[\exp\Big\{h\sum_{x\in\Omega_{N}}\delta_{x}\Big\}\exp\Big\{\diamond\Big(\beta\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big)\Big\}\Big]$ (3.1)
@@
 $\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\overset{(d)}{\longrightarrow}\mathbf{I}_{m}(\psi_{m,k}),\quad\text{as}~N\to\infty.$ (3.7)
+Moreover, for any $p\in\mathbb{N}$ and $m_{1},\cdots,m_{p},k_{1},\cdots,k_{p}\in\mathbb{N}_{0}$, we have the following joint convergence in distribution
+$\left(\mathbf{I}_{m_{1}}^{(N)}(f_{m_{1},k_{1}}^{(N)}),\cdots,\mathbf{I}_{m_{p}}^{(N)}(f_{m_{p},k_{p}}^{(N)})\right)\overset{(d)}{\longrightarrow}\left(\mathbf{I}_{m_{1}}(\psi_{m_{1},k_{1}}),\cdots,\mathbf{I}_{m_{p}}(\psi_{m_{p},k_{p}})\right),\quad\text{as}~N\rightarrow\infty.$ (3.8)
 Here
@@
 Proof.
-We shall focus on the case $11-H<{\alpha}<$ 1 and the proofs for the other two cases are similar.
+We shall focus on the case $1-H<{\alpha}<1$ and the proofs for the other two cases are similar.
 Firstly, observe that $\psi_{m,k}(t_{1},\cdots,t_{m})$ belongs to the space $|\mathcal{H}|^{\otimes m}$ due to the following lemma.
@@
 There exists a constant $C=C_{\alpha,H}$ such that for all $m,k\in\mathbb{N}$,
-$\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}^{2}\leq\frac{C^{m+k}[(m+k)!]^{2H}}{\Gamma\left((m+k)\frac{\alpha+H-1}{H}+1\right)^{2H}}.$
+$\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}^{2}\leq\frac{C^{m+k}[(m+k)!]^{2H}}{\Gamma\big((m+k)\frac{\alpha+H-1}{H}+1\big)^{2H}}.$
 Proof of Lemma 3.2.
@@
 $Y_{m,k}^{(N)}=\mathbf{I}_{m}^{(N)}(\psi_{m,k})-\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})=\mathbf{I}_{m}^{(N)}(\psi_{m,k}-f_{m,k}^{(N)}),$
-and then it suffices to show $Y_{m,k}^{(N)}\rightarrow 0$ in $L^{2}$ as $N\rightarrow\infty$. By Lemma 2.4 and noting the symmetry of $f_{m,k}^{(N)}$ and $\psi_{m,k}$, we get
-$\mathbb{E}\Big{[}\Big{(}Y_{m,k}^{(N)}\Big{)}^{2}\Big{]}\leq m!C^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}.$
-Thus, it remains to show
-$\lim_{N\to\infty}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}=0.$
-Define
+and then it suffices to show $Y_{m,k}^{(N)}\rightarrow 0$ in $L^{2}$ as $N\rightarrow\infty$.
+By Lemma 2.5 and noting the symmetry of $f_{m,k}^{(N)}$ and $\psi_{m,k}$, we get $\mathbb{E}[(Y_{m,k}^{(N)})^{2}]\leq m!C^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}.$ Thus, it remains to show $\lim_{N\to\infty}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}=0.$ Define
 $g_{m,k}^{(N)}(t_{1},\cdots,t_{m+k}):=\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}{\mathbb{P}}\left(\{n_{1},\cdots,n_{m+k}\}\subset\tau\right)\prod_{i=1}^{m+k}\mathbf{I}_{(\frac{n_{i}-1}{N},\frac{n_{i}}{N}]}(t_{i}).$ (3.10)
@@
 $\frac{C_{\alpha}}{A^{\prime}(t-s)^{1-{\alpha}-\epsilon}}\leq\frac{L(N)}{N^{{\alpha}-1}}{\mathbb{P}}_{\tau}\left(N(t-s)\in\tau\right)\leq\frac{A^{\prime}C_{\alpha}}{(t-s)^{1-{\alpha}+\epsilon}},~\text{ for all}~t,s\in\Omega_{N}.$
-Together with (1.6) and (1.7), we have that ($t_{i}$ is unnecessarily belonging to $\Omega_{N}$).
+Together with (1.6) and (1.7), we have that for $t_{i}\in(\frac{n_{i}-1}{N},\frac{n_{i}}{N}],i=1,\dots,m+k$,
 $g_{m,k}^{(N)}(t_{1},\cdots,t_{m+k})\leq C^{m+k}\left(\phi_{m+k}(t_{1},\cdots,t_{m+k})\right)^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}.$ (3.13)
@@
 $\int_{[0,1]^{m}}\left|f_{m,k}^{(N)}-\psi_{m,k}\right|^{\frac{1}{H}}{\mathrm{d}}t_{1}\cdots{\mathrm{d}}t_{m}\leq\int_{[0,1]^{m+k}}\left|g_{m,k}^{(N)}-\phi_{m+k}\right|^{\frac{1}{H}}{\mathrm{d}}t_{1}\cdots{\mathrm{d}}t_{m+k},$
-and thus by Lemma A.4, $\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{L^{1/H}(\mathbb{R}^{m})}^{2}\to 0$, as $N\to\infty$. ∎
+and thus by Lemma A.4, $\|f_{m,k}^{(N)}-\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}^{2}\leq C_{H}^{m}\|f_{m,k}^{(N)}-\psi_{m,k}\|_{L^{1/H}(\mathbb{R}^{m})}^{2}\to 0$, as $N\to\infty$.
+The joint convergence in distribution (3.8) is then a direct consequence of (2.17) in Proposition 2.7 and $Y_{m_{i},k_{i}}^{(N)}\to 0$ in $L^{2}$ as $N\to\infty$ above for each $i=1,\cdots,p$. ∎
 Now we are ready to prove our first main result.
 Proof of Theorem 1.1.
-We assume (A1) and $\alpha>\frac{1}{2}$, i.e., $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$. Note that
+Case (1): We assume (A1) and $\alpha>\frac{1}{2}$, i.e., $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$. Note that
 $\tilde{Z}_{\hat{\beta},\hat{h}}$ $=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod\limits_{i=1}^{k}\diamond\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right)$
 $=\sum_{k=0}^{\infty}\sum_{p=0}^{k}\frac{\hat{\beta}^{p}\hat{h}^{k-p}}{k!}{\binom{k}{p}}\mathbf{I}_{p}\bigg(\int_{[0,1]^{k-p}}\phi_{k}(t_{1},\cdots,t_{k}){\mathrm{d}}t_{p+1}\cdots{\mathrm{d}}t_{k}\bigg)$
-$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbf{I}_{p}\bigg{(}\int_{[0,1]^{q}}\phi_{k}(t_{1},\cdots,t_{k}){\mathrm{d}}t_{p+1}\cdots{\mathrm{d}}t_{k}\bigg{)}$
-$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbf{I}_{p}(\psi_{p,q})=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}(\psi_{m,k}),$
+[…]
 where $\psi_{m,k}$ is given in (3.9).
@@
 $\tilde{Z}_{N}^{(R)}:=\sum_{k+m\leq R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\quad\text{and}\quad\tilde{Z}^{(R)}:=\sum_{k+m\leq R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}(\psi_{m,k}).$
-By Proposition 3.1, we have that $\tilde{Z}_{N}^{(R)}\overset{(d)}{\longrightarrow}\tilde{Z}^{(R)}$ as $N\to\infty$.
+By (3.8) in Proposition 3.1 and the Cramér-Wold Theorem, we have that $\tilde{Z}_{N}^{(R)}\overset{(d)}{\longrightarrow}\tilde{Z}^{(R)}$ as $N\to\infty$.
 To complete the proof, we will apply Lemma A.3.
@@
 Step (1): We have
-$\Big{\|}\tilde{Z}^{(R)}-\tilde{Z}_{\hat{\beta},\hat{h}}\Big{\|}_{L^{2}}$ $=\Big{\|}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\Big{\|}_{L^{2}}\leq\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\Big{\|}\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\Big{\|}_{L^{2}}$
-$\leq C^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\|\psi_{m,k}\|_{|\mathcal{H}|^{\otimes m}}\leq C_{H}^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}$
-$\leq\sum_{k+m>R}(C_{\hat{\beta},\hat{h}}^{\alpha,H})^{m+k}\frac{[(m+k)!]^{H}}{\sqrt{m!}k!\Gamma\left((m+k)\frac{\alpha+H-1}{H}+1\right)^{H}},$
-where the second inequality is due to the Riemann sum approximation and the third inequality follows from Lemma 3.2.
-By Stirling’s formula, each term in the sum on the right-hand side is bounded above by
-$C^{m+k}\frac{(m+k)^{H/2}}{\sqrt{m^{1/2}k}\left((m+k)\frac{\alpha+H-1}{H}\right)^{H/2}}\frac{(m+k)^{(m+k)H}}{m^{m/2}k^{k}(m+k)^{(m+k)(\alpha+H-1)}}\leq C^{m+k}\frac{(m+k)^{(m+k)(1-\alpha)}}{m^{m/2}k^{k}}.$ (3.13)
+$\Big\|\tilde{Z}^{(R)}-\tilde{Z}_{\hat{\beta},\hat{h}}\Big\|_{L^{2}}$ $=\Big\|\sum_{k+m>R}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbf{I}_{m}(\psi_{m,k})\Big\|_{L^{2}}\leq\sum_{k+m>R}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{m!k!}\Big\|\mathbf{I}_{m}(\psi_{m,k})\Big\|_{L^{2}}$ (3.14)
+$\leq C_{H}^{m+k}\sum_{k+m>R}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{\sqrt{m!}k!}\|\phi_{m+k}\|_{L^{1/H}(\mathbb{R}^{m+k})}$
+$\leq\sum_{k+m>R}(C_{\hat{\beta},\hat{h}}^{\alpha,H})^{m+k}\frac{[(m+k)!]^{H}}{\sqrt{m!}k!\Gamma\big((m+k)\frac{\alpha+H-1}{H}+1\big)^{H}},$
+where the second and the last inequality is due to Lemma 3.2. By Stirling’s formula, each term in the sum on the right-hand side is bounded above by
+$C^{m+k}\frac{(m+k)^{H/2}}{\sqrt{m^{1/2}k}\left((m+k)\frac{\alpha+H-1}{H}\right)^{H/2}}\frac{(m+k)^{(m+k)H}}{m^{m/2}k^{k}(m+k)^{(m+k)(\alpha+H-1)}}\leq C^{m+k}\frac{(m+k)^{(m+k)(1-\alpha)}}{m^{m/2}k^{k}},$ (3.15)
 Note that for $a,b,c>0,(a+b)^{c}\leq(2\max\{a,b\})^{c}\leq 2^{c}(a^{c}+b^{c})$. Hence, the right-hand side of (3.15) is bounded above by:
-$C^{m+k}\Big{(}\frac{m^{(1-\alpha)m}m^{(1-\alpha)k}}{m^{m/2}k^{k}}+\frac{k^{(1-\alpha)k}k^{(1-\alpha)m}}{m^{m/2}k^{k}}\Big{)},$
+$C^{m+k}\Big(\frac{m^{(1-\alpha)m}m^{(1-\alpha)k}}{m^{m/2}k^{k}}+\frac{k^{(1-\alpha)k}k^{(1-\alpha)m}}{m^{m/2}k^{k}}\Big)$
 which is summable in $m,k$ assuming $\alpha>\frac{1}{2}$ (this is where the additional condition $\alpha>\frac{1}{2}$ is used) by Lemma A.6.
@@
 Step (2): To prove the uniform convergence $\tilde{Z}_{N}^{(R)}\overset{L^{2}}{\longrightarrow}\tilde{Z}_{N}$ as $R\to\infty$, recalling (3.4) and (3.6),
-$\begin{split}&\Big{\|}\tilde{Z}^{(R)}_{N}-\tilde{Z}_{N}\Big{\|}_{L^{2}}^{2}=\mathbb{E}\Big{[}\Big{(}\sum_{k+m>R}\frac{1}{m!k!}S_{m,k}^{(N)}\Big{)}^{2}\Big{]}\\ =&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}\mathbb{E}\left[\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\mathbf{I}_{m}^{(N)}(f_{m,k^{\prime}}^{(N)})\right]\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}C^{m}m!\|f_{m,k}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}\|f_{m,k^{\prime}}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}.\end{split}$ (3.14)
-where the inequality follows from Cauchy-Schwartz inequality and Lemma 2.4. Recalling (3.9) and (3.10), by Jensen’s inequality and Lemma A.4, then (3.14) is bounded above by
-$\begin{split}&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\|g_{m,k}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\|g_{m,k^{\prime}}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\Big{\|}\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\Big{\|}\phi_{m+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum_{k+m>R}C^{m+k}\frac{\hat{\beta}^{m}\hat{h}^{k}}{\sqrt{m!}k!}\Big{\|}\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m+k})}\sum_{k^{\prime}+m^{\prime}>R}C^{m^{\prime}+k^{\prime}}\frac{\hat{\beta}^{m^{\prime}}\hat{h}^{k^{\prime}}}{\sqrt{m^{\prime}!}k^{\prime}!}\Big{\|}\phi_{m^{\prime}+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big{\|}_{L^{1/H}(\mathbb{R}^{m^{\prime}+k^{\prime}})},\end{split}$
-where the second inequality follows from (3.12) and the term on the right-hand side goes to 0 as $R\to\infty$ by choosing $\epsilon$ sufficiently small such that $(\alpha-\epsilon)+H>1$ and ${\alpha}-\epsilon>\frac{1}{2}$ and the result in Step (1).
-∎
+$\begin{split}\Big\|\tilde{Z}^{(R)}_{N}-\tilde{Z}_{N}\Big\|_{L^{2}}^{2}&=\mathbb{E}\Big[\Big(\sum_{k+m>R}\frac{1}{m!k!}S_{m,k}^{(N)}\Big)^{2}\Big]\\ &=\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{\hat{h}^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}\mathbb{E}\left[\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)})\mathbf{I}_{m}^{(N)}(f_{m,k^{\prime}}^{(N)})\right]\\ &\leq\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{|\hat{h}|^{k+k^{\prime}}\hat{\beta}^{2m}}{k!(k^{\prime})!(m!)^{2}}C^{m}m!\|f_{m,k}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}\|f_{m,k^{\prime}}^{(N)}\|_{|\mathcal{H}|^{\otimes m}}.\end{split}$ (3.16)
+where the inequality follows from the Cauchy- Schwarz inequality and Lemma 2.5. Recalling (3.10) and (3.11), by Jensen’s inequality and Lemma A.4, then (3.16) is bounded above by
+$\begin{split}&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\|g_{m,k}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\|g_{m,k^{\prime}}^{(N)}\|_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum\limits_{k+m>R}\sum\limits_{k^{\prime}+m>R}\frac{C^{2m+k+k^{\prime}}}{k!(k^{\prime})!m!}\Big\|\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m+k})}\Big\|\phi_{m+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m+k})}\\ \leq&\sum_{k+m>R}\frac{C^{m+k}}{\sqrt{m!}k!}\Big\|\phi_{m+k}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m+k})}\sum_{k^{\prime}+m^{\prime}>R}\frac{C^{m^{\prime}+k^{\prime}}}{\sqrt{m^{\prime}!}k^{\prime}!}\Big\|\phi_{m^{\prime}+k^{\prime}}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}\Big\|_{L^{1/H}(\mathbb{R}^{m^{\prime}+k^{\prime}})}.\end{split}$
+where the first inequality follows from (3.13).
+By the same estimate as in (3.14) above, the right-hand side of the last inequality goes to 0 as $R\to\infty$ by choosing $\epsilon$ sufficiently small such that $(\alpha-\epsilon)+H>1$ and ${\alpha}-\epsilon>\frac{1}{2}$.
+Case (2): We assume (A1) and $\alpha=\frac{1}{2}$, i.e., $H\in(\frac{1}{2},1)$ and $\alpha=\frac{1}{2}$.
+The proof is almost the same as in Case (1) and we indicate only the necessary modifications.
+First, noting that by the assumption $0<\inf_{n\in\mathbb{N}}L(n)\leq\sup_{n\in\mathbb{N}}L(n)<+\infty$ for this case, the exponent $\epsilon$ in (3.12) vanishes.
+Then by (3.14), (3.15) and (3.16), it suffices to show that
+$\sum\limits_{m,k=0}^{\infty}\big(\hat{\beta}C_{H}\big)^{m}\big(|\hat{h}|C_{H}\big)^{k}\frac{[(m+k)!]^{H}}{\sqrt{m!}k!\Gamma((m+k)(\frac{2H-1}{2H})+1)^{H}}\leq\sum\limits_{m,k=0}^{\infty}\big(\hat{\beta}C_{H}\big)^{m}\big(|\hat{h}|C_{H}\big)^{k}\frac{(m+k)^{\frac{m+k}{2}}}{m^{m/2}k^{k}}$ (3.17)
+converges for $\hat{\beta}>0,|\hat{h}|>0$ small enough.
+Note that
+$\frac{(m+k)^{\frac{m+k}{2}}}{m^{m/2}k^{k}}\leq\frac{(2k)^{k}}{m^{m/2}k^{k}}\mathrm{I}_{\{k\geq m\}}+\frac{(2m)^{\frac{m+k}{2}}}{m^{m/2}k^{k}}\mathrm{I}_{\{k\leq m\}}\leq\frac{2^{k}}{m^{m/2}}\mathrm{I}_{\{k\geq m\}}+\frac{2^{m}m^{\frac{k}{2}}}{k^{k}}\mathrm{I}_{\{k\leq m\}}.$
+Then the right-hand side of (3.17) is bounded by
+$\begin{split}&\sum\limits_{m=0}^{\infty}\frac{(\hat{\beta}C_{H})^{m}}{m^{m/2}}\sum\limits_{k=m}^{\infty}(2|\hat{h}|C_{H})^{k}+\sum\limits_{m=0}^{\infty}(2\hat{\beta}C_{H})^{m}\sum\limits_{k=0}^{m}\frac{(C_{H}|\hat{h}|\sqrt{m})^{k}}{k!}\\ \leq&\sum\limits_{m=0}^{\infty}\frac{(\hat{\beta}C_{H})^{m}}{m^{m/2}}\cdot\frac{1}{1-2|\hat{h}|C_{H}}+\sum\limits_{m=0}^{\infty}e^{C_{H}|\hat{h}|\sqrt{m}}(2C_{H}\hat{\beta})^{m}<+\infty,\end{split}$
+given that $2C_{H}|\hat{h}|<1$ and $2C_{H}\hat{\beta}<1$.
+Case (3): We assume $\alpha<\frac{1}{2}$. Denote by $\tau^{\prime}$ an independent copy of $\tau$ and we have that
+$\begin{split}\mathbb{E}\Big[\tilde{Z}_{N}^{2}\Big]&=\mathbb{E}_{\tau,\tau^{\prime}}\Big[\exp\Big\{\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}+h_{N}\sum\limits_{n=1}^{N}\big(\mathbf{1}_{\{n\in\tau\}}+\mathbf{1}_{\{n\in\tau^{\prime}\}}\big)\Big\}\Big]\\ &\geq\exp\Big\{\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)-2N|h_{N}|\Big\}{\mathbb{P}}(\{1,\cdots,N\}\subset\tau\cap\tau^{\prime})\\ &\geq\exp\Big(C\hat{\beta}^{2}\frac{1}{N^{2(\alpha+H-1+\epsilon)}}\sum\limits_{n,m=1}^{N}|n-m|^{2H-2}-C|\hat{h}|N^{1-\alpha+\epsilon}\Big)({\mathbb{P}}(\tau_{1}=1))^{2N}\\ &\geq\exp\Big(C\Big(N^{2(1-\alpha-\epsilon)}-N^{1-\alpha+\epsilon}\Big)\Big)({\mathbb{P}}(\tau_{1}=1))^{2N},\end{split}$
+which tends to infinity as $N\to\infty$ if $\alpha<\frac{1}{2}$, by choosing $\epsilon>0$ small enough in (3.12). ∎
 Remark 3.3.
-Note that $\alpha=\frac{1}{2}$ is indeed a criticality. For simplicity, assume $h_{N}\equiv 0$ and $L(\cdot)\equiv 1$ in (1.11). Denote by $\tau^{\prime}$ an independent copy of $\tau$. Then
-$\begin{split}\mathbb{E}\Big{[}\big{(}\tilde{Z}_{N}\big{)}^{2}\Big{]}&=\mathbb{E}_{\tau,\tau^{\prime}}\Big{[}\exp\Big{\{}\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{\}}\Big{]}\\ &\geq\exp\Big{\{}\beta_{N}^{2}\sum\limits_{n,m=1}^{N}\gamma(n-m)\Big{\}}{\mathbb{P}}(\{1,\cdots,N\}\subset\tau\cap\tau^{\prime})\\ &\geq\exp\Big{(}C\hat{\beta}^{2}\frac{1}{N^{2(\alpha+H-1)}}\sum\limits_{n,m=1}^{N}|n-m|^{2H-2}\Big{)}({\mathbb{P}}(\tau_{1}=1))^{2N}\\ &\geq\exp(CN^{2(1-\alpha)})({\mathbb{P}}(\tau_{1}=1))^{2N},\end{split}$
-which tends to infinity as $N\to\infty$ if $\alpha<\frac{1}{2}$. In this case, $\{\tilde{Z}_{N}\}$ is not $L^{2}$ uniformly bounded.
+We now comment on the case $\alpha=\tfrac{1}{2}$, which represents the threshold between the $L^{2}$ regime and the non-$L^{2}$ regime.
+However, this should not be regarded as the marginal case between the disorder-relevant and disorder-irrelevant regimes, since in Section 3.2 we show that the subsequential limit of $\{\tilde{Z}_{N}\}_{N}$ can be nonzero even for some $\alpha<\tfrac{1}{2}$.
+Moreover, the case $\alpha=\tfrac{1}{2}$ can still be treated within the framework of [1, 9], where disorder is relevant, rather than that of [10], where disorder is marginally relevant.
+These two frameworks are fundamentally different.
+In comparison with the parabolic Anderson model on $\mathbb{R}^{2}$, where the $L^{2}$ solution exists only locally while the $L^{1}$ solution exists globally (see [28]), we believe that when $\alpha=\tfrac{1}{2}$, the $L^{2}$ regime of our model exists only locally and the $L^{1}$ regime exists globally (see Remarks 1.41.6, and Section 3.2) We also note that our assumption on $L(\cdot)$ may not be optimal.
+To allow for a general slowly varying function $L(\cdot)$, one would need to verify the dominated convergence theorem carefully.
+Since our main purpose here is to highlight the critical nature of $\alpha=\tfrac{1}{2}$ by a simple argument, we are content with the present restriction on $L(\cdot)$.
 3.2. Some discussion on uniform integrability
-The Weinrib-Halperin prediction suggests that our pinning model should be disorder relevant under (A1), i.e., $H\in(\frac{1}{2},1)$ and $\alpha+H>1$.
-This is partially confirmed by our results in the following sense: Proposition 3.1 yields the weak convergence of each chaos under (A1); while for the weak convergence of the partition functions, an extra condition $\alpha>\frac{1}{2}$ is imposed in order to guarantee the $L^{2}$-convergence of the Wiener chaos expansion (see Step (1) in the proof of Theorem 1.1 and Remark 3.3).
-This implies that when $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$, the family $\{\tilde{Z}_{N}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\}_{N\in\mathbb{N}}$ of the rescaled partition functions is bounded in $L^{2}$-norm and in particular, for the limiting distribution we have $\mathbb{E}[\tilde{Z}_{\hat{\beta},\hat{h}}]=1$.
-Inspired by the Weinrib-Halperin prediction, we conjecture that the family $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is uniformly integrable and converges weakly to an $L^{1}$-integrable limit under (A1).
-In the rest of this section, we shall prove the uniform integrability of $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ and characterize its weak limit, under a stronger condition:
+[…]
+Considering the above discussion together with Remarks 1.6 and 3.3, we conjecture that the family $\{\tilde{Z}_{N}\}_{N}$ converges weakly to some $L^{1}$-integrable random variable under (A1).
+In the rest of this section, we shall prove the uniform integrability of $\{\tilde{Z}_{N}\}_{N}$ and characterize its weak limit, under a stronger condition:
 Assumption (A2).
@@
 Assuming (A2), we have
-$\sup_{N}\beta_{N}^{2}\mathbb{E}\Big{[}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{]}<\infty,$ (3.15)
-and
-$\sup_{N}\mathbb{E}\Big{[}\exp\Big{\{}h_{N}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}\Big{\}}\Big{]}<\infty.$ (3.16)
+$\sup_{N}~\beta_{N}^{2}\mathbb{E}\Big[\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big]<\infty\quad\text{and}\quad\sup_{N}~\mathbb{E}\Big[\exp\Big\{h_{N}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}\Big\}\Big]<\infty.$ (3.18)
 Proof.
-We prove the result for the case $\alpha<1$ and the other cases can be proven in a similarly way. For (3.15), we have
-$\quad\beta_{N}^{2}\mathbb{E}\Big{[}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{]}$
-$=\hat{\beta}^{2}\sum_{n,m=1}^{N}N^{2-2H}\gamma(n-m)\frac{L(N)^{2}}{N^{2{\alpha}-2}}{\mathbb{P}}_{\tau}(n\in\tau,m\in\tau)\frac{1}{N^{2}}$
+We prove the result for the case $\alpha<1$ and the other cases can be proven in a similar way. For the first bound in (3.18), we have
+$\quad\beta_{N}^{2}\mathbb{E}\Big[\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big]$ $=\hat{\beta}^{2}\sum_{n,m=1}^{N}N^{2-2H}\gamma(n-m)\frac{L(N)^{2}}{N^{2{\alpha}-2}}{\mathbb{P}}_{\tau}(n\in\tau,m\in\tau)\frac{1}{N^{2}}$
 $\leq C\int_{0}^{1}\int_{0}^{1}|s-t|^{2H-2}\phi_{2}^{\frac{1-{\alpha}+\epsilon}{1-{\alpha}}}(s,t){\mathrm{d}}s{\mathrm{d}}t<\infty$
-for $\alpha+2H>2$, where $\phi_{2}$ is defined in (1.12) and the inequality follows from (1.6), (1.7) and (3.12). For (3.16), we have that for some $\alpha^{\prime}\in(0,\alpha)$
-$\begin{split}\mathbb{E}\Big{[}\exp\Big{\{}\sum_{n=1}^{N}h_{N}\sigma_{n}\Big{\}}\Big{]}&=\sum\limits_{k=0}^{\infty}\frac{1}{k!}(h_{N})^{k}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{k}}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{k}\in\tau)\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{\Gamma(k\alpha^{\prime})}<\infty,\end{split}$ (3.17)
-where we have used (3.12) and Lemma A.5. ∎
+for $\alpha+2H>2$, where $\phi_{2}$ is defined in (1.13) and the inequality follows from (1.6), (1.7) and (3.13). For the second bound in (3.18), we have that for some $\alpha^{\prime}\in(0,\alpha)$
+$\begin{split}\mathbb{E}\Big[\exp\Big\{\sum_{n=1}^{N}h_{N}\mathbf{1}_{\{n\in\tau\}}\Big\}\Big]&=\sum\limits_{k=0}^{\infty}\frac{1}{k!}(h_{N})^{k}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{k}}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{k}\in\tau)\\ &=\sum\limits_{k=0}^{\infty}\frac{\hat{h}^{k}}{k!}\int_{[0,1]^{k}}g_{0,k}^{(N)}(t_{1},\cdots,t_{k})\prod\limits_{i=1}^{k}{\mathrm{d}}t_{i}\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{\Gamma(k\alpha^{\prime})}<\infty.\end{split}$ (3.19)
+where we have used (3.10), (3.13), (1.13) and Lemma A.5. ∎
 Proposition 3.5.
@@
 Proof.
-Recall that
-$\tilde{Z}_{N}=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{\}}\exp\Big{\{}\diamond\Big{(}\beta_{N}\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big{)}\Big{\}}\Big{]},$
-and we have
+Recall (1.11) and we have
 $\mathbb{E}_{\omega}\Big[\Big(\tilde{Z}_{N}\Big)^{2}\Big]=\mathbb{E}_{\tau,\tau^{\prime}}\Big[\exp\Big\{h_{N}\sum_{n=1}^{N}(1_{\{n\in\tau\}}+\mathbf{1}_{\{n\in\tau^{\prime}\}})\Big\}\exp\Big\{\beta_{N}^{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big\}\Big].$
@@
 $\mathbb{E}_{\omega}\Big[\Big(\tilde{Z}_{N}\Big)^{2}\Big]$ $\leq\mathbb{E}_{\tau,\tau^{\prime}}\bigg[\exp\Big\{h_{N}\sum_{n=1}^{N}1_{\{n\in\tau\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big\}$
-×$\quad\times\exp\Big{\{}h_{N}\sum_{n=1}^{N}1_{\{n\in\tau^{\prime}\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau^{\prime}\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{\}}\bigg{]}.$
+$\quad\times\exp\Big\{h_{N}\sum_{n=1}^{N}1_{\{n\in\tau^{\prime}\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau^{\prime}\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big\}\bigg].$
 Denote
 $A_{l}^{N}:=\Big\{h_{N}\sum_{n=1}^{N}1_{\{n\in\tau\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\leq l\Big\}.$
-Lemma 3.4 and Markov’s inequality together yields:
+Lemma 3.4 and Markov’s inequality together yield:
 $\lim_{l\to\infty}\sup_{N}{\mathbb{P}}\left((A_{l}^{N})^{c}\right)=0.$ (3.20)
@@
 $\tilde{Z}_{N}(A_{l}^{N}):=\mathbb{E}_{\tau}\Big[\exp\Big\{h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\mathbf{1}_{A_{l}^{N}}\Big\}\exp\Big\{\diamond\Big(\beta_{N}\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\mathbf{1}_{A_{l}^{N}}\Big)\Big\}\Big].$
-Then, by Lemma A.7, we get
-$\left|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})\right|\leq\mathbb{E}_{\tau}\left[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\right]+{\mathbb{P}}((A_{l}^{N})^{c}),$
-where we denote
+Then, by Lemma A.7, we get $|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})|\leq\mathbb{E}_{\tau}[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}]+{\mathbb{P}}((A_{l}^{N})^{c})$, where
 $\mathscr{E}_{N}:=\exp\Big\{h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big\}\exp\Big\{\diamond\Big(\beta_{N}\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big)\Big\}.$
 Thus we have
-$\mathbb{E}_{\omega}\left|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})\right|$ $\leq\mathbb{E}_{\omega}\mathbb{E}_{\tau}\left[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\right]+{\mathbb{P}}((A_{l}^{N})^{c})$ (3.19)
-$=\mathbb{E}_{\tau}\Big{[}\mathbf{1}_{(A_{l}^{N})^{c}}\exp\Big{(}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{)}\Big{]}+{\mathbb{P}}((A_{l}^{N})^{c})$
-$\leq{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}\mathbb{E}_{\tau}\Big{(}\Big{[}\exp\Big{(}2h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{)}\Big{]}\Big{)}^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c})$
-$\leq C{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c}),$
+$\mathbb{E}_{\omega}\big|\tilde{Z}_{N}$- $\tilde{Z}_{N}(A_{l}^{N})\big|\leq\mathbb{E}\big[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\big]+{\mathbb{P}}((A_{l}^{N})^{c})=\mathbb{E}_{\tau}\Big[\mathbf{1}_{(A_{l}^{N})^{c}}\exp\Big(h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big)\Big]+{\mathbb{P}}((A_{l}^{N})^{c})$ (3.21)
+$\leq{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}\mathbb{E}_{\tau}\Big(\Big[\exp\Big(2h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big)\Big]\Big)^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c})\leq C{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c}),$
 where the constant $C$ in the last inequality is independent of $N$ by (3.18).
@@
 ∎
-A direct consequence of the uniform integrability is the following.
+In the proof of Proposition 3.5, we applied the Cauchy-Schwarz inequality to decouple $\tau$ and $\tau^{\prime}$, which requires the stronger Assumption (A2).
+It seems that this is a technical issue and (A1) should be sufficient for the uniform integrability.
+A direct consequence of the uniform integrability is the following, which characterizes the possible weak limit of $\{\tilde{Z}_{N}\}_{N}$.
 Proposition 3.6.
@@
 ∎
-In the proof of Proposition 3.5, we apply Cauchy-Schwartz inequality to decouple $\tau$ and $\tau^{\prime}$, which requires the stronger Assumption (A2).
-We believe this is a technical issue and (A1) should be sufficient for the uniform integrability.
-When $\alpha<\frac{1}{2}$, though $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is tight, we are not able to prove the weak convergence as in the proof of Theorem 1.1.
-This is because, when $\alpha<\frac{1}{2}$, $\tilde{Z}_{N}$ is not uniformly bounded in $L^{2}$ (see Remark 3.3) and thus $\mathbb{E}[\mathcal{Z}^{2}]=\infty$ for any weak limit $\mathcal{Z}$.
-Therefore, we cannot employ the theory of Wiener chaos, which applies only to $L^{2}$-integrable random variables, to obtain a weak limit.
 4. The Stratonovich case
@@
 $Z_{\hat{\beta},\hat{h}}$ $=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{[0,1]^{k}}\phi_{k}(t_{1},\cdots,t_{k})\prod_{i=1}^{k}\left(\hat{\beta}W({\mathrm{d}}t_{i})+\hat{h}{\mathrm{d}}t_{i}\right)$ (4.1)
-[…]
+$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbb{I}_{p}\bigg(\int_{[0,1]^{q}}\phi_{k}(t_{1},\cdots,t_{k})dt_{p+1}\cdots dt_{k}\bigg)$
 $=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbb{I}_{p}(\psi_{p,q})=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbb{I}_{m}(\psi_{m,k}).$
@@
 Lemma 4.1.
-Assume (A2). Then we have $\mathrm{Tr}^{j}{\psi}_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $j=0,1,...,\lfloor\frac{m}{2}\rfloor$. Moreover, we have
+Assume (A2). Then we have $\mathrm{Tr}^{j}{\psi}_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $j=0,1,...,\lfloor\frac{m}{2}\rfloor$. Moreover,
 $\|\mathrm{Tr}^{j}{\psi}_{m,k}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}\leq\|\mathrm{Tr}^{j}{\phi}_{m+k}\|^{2}_{|\mathcal{H}|^{\otimes(m+k-2j)}}.$ (4.2)
@@
 We only prove the case $\alpha\in(0,1)$. When $k=0$, we have $\psi_{m,0}=\phi_{m}$ and thus
-$\|\mathrm{Tr}^{j}{\phi}_{m}$ ||($\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg{(}\int_{[0,1]^{2j}}\phi_{m}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
-×$\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\phi_{m}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
-First, we split $[0,1]^{2m}$ into $(2m)!$ regions such that in each region, the numbers $\{t_{i},t_{i}^{\prime},s_{j},s_{j}^{\prime}:i=1,\dots,m-2j,j=1,\dots,2j\}$ are listed in an increasing order.
-To get an upper bound, we replace, for instance, each $|t_{i}-t_{i}^{\prime}|^{2H-2}$ by $|t_{i}^{*}-t_{i}^{\star}|^{2H-2}$ where $t^{*}_{i}:=\max\{t_{i},t_{i}^{\prime}\}$ and $t_{i}^{\star}$ is the largest number that is smaller than $t^{*}_{i}$ (it could be some $t_{j}$, $t^{\prime}_{j}$, $s_{j}$ or $s^{\prime}_{j}$), and we perform the same operation for all the other $|s_{2i}-s_{2i-1}|^{2H-2}$, $|s^{\prime}_{2i}-s^{\prime}_{2i-1}|^{2H-2}$ and the factors in $\phi_{m}$ (see Figures 1 and 2 for illustration).
-Then on each region, the above integral is smaller than an integral with $2m$ variables in Lemma A.5 with $m$ exponents equal to ${\alpha}-1$ and $m$ exponents equal to $32H+{\alpha}-$ 3.
-Noting that $12H+\alpha-3>-$ 1 under (A2), we get by Lemma A.5
+$\|\mathrm{Tr}^{j}{\phi}_{m}$ $\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg(\int_{[0,1]^{2j}}\phi_{m}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
+$\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\phi_{m}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)$
+First, we split $0,1]^{2m}$ into $(2m)!$ regions such that in each region, the numbers $\{t_{i},t_{i}^{\prime},s_{j},s_{j}^{\prime}:i=1,\dots,m-2j,j=1,\dots,2j\}$ are listed in an increasing order.
+To get an upper bound, we replace, for instance, each $|t_{i}-t_{i}^{\prime}|^{2H-2}$ by $|t_{i}^{*}-t_{i}^{\star}|^{2H-2}$ where $t^{*}_{i}:=\max\{t_{i},t_{i}^{\prime}\}$ and $t_{i}^{\star}$ is the largest number that is smaller than $t^{*}_{i}$ (it could be some $t_{j}$, $t^{\prime}_{j}$, $s_{j}$ or $s^{\prime}_{j}$), and we perform the same operation for all the other $|s_{2i}-s_{2i-1}|^{2H-2}$, $|s^{\prime}_{2i}-s^{\prime}_{2i-1}|^{2H-2}$ and the factors in $\phi_{m}$.
+Then on each region, the above integral is smaller than an integral with $2m$ variables in Lemma A.5 with $m$ exponents equal to ${\alpha}-1$ and $m$ exponents equal to $2H+{\alpha}-3$.
+Noting that $2H+\alpha-3>-1$ under (A2), we get by Lemma A.5
 $\|\mathrm{Tr}^{j}{\phi}_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}\leq(2m)!\frac{\Gamma({\alpha})^{m}\Gamma({\alpha}+2H-2)^{m}}{\Gamma(m(2H+2\alpha-2)+1)}<\infty.$ (4.3)
-[…]
-Figure 1.
-An example for $m=6$ and $j=2$ before modifications, the red segments represent the renewal kernels (with exponent $\alpha-1$) and the blue segments represent the kernels of the fractional Brownian motion (with exponent $22H-$ 2).
-[…]
-Figure 2. After modifications, the red segments represent the renewal kernels and the blue segments represent the kernels of the fractional Brownian motion.
-When $k>0$, we have
-$\quad\|\mathrm{Tr}^{j}{\psi}_{m,k}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}$
-$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg{(}\int_{[0,1]^{2j}}\psi_{m,k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
-×$\quad\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\psi_{m,k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
-$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}$
-×$\qquad\times\bigg{(}\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j},x_{1},\cdots,x_{k})\mathrm{d}\boldsymbol{x}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
-×$\qquad\times\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime},x_{1}^{\prime},\cdots,x_{k}^{\prime})\mathrm{d}\boldsymbol{x^{\prime}}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
-$\leq\int_{[0,1]^{2(m+k-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\prod_{p=1}^{k}|x_{p}-x_{p}^{\prime}|^{2H-2}\mathrm{d}x_{p}\mathrm{d}x_{p}^{\prime}$
-×$\quad\times\bigg{(}\int_{[0,1]^{2j}}\phi_{m+k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j},x_{1},\cdots,x_{k})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
-×$\quad\times\int_{[0,1]^{2j}}\phi_{m+k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime},x_{1}^{\prime},\cdots,x_{k}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
-$=\|\mathrm{Tr}^{j}{\phi}_{m+k}\|^{2}_{|\mathcal{H}|^{\otimes(m+k-2j)}}.$
+When $k>0$, by denoting $\boldsymbol{s}:=(s_{1},\cdots,s_{2j})$, $\boldsymbol{t}:=(t_{1},\cdots,t_{m-2j})$, $\boldsymbol{x}:=(x_{1},\cdots,x_{k})$ and $\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime},\boldsymbol{x}^{\prime}$ correspondingly, we have
+$\quad\|\mathrm{Tr}^{j}{\psi}_{m,k}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg(\int_{[0,1]^{2j}}\psi_{m,k}(\boldsymbol{s},\boldsymbol{t})$
+$\qquad\qquad\qquad\qquad\qquad\times\int_{[0,1]^{2j}}\psi_{m,k}(\boldsymbol{s}^{\prime},t_{1}^{\prime},\boldsymbol{t}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)$
+$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg(\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(\boldsymbol{s},\boldsymbol{t},\boldsymbol{x})\mathrm{d}\boldsymbol{x}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime},\boldsymbol{x}^{\prime})\mathrm{d}\boldsymbol{x^{\prime}}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)$
+$\leq\int_{[0,1]^{2(m+k-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\prod_{p=1}^{k}|x_{p}-x_{p}^{\prime}|^{2H-2}\mathrm{d}x_{p}\mathrm{d}x_{p}^{\prime}\bigg(\int_{[0,1]^{2j}}\phi_{m+k}(\boldsymbol{s},\boldsymbol{t},\boldsymbol{x})$
+$\quad\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\phi_{m+k}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime},\boldsymbol{x}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)=\|\mathrm{Tr}^{j}{\phi}_{m+k}\|^{2}_{|\mathcal{H}|^{\otimes(m+k-2j)}}.$
 This proves (4.2) and hence $\mathrm{Tr}^{j}{\psi}_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ from the first step. ∎
 Remark 4.2.
-Assuming (A2), Lemma 4.1 yields ${\vvvert\psi_{m,k}\vvvert_{m}<\infty}$ (see (2.9)) and hence $\mathbb{I}_{m}(\psi_{m,k})$ is well defined due to Hu-Meyer’s formula (2.8).
-On the other hand, $\alpha+2H>2$ is the optimal condition for ${\vvvert\psi_{m,k}\vvvert_{m}<\infty}$.
+Assuming (A2), Lemma 4.1 yields $\interleave\psi_{m,k}\interleave_{m}<\infty$} (see (2.11)) and hence $\mathbb{I}_{m}(\psi_{m,k})$ is well defined due to Hu–Meyer formula (2.10).
+On the other hand, $\alpha+2H>2$ is the optimal condition for $\interleave\psi_{m,k}\interleave_{m}<\infty$}.
 Indeed, for $k=0,m=2$ and $j=1$, we have
-[…]
+$\mathrm{Tr}^{1}\phi_{2}=2\int_{0<t_{1}<t_{2}<1}\frac{C_{\alpha}^{2}}{t_{1}^{1-{\alpha}}(t_{2}-t_{1})^{1-{\alpha}}}(t_{2}-t_{1})^{2H-2}dt_{1}dt_{2},$
 which is finite if and only if ${\alpha}+2H>2$.
@@
 Proof.
-Recalling
-$Z_{\hat{\beta},\hat{h}}=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbb{I}_{m}(\psi_{m,k}),$
-we have
-$\|Z_{\hat{\beta},\hat{h}}\|_{2}$ $\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\|\mathbb{I}_{m}(\psi_{m,k})\|_{2}$
-$\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\sqrt{2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}}$
+Recalling $Z_{\hat{\beta},\hat{h}}$ from (4.1), we have
+$\|Z_{\hat{\beta},\hat{h}}\|_{2}$ $\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{m!k!}\|\mathbb{I}_{m}(\psi_{m,k})\|_{2}\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{m!k!}\sqrt{2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}}$
 $\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{C^{m+k}}{\sqrt{m!}k!}\sqrt{\frac{[2(m+k)]!}{\Gamma((m+k)(2H+2\alpha-2)+1)}}\,$,
-where the second inequality follows from Lemma 4.4 below and the last inequality is due to (4.2) and (4.3) below.
+where the second inequality follows from Lemma 4.4 below and the last inequality is due to (4.2) and (4.3) above.
 Note that $\alpha+H>\frac{3}{2}$ implies $\zeta=4-2H-2\alpha<1$. By Stirling’s formula, we get
@@
 This together with Lemma A.6 yields $\|Z_{\hat{\beta},\hat{h}}\|_{2}<\infty$. ∎
-Lemma 4.4.
+Lemma $4.$
 Recall $\phi$ in (1.13) and $\psi$ in (3.9). We have
-$\mathbb{E}[\mathbb{I}_{m}(\phi_{m})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\phi}_{2m},$
-and
-$\mathbb{E}[\mathbb{I}_{m}(\psi_{m,k})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}.$
+$\mathbb{E}[\mathbb{I}_{m}(\phi_{m})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\phi}_{2m}\quad\text{and}\quad\mathbb{E}[\mathbb{I}_{m}(\psi_{m,k})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}.$
 Proof.
-For the first inequality, Hu-Meyer’s formula (2.8) implies
-$\mathbb{E}[|\mathbb{I}_{m}(\phi_{m})|^{2}]$ $=\sum_{k=0}^{[\frac{m}{2}]}\frac{1}{(m-2k)!}\left(\frac{m!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.4)
-$=\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}\binom{m}{2k}\binom{2k}{k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}\leq\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}2^{m+2k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$
-$=2^{m}m!\sum_{k=0}^{[\frac{m}{2}]}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}},$
+For the first inequality, Hu–Meyer formula (2.10) implies
+$\mathbb{E}[|\mathbb{I}_{m}(\phi_{m})|^{2}]$ $=\sum_{k=0}^{[\frac{m}{2}]}\frac{1}{(m-2k)!}\left(\frac{m!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}=\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}\binom{m}{2k}\binom{2k}{k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.4)
+$\leq\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}2^{m+2k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}=2^{m}m!\sum_{k=0}^{[\frac{m}{2}]}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}},$
 where the inequality follows from the fact $\binom{m}{n}\leq 2^{m}$. For each term on the right-hand side, we have
-$\|\mathrm{Tr}^{k}{\phi}_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.5)
-$=$ $\int_{[0,1]^{2m}}\prod_{i=1}^{m-2k}|t_{i}-t_{i}^{\prime}|^{2H-2}\prod_{j=1}^{k}|s_{2j}-s_{2j-1}|^{2H-2}|s_{2j}^{\prime}-s_{2j-1}^{\prime}|^{2H-2}$
-×$\times\quad\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}){\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{s^{\prime}}{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{t^{\prime}}$
-$=$ $\mathrm{Tr}^{m}{\Phi}_{m}\leq\mathrm{Tr}^{m}{\phi}_{2m},$
-where
-$\Phi_{m}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime})$
-::=:= $\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime})$
-$\leq$ ${\phi}_{2m}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime}).$
-Substituting (4.5) into (4.4) yields the desired inequality.
+$\|\mathrm{Tr}^{k}{\phi}_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}=$ $\int_{[0,1]^{2m}}\prod_{i=1}^{m-2k}|t_{i}-t_{i}^{\prime}|^{2H-2}\prod_{j=1}^{k}|s_{2j}-s_{2j-1}|^{2H-2}|s_{2j}^{\prime}-s_{2j-1}^{\prime}|^{2H-2}$ (4.5)
+$\times\quad\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}){\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{s^{\prime}}{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{t^{\prime}}$
+$\leq$ $\mathrm{Tr}^{m}{\phi}_{2m}.$
+For the last inequality above, denoting $\boldsymbol{t}:=(t_{1},\cdots,t_{m-2k})$, $\boldsymbol{s}:=(s_{1},\cdots,s_{2k})$ and $\boldsymbol{t}^{\prime},\boldsymbol{s}^{\prime}$ correspondingly, we have that by the symmetry and the expression of $\phi$,
+$\phi_{m}(\boldsymbol{s},\boldsymbol{t})\phi_{m}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})\leq\phi_{2m}(\boldsymbol{s},\boldsymbol{s}^{\prime},t_{1},t^{\prime}_{1},\cdots,t_{m-2k},t^{\prime}_{m-2k}),$
+since each subinterval among $(\boldsymbol{s},\boldsymbol{t})$ and $(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})$ is shortened by the interlacement between $(\boldsymbol{s},\boldsymbol{t})$ and $(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})$.
+Also note that the kernel $K(\cdot)$ in the trace (2.9) is not influenced by the permutation in $\phi$.
+Then substituting (4.5) into (4.4) yields the desired inequality.
 Similar to the above, noting that
-$\quad\psi_{2m,2k}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime})$
-$\geq\psi_{m,k}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\psi_{m,k}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}),$
+$\psi_{m,k}(\boldsymbol{s},\boldsymbol{t})\psi_{m,k}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})\leq\psi_{2m,2k}(\boldsymbol{s},\boldsymbol{s}^{\prime},t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime}),$
 we can prove the second inequality in the same way. ∎
@@
 First, note that Taylor expansion of the partition function $Z_{N}:=Z_{N,\beta_{N}}^{\omega,h_{N}}$ given in (1.9) yields
-$Z_{N}$ $=\mathbb{E}_{\tau}\bigg{[}\Big{(}\sum_{k=0}^{\infty}\frac{1}{k!}\Big{(}\sum_{n=1}^{N}h_{N}\sigma_{n}\Big{)}^{k}\Big{)}\Big{(}\sum_{m=0}^{\infty}\frac{1}{m!}\Big{(}\sum_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{m}\Big{)}\bigg{]}$ (4.6)
-$=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m},\cdots,n_{m+k}\in\tau)\prod_{i=1}^{m}\omega_{n_{i}},$
-$=:\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\mathbb{S}_{m,k}^{(N)}.$
+$Z_{N}$ $=\mathbb{E}_{\tau}\bigg[\Big(\sum_{k=0}^{\infty}\frac{1}{k!}\Big(\sum_{n=1}^{N}h_{N}\delta_{\frac{n}{N}}\Big)^{k}\Big)\Big(\sum_{m=0}^{\infty}\frac{1}{m!}\Big(\sum_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{m}\Big)\bigg]$ (4.6)
+$=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m},\cdots,n_{m+k}\in\tau)\prod_{i=1}^{m}\omega_{n_{i}}=:\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\mathbb{S}_{m,k}^{(N)}.$
 We then prove the weak convergence of each chaos.
@@
 […]
-$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}]$
-$=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}].$
+$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{B\subset\llbracket m\rrbracket,~|B|=m-2j}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}]$
+$=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{B\subset\llbracket m\rrbracket,~|B|=m-2j}\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}].$
 For each $j=0,1,\cdots,\lfloor\frac{m}{2}\rfloor$, all the terms in the summation $\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket,|B|=m-2j\end{subarray}}$ are equal by symmetry.
-Fix $j\in\{0,1,...,\lfloor\frac{m}{2}\rfloor\}$ and note that there are in total $\binom{m}{m-2j}$ subsets of $\llbracket m\rrbracket$ with cardinality $m-2j$.
+Fix $j\in\{0,1,...,\lfloor\frac{m}{2}\rfloor\}$ and note that there are in total $\binom{m}{m-2j}$ subsets of $\llbracket m\rrbracket$ with cardinality m2jm-2j.
 Without loss of generality, we may assume $B=\{2j+1,...,m\}$.
-Then, we have
+By the Wick theorem (see, e.g., [19, Theorem 1.28]), we have
 $\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}]=\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket 2j\rrbracket}}]=\sum_{V}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{n_{l_{1}}}\omega_{n_{l_{2}}}],$
@@
 $\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{n_{l_{1}}}\omega_{n_{l_{2}}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)$
-equal to each other for all pair partitions $V$. Note that there are in total $(2j-1)!!$! pair partitions of $\llbracket 2j\rrbracket$, and thus
-$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\binom{m}{m-2j}(2j-1)!!\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)$ (4.9)
+equal to each other for all pair partitions $V$. Since there are in total (2j1)!!(2j-1)!! pair partitions of $\llbracket 2j\rrbracket$,
+$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\binom{m}{m-2j}(2j-1)!\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)$ (4.9)
 $=\sum_{j=0}^{[\frac{m}{2}]}\frac{m!}{j!(m-2j)!2^{j}}\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau).$
-Compare (4.9) with Hu-Meyer’s formula (2.8) and note that $\mathrm{Tr}^{j}\psi_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $k=0,1,...,\lfloor\frac{m}{2}\rfloor$ by Lemma 4.1.
+Compare (4.9) with Hu–Meyer formula (2.10) and note that $\mathrm{Tr}^{j}\psi_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $k=0,1,...,\lfloor\frac{m}{2}\rfloor$ by Lemma 4.1.
 From now on, without loss of generality, we assume that $m$ is odd and the case of even $m$ can be treated in the same manner.
@@
 $f_{\mathrm{Tr}^{j},m,k}^{(N)}=f_{\mathrm{Tr}^{j},m,k}^{(N)}(t_{1},\cdots,t_{m-2j})$ (4.12)
-$:=N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}\sum_{\boldsymbol{u}\in\llbracket N\rrbracket^{2j}}\sum_{\boldsymbol{q}\in\llbracket N\rrbracket^{m-2j}}\sum_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})\prod_{i=1}^{m-2j}\mathbf{1}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),$
-where we recall that $\mathbf{I}_{m-2j}^{(N)}$ is defined in (2.11), $P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})$ is defined in (3.5), and note that
-$\frac{L(N)^{m+k}}{N^{m(\alpha-1+H)+\alpha k}}=N^{-(m-2j)H}N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}.$
-For $j=0$, we have $Y_{N}=\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)}-\psi_{m,k})$, which converges to 00 in $L^{2}$ as $N\to\infty$ as has been proved in the proof of Proposition 3.1.
+$=N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}\!\!\sum_{\boldsymbol{u}\in\llbracket N\rrbracket^{2j}}\sum_{\boldsymbol{q}\in\llbracket N\rrbracket^{m-2j}}\sum_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})\prod_{i=1}^{m-2j}\mathbf{1}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),$
+where we recall that $\mathbf{I}_{m-2j}^{(N)}$ is defined in (2.13), $P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})$ is defined in (3.5).
+For $j=0$, we have $Y_{N}=\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)}-\psi_{m,k})$, which converges to 0 in $L^{2}$ as $N\to\infty$ as has been proved in the proof of Proposition 3.1.
 For $j=1,...,\frac{m-1}{2}$, by Lemma 2.5 and Lemma A.4, we get
@@
 $g_{\mathrm{Tr}^{j},m,k}^{(N)}(t_{1},\cdots,t_{m+k})\leq C^{m+k}\prod_{l=1}^{j}|t_{2l-1}-t_{2l}|^{2H-2}\big(\phi_{m+k}(t_{1},\cdots,t_{m+k})\big)^{\frac{1-{\alpha}^{\prime}}{1-{\alpha}}}.$ (4.16)
-By a similar argument used in the proof of Lemma 4.1, we can show that the right-hand side of (4.16) belongs to $L^{1/H}$. This is provides the desired $L^{1/H}$ domination.
+By a similar argument used in the proof of Lemma 4.1, we can show that the right-hand side of (4.16) belongs to $L^{1/H}$. This provides the desired $L^{1/H}$ domination.
 Finally, the proof is completed by the continuous mapping theorem and (4.10). ∎
 4.3. $L^{1}$-bound of the rescaled partition functions
-In this subsection, we aim to prove the uniform $L^{1}$-boundedness, under Assumption (A3), for the partition function $Z_{N}$.
-Recall its Taylor expansion (4.6) and by Cauchy-Schwartz inequality, we have that
-$\begin{split}\mathbb{E}[Z_{N}]\leq\sum\limits_{k=0}^{\infty}\sum\limits_{m=0}^{\infty}\frac{1}{k!}\frac{1}{m!}\mathbb{E}_{\tau}\Big{[}\Big{(}\sum\limits_{n=1}^{N}h_{N}\sigma_{n}\Big{)}^{2k}\Big{]}^{\frac{1}{2}}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}.\end{split}$ (4.17)
+In this subsection, we prove the uniform $L^{1}$-boundedness, under Assumption (A3), for the partition function $Z_{N}$.
+Recall its Taylor expansion (4.6) and by the Cauchy-Schwarz inequality, we have that
+$\begin{split}\mathbb{E}[Z_{N}]\leq\sum\limits_{k=0}^{\infty}\sum\limits_{m=0}^{\infty}\frac{1}{k!}\frac{1}{m!}\mathbb{E}_{\tau}\Big[\Big(\sum\limits_{n=1}^{N}h_{N}\delta_{\frac{n}{N}}\Big)^{2k}\Big]^{\frac{1}{2}}\mathbb{E}\Big[\Big(\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{2m}\Big]^{\frac{1}{2}}.\end{split}$ (4.17)
 Using the same argument as in (3.19), we get that for some $\alpha^{\prime}\in(0,\alpha)$,
-$\sum\limits_{k=0}^{\infty}\frac{1}{k!}\mathbb{E}_{\tau}\Big{[}\Big{(}\sum\limits_{k=0}^{\infty}h_{N}\sigma_{N}\Big{)}^{2k}\Big{]}^{\frac{1}{2}}\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k!}\sqrt{\frac{(2k)!}{\Gamma(2k\alpha^{\prime})}}\lesssim\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k^{k\alpha^{\prime}}}<\infty.$ (4.18)
+$\sum\limits_{k=0}^{\infty}\frac{1}{k!}\mathbb{E}_{\tau}\Big[\Big(\sum\limits_{n=1}^{N}h_{N}\delta_{\frac{n}{N}}\Big)^{2k}\Big]^{\frac{1}{2}}\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k!}\sqrt{\frac{(2k)!}{\Gamma(2k\alpha^{\prime})}}\lesssim\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k^{k\alpha^{\prime}}}<\infty.$ (4.18)
 For each term in the summation over $m$ in (4.17), we have that
-$\begin{split}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}=\Big{(}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{2m}}\beta_{N}^{2m}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{2m}\in\tau)\mathbb{E}_{\omega}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]\Big{)}^{\frac{1}{2}}.\end{split}$ (4.19)
+$\begin{split}\mathbb{E}\Big[\Big(\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{2m}\Big]^{\frac{1}{2}}=\Big(\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{2m}}\beta_{N}^{2m}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{2m}\in\tau)\mathbb{E}_{\omega}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]\Big)^{\frac{1}{2}}.\end{split}$ (4.19)
 Similar to the proof of Proposition 4.5, $\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]=\sum_{V}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{l_{1}}\omega_{l_{2}}]$, where $V$ is all pair partitions of $\llbracket 2m\rrbracket$.
-Note that $|V|=(2m-1)!!$!.
+Note that |V|=(2m1)!!|V|=(2m-1)!!.
 By (1.4) and (3.12), a Riemann sum approximation and the kernel modification procedure in the proof of Lemma 4.1, (4.19) is bounded above by
@@
 for some $\alpha^{\prime}<\alpha$ and $\alpha^{\prime}+H>\frac{3}{2}$. Hence, we have that
-$\sum\limits_{m=0}^{\infty}\frac{1}{m!}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}\leq\sum\limits_{m=0}^{\infty}\frac{C^{m}}{m^{m(\alpha^{\prime}+H-\frac{3}{2})}}<\infty.$ (4.21)
+$\sum\limits_{m=0}^{\infty}\frac{1}{m!}\mathbb{E}\Big[\Big(\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{2m}\Big]^{\frac{1}{2}}\leq\sum\limits_{m=0}^{\infty}\frac{C^{m}}{m^{m(\alpha^{\prime}+H-\frac{3}{2})}}<\infty.$ (4.21)
 Combining (4.17), (4.18) and (4.21) we conclude the $L^{1}$-boundedness of $Z_{N}$ and we can also deduce from the proof that
-$\lim_{M\to\infty}\sup_{N\in\mathbb{N}}\sum_{l=M}^{\infty}\sum_{m+k=l}\frac{1}{k!m!}\mathbb{E}\Big{[}\Big{|}h_{N}\sum_{n=1}^{N}\sigma_{n}\Big{|}^{k}\Big{|}\beta_{N}\sum_{n=1}^{N}\omega_{n}\sigma_{n}\Big{|}^{m}\Big{]}=0.$ (4.22)
+$\lim_{M\to\infty}\sup_{N\in\mathbb{N}}\sum_{l=M}^{\infty}\sum_{m+k=l}\frac{1}{k!m!}\mathbb{E}\Big[\Big|h_{N}\sum_{n=1}^{N}\delta_{\frac{n}{N}}\Big|^{k}\Big|\beta_{N}\sum_{n=1}^{N}\omega_{n}\delta_{\frac{n}{N}}\Big|^{m}\Big]=0.$ (4.22)
 Now we are ready to prove our second main result.
@@
 Appendix A Some miscellaneous results
+The following two lemmas about the Wick product can be found in [19, Corollary 3.17, Theorem 3.9]. For Lemma A.1, see also [29, equation (A.5)].
 Lemma A.1.
 Let $(\omega_{n})_{n\in\mathbb{N}}$ be Gaussian. Then
-$\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}=\sum_{j=0}^{[\frac{m}{2}]}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}],$
+[…]
 where $\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}:=\prod_{i=1}^{m}\omega_{n_{i}}$ is an ordinary product and $\omega_{\boldsymbol{n}_{B}}^{\diamond}:=\prod_{n_{i}\in B}\diamond\omega_{n_{i}}$ is a Wick product.
@@
 where $\pi$ is a pairwise partition of the set $\{n_{1},\dots,n_{p},m_{1},\dots,m_{p}\}$ such that for each $\{i,j\}\in\pi$, $i\in\{n_{1},\dots,n_{p}\}$ and $j\in\{m_{1},\dots,m_{p}\}$.
+The following lemma about the commutative diagram of weak convergence can be found in [6, Theorem 3.2, Chapter 1].
 Lemma A.3.
 Consider random variables $Y_{n}^{(N)},Y^{(N)},Y_{n}$ and $Y$ such that $Y^{(N)}_{n}$ converges weakly to $Y_{n}$ as $N\to\infty$, $Y_{n}$ converges weakly to $Y$ as $n\to\infty$, and $Y_{n}^{(N)}$ converges in probability to $Y^{(N)}$ uniformly in $N$ as $n\to\infty$, then we have $Y^{(N)}$ converges weakly to $Y$ as $N\to\infty$.
-The following Hardy-Littlewood-Sobolev inequality was proved in [14].
+The following inequality is borrowed from [23].
 Lemma A.4.
-For $H\in(\frac{1}{2},1]$, the following inequality holds:
+For $H\in(\frac{1}{2},1]$, there exists a constant $C_{H}$ depending only on $H$, such that,
 $\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(\boldsymbol{t})f(\boldsymbol{s})\prod_{i=1}^{m}|t_{i}-s_{i}|^{2H-2}{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{s}\leq C_{H}^{m}\|f\|_{L^{1/H}(\mathbb{R}^{m})}^{2},$
-where $C_{H}$ is a constant depending only on $H$.
-The following result can be obtained by direct computation.
+The following result can be obtained by a direct computation.
 Lemma A.5.
@@
 $\sum_{k+m=1}^{\infty}A^{m+k}\frac{m^{ak}}{m^{bm}k^{ck}}\leq\Big(\sum_{k<k_{0}}\sum_{m\geq 1}+\sum_{m<m_{0}}\sum_{k\geq 1}+\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}\Big)A^{m+k}\frac{m^{ak}}{m^{bm}k^{ck}}.$
-It is not hard to see that the first two sums are finite. For the third term, note that $(bm)^{ck}(ck)^{bm}\leq(bm)^{bm}(ck)^{ck}$. Let $d=c-a$ and we have that
+It is not hard to see that the first two sums are finite. For the third term, note that $(bm)^{ck}(ck)^{bm}\leq(bm)^{bm}(ck)^{ck}$. Let d=cad=c-a and we have that
 $\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}A^{m+k}\frac{m^{ak}}{m^{bm}k^{ck}}\leq\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}A^{m+k}\frac{m^{ak}}{m^{ck}k^{bm}}\Big(\frac{b}{c}\Big)^{bm}\Big(\frac{c}{b}\Big)^{ck}$
-$=$ $\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}\frac{1}{\left(C_{A}m\right)^{dk}\left(C^{\prime}_{A}k\right)^{bm}}\leq\sum_{m\geq m_{0}}\frac{1}{\left(C^{\prime}_{A}k_{0}\right)^{bm}}\sum_{k\geq k_{0}}\frac{1}{\left(C_{A}m_{0}\right)^{dk}}<\infty,$
+$=$ $\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}\frac{1}{(C_{A}m)^{dk}(C^{\prime}_{A}k)^{bm}}\leq\sum_{m\geq m_{0}}\frac{1}{(C^{\prime}_{A}k_{0})^{bm}}\sum_{k\geq k_{0}}\frac{1}{(C_{A}m_{0})^{dk}}<\infty,$
 by choosing $C_{A}^{\prime}k_{0}>1$ and $C_{A}m_{0}>1$. The proof is now completed. ∎
 Lemma A.7.
-Let $X$ be an exponentially integrable random variable on $(\Omega,\mathcal{F},{\mathbb{P}})$ and $A\in\mathcal{F}$. Then we have
-$\mathbb{E}\left[e^{X}-e^{X\mathbf{1}_{A}}\right]={\mathbb{P}}(A^{c})\left(\mathbb{E}\left[e^{X}\Big{|}A^{c}\right]-1\right)=\mathbb{E}\left[e^{X}\mathbf{1}_{A^{c}}\right]-{\mathbb{P}}(A^{c}).$
+Let $X$ be an exponentially integrable random variable on $(\Omega,\mathcal{F},{\mathbb{P}})$ and $A\in\mathcal{F}$. Then
+$\mathbb{E}\left[e^{X}-e^{X\mathbf{1}_{A}}\right]={\mathbb{P}}(A^{c})\Big(\mathbb{E}\Big[e^{X}\Big|A^{c}\Big]-1\Big)=\mathbb{E}\left[e^{X}\mathbf{1}_{A^{c}}\right]-{\mathbb{P}}(A^{c}).$
 Proof.
@@
 $\mathbb{E}\left[f(X)-f(X\mathbf{1}_{A})\right]$ $={\mathbb{P}}(A)\mathbb{E}[f(X)-f(X\mathbf{1}_{A})|A]+{\mathbb{P}}(A^{c})\mathbb{E}[f(X)-f(X\mathbf{1}_{A})|A^{c}]$
-$={\mathbb{P}}(A)\cdot 0+{\mathbb{P}}(A^{c})\mathbb{E}[f(X)-f(0)|A^{c}]$
-$=\mathbb{E}[f(X)\mathbf{1}_{A^{c}}]-f(0){\mathbb{P}}(A^{c}),$
+$={\mathbb{P}}(A)\cdot 0+{\mathbb{P}}(A^{c})\mathbb{E}[f(X)-f(0)|A^{c}]=\mathbb{E}[f(X)\mathbf{1}_{A^{c}}]-f(0){\mathbb{P}}(A^{c}),$
 which yields the desired result by choosing $f(x)=e^{x}$. ∎
@@
 Proof.
-By triangular inequality and Cauchy-Schwartz inequality, we get
+By the triangular inequality and the Cauchy-Schwarz inequality, we get
 $\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}|$ $\leq\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}-Y^{\varepsilon}_{n}\mathbf{1}_{[Y_{n}>k]}|+\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y_{n}>k]}-Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|+\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|$
@@
 $\leq\varepsilon+\sqrt{M_{\varepsilon}}\big[{\mathbb{P}}(Y_{n}>k)^{\frac{1}{2}}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{\frac{1}{2}}\big]+\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|.$
-For any fixed $\varepsilon\in(0,1)$, we may find $K=K(\varepsilon)$ such that for all $k>K$ and $n\in\mathbb{N}$,
-${\mathbb{P}}(Y_{n}>k)^{\frac{1}{2}}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{\frac{1}{2}}<\frac{\varepsilon}{\sqrt{M_{\varepsilon}}}\quad\text{and}\quad\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|<\varepsilon$
-by Markov’s inequality and Cauchy-Schwartz inequality.
-Therefore, for all $k>K$ and $n\in\mathbb{N}$, $\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}|<3\varepsilon$.
+For any fixed $\varepsilon\in(0,1)$, we may find $K=K(\varepsilon)$ such that for all $k>K$ and $n\in\mathbb{N}$, ${\mathbb{P}}(Y_{n}>k)^{1/2}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{1/2}<\varepsilon(M_{\varepsilon})^{-1/2}$ and $\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|<\varepsilon$ by Markov’s inequality and the Cauchy-Schwarz inequality.
+Hence, $\sup_{k\geq K,{n}\in\mathbb{N}}\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}|<3\varepsilon$.
 This proves the uniform integrability of $\{Y_{n},n\in\mathbb{N}\}$.
 ∎
-Acknowledgements.
+Acknowledgments
+We thank Quentin Berger, Rongfeng Sun, and Jinjiong Yu for valuable suggestions on an earlier version of the paper.
+We also thank Julien Poisat for helpful comments and references.
+We are further grateful to the referees for their suggestions and comments, which have significantly improved the paper.
+funding
 J.
-Song is partially supported by NSFC Tianyuan Key Program Project (No.
-12226001), National Natural Science Foundation of China (No.
-12471142), and the Fundamental Research Funds for the Central Universities.
+Song is partially supported by NSFC (No.
+12471142) and the Fundamental Research Funds for the Central Universities.
 R.
-Wei is supported by Xi’an Jiaotong-Liverpool University Research Development Fund (No.
-RDF-23-01-024) and NSFC (No.
-12401170).
-The authors thank Quentin Berger, Rongfeng Sun and Jinjiong Yu for reading the first version of this paper and providing invaluable suggestions.
+Wei is supported by NSFC (No.
+12401170) and Xi’an Jiaotong-Liverpool University Research Development Fund (No.
+RDF-23-01-024).